A Complete Bibliography of Publications in Annals of Statistics (1990–1999)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 March 2019 Version 1.02

Title word cross-reference

1/2 [BZ98]. 2 [GG99b]. 2^{m} [Gho80, Gho91]. 2^{n−k} [Che92a, TW96]. 2^{n−l} [CH96]. 2^{n−m} [CC99]. 2d + 4 [Cas96]. 2s [Swa99]. [Ar(p)] [KS93]. α [PR92, YZ93]. B [She95]. χ̄^{2} [TK97]. c [Det96]. C_{L} [Gir91]. D [Det90, Det93a, Det96, RSW97, DS95b]. D_{1} [Det90]. E [CH91, DS93, Det93b, Hei94, PS93]. G [DW95]. G_{2} [DT99]. k [CAGM97, GEGM99, SV98]. L [Hös96, PK89, PK90, Sha94, Xia94]. L^{∞} [Yu93]. L^{p} [GH95b]. L_{1} [Dat92, GHL99, Kni98]. L_{2} [CAMRRdB99]. L_{p} [CGH91, Hor91, Yin90]. M [CK95, CV92, Gey94, HJ90, HS96a, KT91, KT96a, Kol97, KS97, Lah92, LY94, LZ98b, MM99, Nie92, Nis99, Por85, Por91, RG97, Wel89, Wel90]. M_{m} [Bos98]. R^{d} [Cas96]. P [HR99]. MTP_{2} [Sar98]. n [JMP91, Wu95b]. n^{−1/2} [YD97]. P [BRW91, Dud94, Mar91a, Men94b, Por85, Por91]. p^{2}/n [Por85, Por91]. p × 2 [UM97]. P {X < Y } [YG95]. q [Rub96]. q^{n−m} [SCW97]. R [Cha92, KS93]. S [Dav90, Roc96]. s^{n−k} [CW91]. t [GC92]. T^{2} [CW99, MP90, Pin94]. U [AG92, ACG94, Arc95, Arc96, Cha92, Hel91, HS96b, HJ93, She94, Stu94]. V [AG92]. ε [HF92].

-aberration [DT99]. -admissibility [HR99]. -Ancillarity [CH91]. -consistent [YD97]. -Dimensional [DS95b]. -efficiency [DW95]. -Estimates [CK95, Cha92, KT91]. -Estimation [PK90, Wel90, Gey94, HJ90, KS93, PK89, Wel89, KT96a, Kol97, Nis99]. -Estimators [CV92, Dav90, Lah92, LY94, Nie92, HS96a, KS97, LZ98b, MM99, RG97, Roc96]. -means [CAGM97, GEGM99]. -Mixing [PR92]. -Norms [Hor91]. -Optimal [Det90, Det93b, Hei94, PS93, Det96]. -Optimality [Det93a, DS93]. -optimum [RSW97]. -Processes [Wel90, ACG94, She94, Wel89]. -Quantile [YZ93]. -Quantiles [Arc95, Arc96]. -Robust [She95]. -sample [Rub96]. -spacing [Swa99]. -Statistic [Hel91]. -Statistics [Cha92, HJ93, Sha94, Stu94, Xia94, HS96b, Hös96]. -Values [Mar91a, Men94b]. -Variation [Dud94]. -Wasserstein [CAMRRdB99].

1 [Con99, Pit94]. 1994 [WWG^{+}95].

23 [Ano95]. 25 [Kus98].

99g [Kus98].

A. [DCLP90]. Aberration [CW91, Che92a, CH96, CM98, DT99, SCW97, TW96]. Absolute [FH94]. Absolutely [Sal98]. Abstract [NW97]. acceptable [DL96]. Accuracy [Arc95, Hal92a, HCR^{+}92, BB96, CL96b, Kol99]. Accurate [Loh90]. ACE [Buj90, Koy90]. ACE-Type [Koy90]. Achieving [RB90]. Acknowledgment [NT94]. Action [HL90b, Wij85, Wij93]. acyclic [AMP97]. adaptation [LMS97]. Adapting [WW94]. adaption [MGR98]. Adaptive [BX91, Bre91, BDHT91, Cai99, DKW97, EW95, Far92, Fri91a, Fri91b, Gas93, GH91b, GW91, HJ90, Hen97, Jin92, KP90b, KS96b, Low92a, LN99, MPZ98, MP92, O’S91, Owe91a, PV99, PK89, PK90, Ros93, Sch91, Spo96, Sto91b, Wu95b, CHY99, EL96a, EL96b, LS97, MvdG97a, NvS97, Spo98, Tsy98]. Adaptively [Fan93a]. Addendum [Sch95]. Additive [Cuz92, Dha91, DBS94a, DBS94b, Flu94, LY95b, FHM98, Hua99, OR97]. Additive-Multiplicative [LY95b]. Additivity [Bar93, EHSS95]. adjustment [Sta96]. Admissibility [AP95a, AP95b, BM92, GSM94, Lel93, Mee92, YG95, BS96, HR99, Maj96a]. Admissible [Eat92, KZ92, Maj96b]. Admit [BLEL94]. Affine [LR91]. Affinely [RT92b]. Against [HMM92, HKL92, NT94]. Alarm [APRY93, Yak95, GP97, Yak96a, Yak96b, Yak98]. Algebraic [DS98]. algebras [MN98]. Algorithm [Men94a, DLM99, Gou96, MLN99].

Algorithmic [GP93a]. Algorithms [HM94, MR96, DS98, MT96]. Alignment [WG97]. Allocation [GW92, Ros93, WH90]. Almost [GV91, Stu93, HL93]. Alternating [Buj90]. Alternative [AP95b, NT94, Str99]. Alternatives [CKS94a, GP95, MS83, MS92, CS98, TK97]. Analysis [Alb91, AA90, BGL95, CO90, Cox93, DM92, DL93, DL95, DBS94a, DBS94b, Flu94, FZ94, GJ90, HWG65, HWG95, HBT95, JJ90, JC90, Kan91, LY91b, LY94, LY95b, Mul92, OC95, Rob94b, WMC98, ZK90, BB98, Con99, DJ96, DN98, DT98, Gru96b, Koc96, LK99, LPS99, MT99, Sil96]. Analyze [KG92]. Ancillarity [Ber90, Bro90a, Bro90b, Cas90, CH91, Cop90, Efr90, FR90, Gle90, Hil90, Joh90, Lu90, Mor90, Ste90a, Swe92]. anisotropic [NvS97]. Ann [Kus98]. Ann. [Ano95]. Annealing [Eva91]. ANOCOVA [Sen96]. Anomalies [MS91, MS95]. ANOVA [Hua98, Luo98, WWG^{+}95]. Anscombe [Liu97]. Appears [Ber90, Bro90a, Bro90b, Cas90, Cop90, Efr90, FR90, Gle90, Hil90, Joh90, Lu90, Mor90, Ste90a]. Application [BL91, DB90, ET91, Fan93a, GK94, GJ90, PR92, Wij85, Wij93, vdG95, Düm98, HS96a, HS96b, Hua98, NvS97, SCW97, WWG^{+}95, Yan97]. Applications [ACG94, CD91, CT95, CH94a, DGZ95, DG91, Dat91b, GL91, HWG65, Hal91, HWG95, Kon92, KO94, LD93, She94, TT90, Wor95, Xia94, Yu93, And96, BL96b, CS96, DM98a, KK96, LW99, RT97, Swa99, WS97]. Approach [Cha92, Csi91, DF93, FR92, Kub94, LD95, MN92, Myk92, OAT91, Pol95, Ste95a, Vos91, ZK90, Bei96, Cai99, CF97, HK96, Hsi96, Hsi97, LMS97, Li96, MW99a, STZ98, Wan99]. approaches [AM97b]. approximately [SY98]. approximating [Wal97b]. Approximation [BS91a, BS91b, HI90, Hel91, Jon92, Kon92, Kon94, Low92b, Mar83, Mar86, Mar91b, Neu94, Por85, Por91, Yin90, DLM99, JT99, Neu98]. Approximations [JR94, VMT95, Csö96, HR98, Kol96, RS98]. Arbitrary [DGZ95]. arc [Wu97]. ARCH [DM98a]. Arcing [Bre98]. Area [DG91]. AREs [CL92]. Argument [BG90b]. ARIMA [KT96b]. Arising [BLRS94, Sma97]. ARL [Yak96b]. ARMA [Bai94, DCG99, MGKA95]. ARMAX [HG90]. Armed [Cha94, Sar91]. arms [BCZ^{+}97]. array [Loh96a, Muk99]. Arrays [Che95f, HPS92a, Owe94, PR92, HSS97b, MW95]. Aspects [Dav93, GM93, KM92, Lav92, Lav94]. Assigned [Efr95, Pru91]. Associate [CB91]. Association [BL92, BD93, Jun93, Mag93, CS96]. Assumptions [Mam91b, PR94, Bir97]. Asymmetric [Tak94]. Asymmetrical [HPS92a, MW95]. Asymptotic [AA90, And99, AW93, ACG94, Arc95, BH99, BRF97, BRR98, BF90, BL96a, BZ98, BM98, Cha96, CH94a, CH94b, Chi90, CO90, Dat91a, Dum91, ET91, EL92, FR98, GK94, GW98, GV91, Gir91, Gir98, GL92, GH94, GHL99, GZ93, Hal91, Hey92, HC92, Hua97, IL96, JP99, JC90, Kan91, KS97, Lah95, Lai94, Lar90, LY95a, MSY94, Mur95, Myk92, NT94, Nus96, OAT91, Par98, Phe96, Por85, Por91, PH96, Rit90a, Rit96, Ros93, RT92a, RS95, SW92b, Shi91, SM90, SV98, Ste90b, Str96, Swe92, Tak94, TT90, Wu90, Yaj91, Zam91, Zho92, BGY97, CHST99, CCR96, DM98b, Fal98, HH96, Jia96, Koc96, MLN99, RT97, ST96].

Asymptotical [MT95a]. Asymptotically [BN91, Che95b, CHS96, DW96a, FM93, GG99b, GL98, Huw96, IK91, Maj96b, Ryd94, Tak92, Wei95, Pol99]. Asymptotics [BDJ93, CG91, Dav90, Dav92, Gey94, GG96b, HRC94, HR95, KP90a, Kou96, Lop99, Nie92, PS92, Ric95, SWZ98, SY99, Sto91a, Wan96a, ZF96, vE92, Por97]. attack [WS97]. attempt [CAGM97]. author [Bre98]. Autocorrelation [Aba93, Deg90, HMM92, KL91, SS97]. autocorrelations [DM98a, Lee96]. Automatic [Neu95, CFM97]. Autoregression [Cha90a, KP90b, KO94, KS95, BRF97, HJ99, KN98]. Autoregressions [CT95, Kre90, TT90]. Autoregressive [BMM^{+}91, BP91, Cha93, CL91, DHY95, KS93, Mil92, Cha96, HJ99, KP97, KS96b, LL98, Lin98]. Average [AW93, BH90b, GP97, Nyc90, Tra93, CDS97, GW98, LL98, Yak98]. Averaged [Zam91]. averages [AM97a, HH96]. axes [BF98]. Axiomatic [Csi91].

Backfitting [Luo98, MLN99]. Bahadur [Arc96, Bos98, Cha91, Gre93, Hal91, HS96a, Hes90]. Bahadur-Type [Hes90]. Balance [BSV99]. Balanced [AC91, BHW93, CD91, CB91, KG91, SW92b]. band [CY96]. Bandit [BCZ^{+}97, Sar91]. Bandits [Cha94, Sam92]. Bands [SL94, LHMY96, XM95]. Bandwidth [Chi90, Chi91a, Chi91b, CM91, FM92, FG92, Hal92c, HV90, JMP91, Loa99, Pat93, Che97, HLT95, HLP95, LMS97, Neu95]. Bandwidths [vE92]. Bars [HM91]. Bartlett [BG90b, DHR91, JW96, Myk94, Myk99]. Bartlett-Correctable [DHR91]. barycenters [HM98]. Based [BL92, BH90b, CHP95, DL90, DG92, GZ93, HP95, HV90, Hjo90a, Hjo90b, HG90, Kan92, LvE93, Liu90, MY93, PR94, Roj92, SC91, Tyl94, Was90, XY95, CAMRRdB99, Det99, FR98, GKL98, GKL99, HJ99, Hua97, Kor96, LMS97, LR97, Peñ98, Qin99, RSW97]. Basic [Kon94]. basis [Don97]. Baye [WK90]. Bayes [Bei96, BR90, BRR91, BG90c, BM98, Das91, Dat91a, DB90, DF90, DL90, FSW98, GK94, Haf91, HK96, HW91, Hjo90b, Kar96, Liu96, MSY94, Phe90, Ruk94, SA96, WH91]. Bayesian [BB98, BBW94, BG90b, Bos94, Cho98, Con99, Cox93, DS91, DG91, DW93, Det96, DW96b, DF93, Dos94a, DHL97, EKS91, FS97, GG97, Gas95, GW95b, IR97, Kim99, LH92, Lo93a, Lo93b, SQ96, VW98, Yeh94]. Behavior [Dum91, Ioa94, Por85, Por91, SF94, vZ91, BM98, Jia96, Swa99]. Behaviour [Lah95]. Belief [Gol91, Was90]. Bell [CCL95]. Beneath [Lal99]. Bernoulli [Sam92]. Bernstein [Fre99]. Berry [HJ95, Seo90, Xia94]. Best [FM92, Hey92, YP92, Don97]. best-ortho-basis [Don97]. Beta [Hjo90b, Uhl94, WM97, DGJ97]. Beta-Stacy [WM97]. Between [CH95, HWG65, HWG95, RT97]. Bias [Hal92a, HS93, Low95, MY93, MZ93a, MZ93b, RC94, YZ93, Zam92, Adr98, CF99]. Bias-Robust [MY93, MZ93b, Adr98]. Bias-Variance [Low95]. Biased [BR91b, BNW92, EW95, Qin93]. Bibliography [DCLP90]. Bickel [EL96a, GH91a]. Binary [DF93, BL97]. Binomial [Woo99]. bioequivalence [BHM97]. birth [Phe96]. Bits [Per92]. Bivariate [Akr94, BCFS90, CC94, Mag93, Pru91, OR97, vdL96]. Blackwell [Pro91]. Blind [GG96a, Li95, GG99a]. Block [AC91, CM93, CB91, HKP98, Lah95, LD91, Tju91, Cai99, Lah99]. blocked [MU96]. Blocking [MW99a, CC99]. blocksize [UM97]. Blockwise [Buh94, NNR94, GK96]. Body [CK95]. Bonferroni [NW92]. Boolean [HP96]. Boosting [BFLS98]. Bootstrap [AG91, AG92, Arc95, BHW92, BH93a, BHW93, CA91, FK91, Gas95, Hal92b, Hal92a, HM91, Hel91, HJ93, Kon91, Lah95, LY95a, Lo93a, Mam93, MN92, NNR94, Sri94, BPR99, Büh98, CL96b, Cho98, DJ96, GW98, GK96, Lah99, RS98, Sin98, Str99]. Bootstrap-Confidence [FK91]. Bootstrapped [Buh94]. Bootstrapping [BMM^{+}91, FH92, Lah92, Myk92]. bootstraps [Jam97]. both [HLP95]. Bound [APRY93, GN90, Mar92, Xia94, KK96, Yak96a, Yak96b, ZBCL97]. Boundaries [KST95, MT95a]. Boundary [Oeh92, CFM97, HP96]. bounded [Sal98]. Boundedly [Mat93]. Bounds [BH90b, BL91, CL92, Dat92, DB90, Fin94, GK94, HJ95, HS93, Low93, RB90, RS98, Seo90, Ste90c, Tak94, Bro98, GW98, GMW98]. Box [CL92]. Bradley [CM93, SY99]. Branching [SBH91, Sri94, WW90]. Breakdown [DG92, LZ98b, LR91, Lop92, MM99, Sin98, Tyl94, Dav98]. Broad [Seo90]. Broadband [MS99a]. Brown [DS97, Wan99]. Brownian [SW94b]. Buckley [LY91b].

calculation [DW95]. Calibration [MK94, Gru96b, MZ96, MNS98a]. can [DG90]. Canonical [Buj90, Deg90, Pos92, DN98]. Capacities [Szk92, WK90, KW96]. capacity [CM98]. Capon [Ioa94]. Capon-Type [Ioa94]. Carlo [Asm90, CS97, KL95, LL97a, Owe94, Wal97b]. CART [Don97]. Case [Fos91, GL92, Lin94, Xia94, GG99b, STS96]. Case-Control [GL92]. Cases [Vos91, vE92, CH99]. Categorical [GH95a]. Cauchy [McC96]. causal [IR97, Kos96]. CDF [HK91]. cell [DE98, Hug96, Kan92]. censor [BM97]. Censored [BTZ93, Cha90b, Din93, FR94, GL91, GZ93, HZ94, Jan91, Kim93, LY91a, LY91b, LY91c, LY94, LvE93, Lo93a, Rit90c, Tsi90, Xia95, XY95, Yin92, Yin93a, Zho92, vdV94, CL96b, CLW99, Hsi96, Hsi97, Jam97, LZ98a, MR96, RG97, VA99, WG96, Yan97, ZL96, vdLM98]. Censoring [Akr94, EM90, FZ94, JK95, MBL91, VZ92, Yeh94, GG99b, Hua96, vdL96]. Censorship [CGH91, Neu93, SW93c, Stu95, Yan91, Csö96]. Centered [Ber93]. Central [CGH91, EW95, HT82, HT93, Stu95, Bis98, GEGM99, Loh96a, ZBCL97]. Certain [Haf91, Kha91, Pfa90, vdG93, DGJ97]. chain [ADS96, BS98, FS97, HR99]. Chaining [Eva91]. Chains [Bes94, CG94, Dos94b, Eat92, GW95a, Rob94a, Tie94a, Tie94b, BW99, LL97a]. Change [APRY93, BH92b, DHY95, Dum91, Fer95, FG93, GP94, Har94, HKL92, Loa92, Loa96a, Mul92, PS91, Rit90a, SV95, SW94b, TM93, Yak95, GP97, KPY99, Rai98, Yak96a, Yak96b, Yak97]. Change-Point [Dum91, Fer95, FG93, SV95, Yak95]. Change-Points [Mul92]. Changes [Hor93, CW99, CMT98, PM95]. Channel [DM92]. Channels [DMR90]. chaos [Lal99]. Characteristic [Wor95, HT96]. Characteristics [MZ94, NW92]. Characterization [LvE93, TW96, XY95, AMP97, BM97, CL96a, GH97b, JE97, MW99b]. chart [SS97]. Charts [BH90b]. checks [KS99a, Stu97, STZ98]. Chi [AA90, DB90, IL90, Kim93, LD93]. Chi-Square [AA90, IL90, Kim93, LD93]. Chi-Squared [DB90]. Choice [BS96, HV90, HLT95, HLP95, Neu95, Sil96, Spo98]. Choquet [KW96, WK90]. Chow [Liu97]. chromosomes [YS97]. circle [Wu97]. Circles [LS92b]. circular [Wu97]. Citation [DCLP90]. Class [AA90, CB91, Che95e, Com92, Det92, GV91, GN90, Hes90, JC90, Mas94, MG93, Jam97, Lan96, LK99, WG96]. Classes [Bos94, AMP97, DL97, NvS97, Tsy98]. Classical [Das91, Loa99]. Classification [HT98, KZ95, BSV99, LN96, Nob99]. classifier [Bre98]. Clinical [GL91, Sie93]. clone [YS97]. Closeness [GSM94]. clumps [HP96]. Clusters [Pol95, MNS98b]. Coarse [HR91]. Coarsening [JK95]. Coefficient [Aba93, BH92c, BM94, JV93, BFH96, FZ99, KS96b]. Coefficients [BP91, HMM92, NT94]. Coherent [BRR91, BR94, Bru91, DS99a, DHL97, KW96]. Cohort [BGL95]. Coin [AGW90, EW95]. Collapsibility [Fry90]. columns [MU96]. Combination [DS95b]. combinatorial [Loh96a]. Common [Geo91, Kne94a, Loh91b, ST92]. Compact [Dat91a, Det94, Ric95, Maj96b]. Comparing [HM93, Maj92, Sie93, Bet96]. Comparison [CC94, CM91, EL92, Gol91, Gre96, HM90, Mes91, ST92, SW93b, Ste90d, Ste94, BF98, DM98b, FR98, Gir98, Nak97]. Comparisons [Neu94, Lah99, MPB96, SY99]. complementary [TW96]. Complete [Che95e, Mat93, Mat96]. Completeness [MS93]. Complex [FF91, Sha94]. complexities [LN99]. complexity [DL97]. Component [GW92, Ros95]. Components [CD91, DBS94a, DBS94b, FH91, Flu94, KZ92, FHM98, Hug96, KS99b, Sal98, Sil96]. composite [IKL97, Peñ98]. Composition [Dud94, KDK90]. Compound [Dat91a, Mas93, Maj96b]. Computation [KMZ98, RS95, HK96]. computing [MR96]. concave [Bro98]. Concentration [Ko92, Pol99, Pol98]. Concentrations [Pol95]. Concepts [LS92b]. concerning [BNO96, CY96]. Concomitants [ND94]. Concurvities [DBS94a, DBS94b, Flu94]. Condition [Lar90, Pin94]. Conditional [AP93, Ber92, BBW94, BG90a, BHW92, FR92, FM93, GP93a, Jan91, Jia99, JR94, Jun93, Muk92, Muk93a, Neu93, Stu94, Swe92, AM98, BS98, DS98, HJ96b, Kol96, Kol99, MN98, NO96]. Conditionally [FS97, KM92, SW92a]. Conditions [RT92a, And99, Kni98, MLN99, MW98, VZ97]. Cone [CS96, HW94, MRS92, TK97]. Cones [Mas94, CL96a]. Confidence [BD95, BH93a, cICM92, CH93, Das91, DGZ95, EM99, FK91, FH91, Fin94, GC91, GC92, Hal92b, Hal92a, HJ95, HS95, Hey92, HL90c, HB91, HP94, Kon91, LY95a, MK94, Owe90, PR94, SWZ98, Sho90, SL94, Zha94, BD98, Düm98, GW98, Huw96, LHMY96, Low97, MZ96, MNS98a, Neu95, Qin99, TB97, Wan99, XM95, ZP96]. confounding [Li97]. conic [DCG99]. Conjecture [CM93, CY96, EL96a, Sar98]. conjectures [DGJ97]. Conjugate [BLEL94, HR99]. connection [Don97]. Consensus [GC93]. Conservative [LM94, MZ96]. Considerations [SM90]. Consistency [BL97, Cha93, CH94a, CCL95, Com92, Dat91b, DGKL94, DF90, HJ93, KL95, Koy90, LN96, MN92, Mur94, Ruk94, Sha93, TC85, TC90, WvdG96, vdG93, BSW99, CHY99, Che95d, GGR99, Hes96, Ish98, Jia97, Li96, MW98, PvZ96]. Consistent [Cha90b, DL90, GZ93, Ler92, Nyc90, Ryd94, Ste97, Stu94, TLF95, BR96, Det99, MR96, YD97]. Constancy [NT94]. Constant [BH90a, FM92]. constants [CS97]. Constrained [Fel91, Gey94, HW94, JM92, KT96a, MZ93b, BL96b, Wan96a]. Constraint [HB91, Yak96b]. Constraints [BN91, MRS92, DE98]. construct [ZP96]. Constructing [DGZ95]. Construction [HPS92a, Jac94, FSW98, HSS97b]. Constructions [HM98]. Contaminated [KM92]. Contamination [HS93]. Contents [Ano95]. Contingency [DS95b, RG91, LL97a]. Continuous [AG91, HK91, JK95, Zha91, Bos97, Kar96, Nis99, Sal98]. Contour [Pol95]. contoured [KS99c]. contours [HW97b]. Contrast [Sko90]. Contrasts [KBH93]. Contributions [Ruk90]. Control [BH90b, GL92, Maj92, Fin99, Maj96a, SS97]. Controlling [Ber92]. Converge [DG90, HP92]. Convergence [Ami96, Bai94, BP91, BG93, Bur91, CA91, Cha90b, Che91b, Che95c, DLM99, DL91a, DL91b, DL92, Fan91b, GJ95, GL91, HW97b, Jon92, Kol99, Koo93, KO94, Men94a, MP92, Ros95, Rus95, SW94a, WS95, ADS96, Bro98, CL97, CG95, Dre98, EL99, GG99a, GGS95, HKT97, Hes96, Kar96, Lin98, MT96, NY97, YB99]. Convex [Nie92, And96, DE98, Rie97]. convexity [Kol97]. Correctable [DHR91, JW96]. Correction [BS91b, Bel91, BG90b, CR94, Cuz90, DL95, DR98, Gho91, GKL99, HWG95, HZ92, HT93, Kus98, Mar86, Mar91b, MS92, MS95, PK90, Por91, Sch95, Ste99, TC90, Wel90, Wij93]. corrections [CFM97]. correctness [BPR99, GK96]. Correlated [MU91, Par98, UM97]. Correlation [BD93, TM93]. Correlations [KBH93]. Costs [WH90]. counterexample [CY96]. counterexamples [BNO96]. Counting [CH94b, LY95b, MU90, vdG95, Kim99]. coupling [HT98]. Covariance [BDJ93, CD91, HFH94, Kur93, LH92, Loh91a, LR91, Ste90c, YB94, KS99c, KK96, MPZ98]. Covariate [BD93, MNS95, ZK90, NLB98]. Covariates [CKW95, DS95a, Sar91, Zha91]. Coverage [Ber92, FK91, Hal92a, HJ95, Lo92]. Cox [BGY97, BGL95, CH91, CL92, Dab97, GL92, Hua99, MM97, O’S93, Sas93]. Cramér [EL92, IL90]. credible [Cho98]. Criteria [CH90a, EH92, MSY94, DS99a]. Criterion [DP94, FG94, Kor96]. Critical [Sri94]. Cross [Pat93, Ste90d, Zha93, vE92, Gir98, KLW97]. cross-match [KLW97]. Cross-Validation [Pat93, vE92, Gir98]. Crossing [HK91, KL91]. Cube [KP90a]. Cubic [LM90]. Cumulant [Che90a, VMT95]. Cumulative [Hjo90a, MZ94, SZ98, HO97]. Curie [CG91]. Curve [FH94, HMNT97, Kne94a, Mac95, MG93, HKP98, HT96]. Curves [BD93, FR92, HM90, KG92, SM90, DS96, GW98, GW99, HR98, PH96, WG97]. cusum [Bei96, SW93b, Rit90b].

Data [AB90, BGL95, Cha90b, CS94, CCL95, DM92, Din93, Dos94a, Ell91, FR94, GH95a, GZ93, HL93, HV90, HR91, Hjo90a, Hjo90b, Hsi91, IKL97, Jan91, JJ90, KL95, Kim93, KG92, Kne94a, LY91a, LY91b, LY91c, LY94, LD93, LD95, Liu90, LS92b, Lo93a, Rit90c, Stu93, Tho93, TC85, TC90, Tsi90, Xia95, Yin92, Yin93a, Yu93, Zho92, vdV94, BL97, CGWW98, CL96b, CLW99, DS99b, FH96, HLT95, HR98, HK96, Hsi96, Hsi97, Hug96, IL96, Jam97, LHMY96, LZ98a, LPS99, LN96, MPB96, MR96, Niu98, Nob96, PM95, RG97, Res97, SZ98, WG96, Wan96b, ZL96, vdLM98]. data-dependent [Nob96]. Data-Driven [CS94, HV90, IL96, LN96]. datasets [HW97b]. death [Phe96]. Decision [Mas93, Rit90b, Maj96b]. Decomposable [DL93, DL95]. Decomposition [BG90b, Mas94]. Deconvolution [Che90a, Fan91b, Fan93a, Kim98, Li95, SC91, GG96a, GG99a, JvEvZ98, PV99]. Deconvolution-Based [SC91]. Deficiency [Xia95]. Defined [Nie92]. Definite [BLRS94]. Degenerate [HJ93, She94]. Degree [Det92, Det95]. Degroot [GC93]. delay [Poo98]. Deletions [Neu94]. demixing [Hen97]. Densities [CGH91, HHM95, HS95, MP92, Roe92, Zha90, Zha95, Bir97, Bro98, Sal98, XM95]. Density [BS91a, BS91b, BM95, Chi91a, CM95a, Dat92, Dev92, Dev94, DJKP96, EL95, GQ93, Hal92a, Hal92c, Hal93, HP95, HV90, HI90, Hen90, HG95, Hor91, HR95, HZ94, Ioa94, Low92a, Mar92, PR92, Pol95, RC94, Sko90, TS92a, Was92, Wij85, Wij93, Wu95b, Yu93, Che95d, Che97, DL96, EL96a, ET96, EL99, GGR99, HLT95, HMNT97, HJ96a, KP96, Kim98, KC98, Lau96, Loa96b, LN96, Neu98, Nus96, PV99, Pol98, SW99, Tsy97]. Departure [BD95, Tan93, FH99]. Departures [BH90a]. Dependence [BCFS90, CM95a, HMM92, Rob95a, Rob95b, RT92a, ST92, CCR96, HKT97, Hos97, MS99a, RH97]. Dependencies [DBS94a, DBS94b, Flu94]. Dependent [CM91, CM95b, Dah95, EM90, HV90, Hsi91, Hsi93, KO94, Myk92, NL95, Sam92, Yu93, ZK90, BM97, FZS96, HLT95, HLP95, Kit97, Neu98, Nob96, Rob97]. Depth [DG92, Liu90, LS92b, BH99, HW97b, LPS99]. Derivative [EFM95, CDS97, EL96a, KS96a]. Derivatives [Dud94, Wu95b, Efr98]. descent [DS99b]. description [BS98]. descriptive [LPS99]. Design [AC91, EFM95, FF91, MSY94, MU91, DS99a, HMNT97, HT97, RSW97, TRYV96]. designed [ET96, GP97]. Designing [SZ93a]. Designs [BC92, CM93, CW91, Che92a, CB91, Che95e, Cos93, DS91, Det90, Det92, Det93b, Det94, DHS95, Det95, EW95, EKS91, HZ90, HZ92, Hei94, LD91, Maj92, PT91, PS93, Ros93, Tan93, Tju91, BS99, BSV99, CH96, CHY99, CC99, CM98, CMT98, DT99, Det96, DW96b, DR97, DR98, DH98, DHI99, HS96a, HSS96, Kus97, Kus98, Loh96a, Maj96a, MU96, Rit96, STS96, Ste95b, Ste99, SCW97, TW96, UM97, Wu97]. Detecting [APRY93, GP94, KPY99, SW93b, SW94b, Vos91, Yak96a]. Detection [PS91, SZ93a, SV95, Yak95, CW99, Poo98, Yak96b, Yak97]. Determinant [BDJ93, KS96a]. determination [YB99]. Deterministic [Dev94]. Deviation [FH94, Pfa90, Rub96]. Deviations [Din93, IL90, Kor96, Myk99]. Diabetic [WWG^{+}95]. Diaconis [HJ96b, Pro91]. diagnosis [Bro98]. Diagnostics [KS94, GW95b]. Diameter [Das91]. Dichotomous [MS93]. Difference [MT95b, Det99]. Differentiability [Sha93]. Differentiable [Van91]. Differential [CMS94, OAT91, DS97]. difficult [CH99]. Diffuse [De 91]. Diffusion [Lar90]. digraphs [AMP97]. Dilation [SW93a]. Dimension [CKW95, CLW99, Ell95, HP96, Roc96]. Dimensional [DS95b, Fan93a, HL93, Mam93, MU91, PS95, SF94, WS91, FHM98, Fre99, Li97, MPB96, Mam96, Str99, VW98]. dimensionality [NY97]. Dimensions [DGZ95, Pol99]. Diptych [Eat92]. Direct [FHM98, ZP96]. Direction [HS92]. Directional [CS98, LR92, LS92b, Fin99, HK96]. directions [BF98]. Dirichlet [CR94, Cha94, CR90, GH97b, GGR99]. discontinuities [HMT96, Spo98]. Discontinuous [MS99b, Qiu98, RS95]. Discounting [Cha94]. Discrete [Cho91, NW92, Zha95, Con99, GG96a, GG99a]. discrete-time [Con99]. Discriminant [HBT95]. Discriminants [LR93]. Discrimination [Det94, DR97, DR98, MT99]. Discussion [BX91, Ber90, Bes94, Bre91, BDHT91, Buj94, Cas90, CG94, Cop90, Dos94b, Efr90, Flu94, FR90, Gle90, GH91b, GW91, Has94, Hil90, Joh90, Lu90, Mor90, O’S91, Owe91a, Rob94a, Sch91, Ste90a, Sto91b, Bre98, LPS99]. disease [WS97]. dispersion [Adr98, DS99a]. Distance [BM94, CH95, Dha91, Lin94, Yan91, Yin92, BG97, CAMRRdB99, Wal97b]. Distances [CGH91]. Distributed [Zho91, GKL98, GKL99, MW98]. Distribution [Aba93, Akr94, AP93, AGW90, Bar91, BF90, BH93b, Cha93, CR90, CR94, Cuz92, DH93, ET91, Fel91, GP94, GH94, HM94, Kan91, Ko92, LY91a, Ler92, LvE93, ND94, Pit94, ST92, XY95, Yeh94, YsC91, YP92, Adr98, And99, AM98, CCR96, Eri96, FH96, FH99, GH97b, GP97, Hua97, Kol96, KS99c, Pol99, RT97, TK97, Woo99, Yak96a, Yak97]. distribution-free [FH96, Pol99]. Distributions [And93, AG91, BH92a, BH92c, Bes94, BRW91, BCFS90, BHW92, BB93, BTZ93, CG94, Chi90, DB90, DG90, Dos94b, FM93, FW95, GD92, Gas95, HKL92, IK91, KM90, Lav92, Lav94, Loh91b, Mas94, MSW92, Myk92, Ric95, Rit90a, Rob94a, RT92b, SF94, Sun94, TLF95, Tie94a, Tie94b, Uhl94, VZ92, Zha90, AM97b, BG97, BH99, BSW99, BM97, BFH96, BB96, CC97, CL96a, DS98, GGS95, Har96, HM98, Kni98, Kol99, KM97, LM98, LL98, Ste97, Xu97]. disturbance [CHST99]. Divergence [CMS94]. domain [DJ96]. Dominance [MRS92]. Domination [Geo91]. Donoho [GH97a]. Dose [SM90]. Dose-Response [SM90]. double [ZBCL97]. double-indexed [ZBCL97]. Doubly [Cha90b, EW95, GZ93, Sto91a, MR96, RG97, ZL96]. Down [MP90, FR98]. Downturns [SM90]. Drift [SW94b]. Driven [CS94, HV90, KL95, IL96, IKL97, LN96]. Drugs [HL90b]. Dual [Myk95a, Wal97b]. duality [DE98]. Dynamic [APRY93, GW92, SW94b, Yak96a, KMZ98, WG97]. each [BM97]. Eaton [HR99]. Echelon [Pos92]. ECM [Men94a]. edges [Don99]. Edgeworth [BR91a, BGvZ97, HJ95, Hel91, KM90, PvZ98]. Effect [Gho80, Gho91, Hal92a, MBL91, Mur94, TM93, NLB98]. effectiveness [BFLS98]. Effects [BL92, Dha91, ZK90, BL97, IR97, MU96, STS96]. Efficiencies [Fan93b]. Efficiency [CH94b, GW95a, Hal91, Lin94, LS92a, MZ93b, RS95, Ruk94, Ste90c, DW95, Li96, MOT97, Str96]. Efficiency-Constrained [MZ93b]. Efficient [BRW91, CS94, Che95b, Cuz92, Dah95, GP94, Hua96, Hua99, IK91, KST95, Lau96, Lop92, Rit90a, Sch93, Sch95, Van96, vdL96, vdLM98, GL98, HK96, Wei95]. Efron [HJ96b]. Eigenvalues [ET91]. Electrical [Tju91]. Elfving [Det93a]. Elicitation [GD92]. Ellipsoid [Dav92]. Ellipsoidal [RT92b]. Elliptical [TLF95]. elliptically [KS99c]. Emission [CH94a, JS90, O’S95]. Empirical [Bai94, BCFS90, Buh94, CH93, CH95, DG90, DHR91, GW95a, GP93b, Hal90, Hes90, Hsi96, Jin92, Kit97, Mam96, NNR94, Owe90, Owe91b, PvZ98, Qin93, QL94, Qin99, RC94, VZ92, WH91, YsC91, vZ91, Cab96, CC97, EM99, Fal98, GMW96, HK96, HH96, Hsi97, JW96, Kar96, LW99, Lin98, LN99, RS98, Wal97b]. Empiricals [KO94, Kou96]. Engineering [CH90a]. Entropy [Csi91, GG97, HO97, NY97]. entry [BGY97]. Envelopes [HS95, Was90]. Environment [LK93]. Epi [Hes96]. Epi-convergence [Hes96]. Epidemiological [WWG^{+}95]. Equality [Kur93]. Equations [CH90b, FR94, QL94, DS97, Kus97, Kus98, Li96]. equilibrium [LL97a]. equivalence [AMP97, BL96a, Nus96]. Equivalent [Mes91, PS95]. Equivariance [Mas93]. Equivariant [Kub94, LR91]. Erdos [AGW90]. ergodic [MYG96, Nob99]. Error [BTZ93, Cuz92, HP95, HM91, MW92, Mar92, Nyc90, SC91, ZBCL97, CF99, GHL99, GH95b]. Errors [BH93a, CKW95, CV92, CM91, CM95b, FT93, MU91, Muk93b, Sri92, Yaj91, Cha96, Fin99, HLP95, Huw96, KS97, LHC99, LL98, Rob97, RS98]. Errors-in-Variables [BH93a, CV92, Huw96, LHC99]. Esseen [HJ95, Seo90, Xia94]. Esseen-Type [Seo90]. Essential [Jun93]. Estimate [CH94a, Mes91, Nyc90, RT92a, KLW97, Stu96, ZF96]. Estimated [HB91, Ste93, Wan99]. Estimates [Bur91, CK95, Cha91, Cha92, Chi90, Cuz92, Dev94, DGKL94, DF90, DG92, FH92, HHM95, KT91, Lai94, MY93, Ryd94, SW94a, Wu95b, YZ93, Bos98, BM98, DS97, DL96, Dre98, DE98, LMS97, NY97, Str96]. Estimating [BC92, BH92c, CH90b, EFM95, FH99, FR94, HZ94, KP96, LY91a, LM94, Loh91a, Loh91b, Mam91a, Muk92, Niu98, Pol95, QL94, SdH99, Sri92, Ste90d, Tak92, Tsi90, WS97, Wor95, Wu95a, Zha90, Zha95, vdG90, CS97, Li96]. Estimation [AB90, Akr94, BM94, BR90, BG90a, BRW91, BR91b, BM95, Bir97, BHW92, BB93, BB96, BTZ93, Buj94, CD91, CH90a, CH90b, Cha92, Che91a, Che95b, Chi91a, Cho91, Cox91, CL91, CM95a, Dah95, DG91, Dat92, Dev92, Dha91, DS95a, Don94, Dos94a, Efr95, EL95, EL92, Fan91a, FH94, Far92, Fel91, FG93, FW95, FF91, GK94, Gas93, Gey94, GM92, GJ95, GP93b, GQ93, HJ90, Hal92a, HM94, HFH94, HV90, HI90, Has94, HS92, HS93, HY98a, HW91, Hen90, HG95, HT82, HT93, Hsi91, Hsi93, HG90, HCR^{+}92, HP94, IK91, Jin92, JS90, JM92, Joh94, Kne94a, KE95, Ko92, KST95, KS93, KM92, Lel93, Ler92, LHC99, Loh90, Low92b, Low95, Mac95, Mag93, MZ93a, MZ93b, Mar92, MN98, MG93, MGKA95, MP92, NF91, NL95, O’S93, OAT91, OC95]. Estimation [Pat93, Phe90, Pit94, PR92, PK89, PK90, Rit90a, Rit90c, RG91, Rob95a, Roe92, RC94, Sch93, SM94, Sch95, Spo98, SBH91, Sto94a, Sto94b, TS92a, Tra93, TS92b, WH91, WW90, Wel89, Wel90, WS91, WH90, Xia94, Yan91, YB94, Yin92, Yin93a, Yin93b, YsC91, Yu93, Zam92, vdG95, vdV94, Adr98, BZ96, BFH96, BS96, BL96b, BLZ97, CF99, Cai99, CM98, CF97, DW96a, DJKP96, DJ98, Don99, DHS97, DHL97, DKW97, EL96a, EL96b, Efr98, ET96, Eri96, FHM98, FZ99, GG99b, GGR99, GL98, GH95b, HLT95, HMNT97, HW97a, HKP98, HJ96a, HT96, Hua96, Hua98, Hua99, Ish96a, Ish96b, Jia96, Jia97, JT99, KT96a, Kim98, KT96b, Kol97, KP97, KC98, KS96b, KS99b, KS99c, Lau96, Lee96, LS97, Loa96a, Loa96b, LN96, MPZ98, MvdG97b, McC96, MC96, NvS97, Nis99]. estimation [Nob96, Nob99, Nus96, Phe96, Pol98, PH96, Rai98, Rie97, Sen96, Ste97, Sun96, Tsy97, Tsy98, Van96, Wan96b, Wei95, Woo99, Xu97, YD97, Zha97, vdL96, vdLM98, vdV96]. Estimations [Hsi97]. Estimator [BH93b, BDJ93, Cha93, Cha90b, CCL95, Dav92, DS95b, GV91, HP92, HF92, HRC94, HR95, Ioa94, Kan91, Kim92, LY91b, Low92a, Mee92, MBL91, Muk93a, Pru91, SS94, Stu93, Xia95, Yan92, YP92, YG95, Zho91, Zho92, And99, BRR98, CL97, EL99, GH97a, GT99, GMW96, GHL99, Hes96, JP99, KK96, PV99, ST96, WvdG96, Yan97]. Estimators [ACG94, BP91, BF90, BS91c, CHP95, CS93a, CS94, CV92, CG91, Com92, CO90, CGH91, Dat91a, Dav90, DH93, Dum91, GM92, GW95a, GZ93, Haf91, Hal92c, Hal93, HP95, Har94, HP92, Hjo90b, Hor91, KZ92, KP90b, Kub94, Lah92, LY94, LB93, LR91, Lop92, Nie92, Pfa90, Por85, Por91, RS95, Sas93, She94, Sko90, Ste95a, Tra93, TC85, TC90, WC93, Xia95, vdG93, BG97, BH99, BD98, Bos97, BL97, Cha96, CHY99, Che95d, Che97, Det99, Dre95, ES96, FSW98, Gou96, Grü96a, HKT97, HS96a, JvEvZ98, Kim99, Kni98, Kou96, KS97, LZ98b, LL98, Lop99, MM99, MR96, Neu98, PvZ96, RG97, Rie97, Roc96, SY98, Wan96a, WG96]. Ethical [WH90]. Euclidean [CK95, LK93, MN98]. Euler [NW92]. Even [Det92]. event [SdH99]. evolutionary [NvS97]. EWMA [SS97, SW93b]. Exact [FH91, Gas95, MW92, Mas94, MK94, Ric95, RS95, DM98b, KMZ98, TB97, Wu97]. Excess [Pol95]. Excessive [IL90]. exchangeable [FR98]. Exclusion [NW92, NW97]. Excursion [Wor95]. Existence [Low92a, BR96, HSS97b, MLN99, Min97, MW95]. Expansion [BR91a, HJ95, Hel91, BGvZ97, HH96, KS97]. Expansions [AW93, Hen90, Myk92, Myk95b, Sun94, PvZ98, RT97, SV98]. Expectation [Kha91]. expectations [Gus96]. Expected [SZ93b]. Experimental [DS91, PT91, DS99a]. Experiments [LD91, SZ93a, ST92, Ste94, BZ98, DR97, DR98, Gre96, IR97, Maj96b]. experts [JT99]. explanation [BFLS98]. Explanatory [DS95a]. Explicit [Ste95a]. Exploratory [Kan91, Tho93]. Exploring [Bes94, CG94, Dos94b, Rob94a, Str99, Tie94a, Tie94b]. Explosive [KP90b]. exponent [FH99]. Exponential [Asm90, BLEL94, BS91a, BS91b, BB93, Cho91, CKS94b, Com92, Dat91a, Din95, Ish98, JW96, LvE93, LM90, Mas94, Ste95a, Sun94, WH91, XY95, Yeh94, Zha95, vdG95, BB96, Cas96, ET96, Has99, Ish99, JT99, Kar96, Poo98, VW98, WWG^{+}95]. Exponentiality [HKL92]. Exponents [DL92, MM99]. extended [SHKT97]. Extension [Cha92, MW99b]. externally [FS97]. Extracting [Per92]. Extraction [Bel84, Bel91]. Extremal [DS96, Hsi93, Pin94]. Extreme [HMM92, CMT98, Dre95, Dre98, Fal95, HW97a].

Factor [AA90, Kan91, DL96]. Factorial [CW91, Che92a, Cos93, Jac94, LD91, BS99, CM98, CMT98, DT99]. Factorials [Gho80, Gho91, MW99a]. Factors [DB90, DL97]. failure [vdLM98]. False [APRY93, Yak95, GP97, Yak96a, Yak96b, Yak98]. Families [Asm90, BS91a, BS91b, Cho91, Com92, Dat91a, Din95, IK91, LM90, Mat93, Ste95a, WH91, Cas96, ET96, Has99, Ish99, VW98, WWG^{+}95]. Family [BLEL94, BB93, CKS94b, Sun94, Zha95, BB96, JT99, Kar96]. Fast [Gir91]. Fiducial [DW93]. Field [SZ93b, SW95, Ste90b, Wor95]. Fields [Com92, CW99, GMW96, Ste97, Ste95b, Ste95c, Ste99]. Filter [De 91]. Filtering [TA91]. Filtration [KL92]. Finite [BNW92, BH90b, Che95c, DH93, HS95, Nau92, SF94, Tyl94, Che98, GT99, Ish98, Sho96]. Finite-Dimensional [SF94]. Finite-Sample [HS95]. finite-variance [GT99]. First [BMM^{+}91, EFM95, KL91, MZ94, KPY99]. First-Order [BMM^{+}91, KL91, KPY99]. Fisher [EJ90, Ko92, LD93, LL97b, Tho96]. Fit [And93, CC94, DB90, EL92, EH92, GH91a, Hjo90a, KL95, Khm93, Kim93, Kon92, LD93, RG91, CAMRRdB99, Düm98, MM97, MNS98b, Peñ98, Pol99, Spo98, VW98]. Fits [HM93]. Fitters [Ell95]. Fitting [Dah97, Ell91, LD95, OR97, Rus95, Qiu98]. Fixed [BBW94, cICM92, TRYV96, CHY99, HSS97a, HJ96b]. Fixed-design [TRYV96]. Fleming [Bet96]. Flexible [Sto91a, NLB98, ST96]. flow [Phe96]. Following [CL92, Che95a]. Form [Haf91, Det99]. Forms [Pos92]. Formulae [HP95]. Fourier [DH98, Hen90, RSW97, Zha90]. Fractal [CHP95]. Fractional [CT95, CW91, Che92a, Cos93, BS99, CM98, DT99, KT96b, MW99a]. Frailty [Mur95, Par98]. Free [CM93, Cos93, DL91c, LD91, FH96, Pol99]. frequency [DJ96]. Frequentist [BR90, BBW94]. Fully [Bru91, Liu97]. Function [BR91a, Bar93, BH93b, Buj94, Cha90b, CH91, Deg90, DGKL94, DH93, EL92, FW95, GP93b, Has94, HL90a, LY91a, Mag93, Mam91a, Muk92, MZ95, Pit94, Ste90c, Sto94a, Sto94b, Tra93, TS92b, VMT95, Wan96b, WC93, YsC91, YP92, vdG90, BLZ97, HSS97a, Hua97, LK99, NvS97, Rie97, SZ98, Spo98, WG96]. function-indexed [LK99]. Functional [BH93a, Buj90, GM92, Low92b, Low95, SY98, Zha94, Bis98, BL96b, Cho98, Det99, EL96b, Hua98, Sil96]. Functionals [BRW91, BM95, DS95a, Fan91a, GV91, Sha93, Van91, Dav98, GG99b, KP96, Lau96]. Functions [BS91a, BS91b, BD93, BLRS94, CR90, CR94, GH95a, HFH94, HW94, LM90, LM94, Roj92, She95, Tak94, TA91, WH91, Was90, Yin90, Zam91, DM98b, Efr98, HMT96, Kol96, LHMY96, MGR98, Owe97, Pol98, Tsy98]. future [Tho96].

G [Con99, Pit94]. Gain [KL91]. gamma [Par98]. gamma-frailty [Par98]. Gaussian [Che90a, Che92b, Cox91, Csö96, GT99, KL92, KMZ98, Maj96b, Nis99, Nus96, Rob95a, RS95, SW95]. Gaussians [Ami96]. GCV [Gir91, Gir98]. General [BLEL94, DGZ95, FW95, HM94, JK95, JC90, KZ95, LY95b, Lo92, MN92, MBL91, PR92, QL94, vZ91, And99, BRR98, BGY97, BL97, CCR96, DHS97, Efr96, HS96a, Kni98, LK99, LL98, YD97]. Generalization [Det90, VZ97, WM97, Yan97]. Generalizations [BS91c]. Generalized [CH94b, Che95b, CV92, EM92, Has99, HJ93, LD93, MM97, SV95, Ste90d, Tho93, TC85, TC90, CGWW98, CHY99, Che95d, GEGM99, Hös96, Ish98, Jia99, Kol99, KK96, Lan96]. Generating [VMT95]. generation [HJ96b]. Genetic [Ste94, LL97a]. Geo [Con99]. Geo/G/1 [Con99]. geodesics [Sma97]. Geometric [DS95b, PS95, Vos91, MW99a, Woo99]. Geometrical [OAT91]. Geometrizing [DL91a, DL91b]. Geometry [CMS93, CMS94, DS93, CW99]. GI [Pit94]. GI/G/1 [Pit94]. Gibbs [Ami96, DHL97, GMW96, GMW98, Kol99, Ros95, Yu93]. Given [BLEL94, PS95, BR96, CG95, DS97]. Global [DS95a, Hal92c, GH97b, HR98]. Globally [HS93]. GM [SY98]. GM-estimators [SY98]. Good [Dev94, TB97, EL99]. Goodness [And93, CC94, EL92, EH92, GH91a, Hjo90a, KL95, Khm93, Kim93, Kon92, LD93, CAMRRdB99, Düm98, MM97, Peñ98, Pol99, VW98]. Goodness-of-Fit [CC94, EL92, EH92, KL95, Kim93, Kon92, LD93, Düm98, MM97, Peñ98, Pol99, VW98]. Graham [Pro91]. Granulometric [Wal97a]. graph [BS98]. Graphical [DL93, DL95, Fry90, GP93a, Tho93]. graphics [LPS99]. graphs [FS97, Sma97]. Greedy [Jon92]. Grenander [GHL99]. gridded [HR98]. Group [CE90]. Groups [Szk92]. guaranteed [KP97].

H. [DGJ97]. Haar [KP96]. Hadwiger [Wor95]. half [GG96b]. half-normal [GG96b]. Halfspace [DG92, Wal97b]. Hall [CY96]. Hannan [HG90]. Hastings [MT96]. Hazard [EJ90, Hjo90a, LY95b, NL95, Pat93, FGK97, Peñ98, ST96]. Hazards [BGL95, CH94b, Mur94, Hua96]. Heavy [Lah95, Muk93b, Res97, SM94, DM98a]. Heavy-Tail [Lah95]. Heavy-Tailed [Muk93b, SM94, DM98a]. Hellinger [Lin94, Yan91, Yin92, vdG93]. Hellinger-Consistency [vdG93]. Hellinger-Type [Yin92]. heterogeneous [WS97]. heteroscedastic [DW96b, KK96, LL98]. Heteroscedasticity [MZ95]. Heuristics [Bre96]. Hidden [Ryd94, BRR98]. Hierarchical [BR90, JT99, BS96, Liu96]. High [HL93, KDK90, Lop92, Mam93, Dav98, Li97, Mam96, Roc96, Str99]. high-dimensional [Li97, Mam96, Str99]. High-Scoring [KDK90]. Higher [Kol96, Pol99]. Higher-order [Kol96]. Highly [Lop92]. Hilbert [Jon92]. Histogram [Kan92, Nob96, Wu90, LN96]. History [Hjo90a, Hjo90b, LD93, LD95]. HIV [WS97]. Hodges [Pro91]. homoscedastic [Hsi97]. honest [Huw96]. Hotelling [CW99, JJ90, MP90, Pin94]. Hunt [Szk92]. Hyper [DL93, DL95]. hypercube [Loh96b]. Hyperrectangles [DLM90]. Hypotheses [Gol91, HW94, Kre90, LB95, MS91, Ruk94, IKL97, MS95]. Hypothesis [BBW94, BM92, BN91, DP94, Peñ98, Spo96].

I.D. [Xia94]. Ideas [HI90]. identically [MW98]. Identifiability [Ish96a]. Identification [MZ94, Pos92, DH98]. Identifying [Det95]. Identities [Myk94, NW92, Myk99, NW97, SCW97]. identity [DS99b]. IDMRL [HKL92]. Ignorability [HR91]. II [Por91, DL91a, Por85, XY95]. IID [Efr96]. III [DL91b]. ill [GG97]. ill-posed [GG97]. Illustrations [JJ90]. Image [O’S95, Koc96]. Images [Wor95]. Immigration [SBH91, Sri94, WW90]. Imperfect [HPS92b]. Implications [DLM90, FH94]. Importance [BHW93, Hal91, NW97]. Improper [BR94]. Improved [BH90b, GC91, HHM95, KS99b, Muk93a, Sho90, Sri92, NW97]. Improvement [BH93b, KS99c]. Improving [Kub94, Liu97, SS94]. Impurity [BPKN92]. imputations [Liu96]. Inadmissibility [CS93b, MP90, Vel92]. Incidental [Pfa93, Str96]. Inclusion [NW92, NW97]. inclusion-exclusion [NW97]. Inclusion-Exclusion-Bonferroni [NW92]. Incomplete [AC91, CB91, Dos94a, Hös96, Lar90, Mat93, TC85, TC90, LHMY96]. Incomplete-Block [CB91]. Incorporating [Mur94]. Incorrect [Ste90c, Ste90b, BM98]. Increasing [GC92]. Independence [AP93, GP93a, Jun93, MNS95, AM98, BS98, DN98, GH97b, MN98, NO96]. Independent [HL90b, PT91, Zho91, MW98, Neu98]. Indeterminate [Nau92]. Index [HHI93, Hsi91, Hsi93, BZ98, Dre95, Dre98, Fal95, HKT97, HSS97a, Wei95]. indexed [LK99, ZBCL97]. indexes [HSS97b]. indicators [vdLM98]. Indices [GW92]. individual [CBL99]. Inefficient [Loc91]. Inequalities [BG90c, BLRS94, LB95, NW92, She94, WS95, Yan92, vdG95, And96, AGSC98, NW97, Sar98]. Inequality [BL91, ET91, BL96b, Cai99, Nis99, SA96]. Inference [BRR91, BNW92, Bru91, CT95, Cox93, Csi91, DW93, DKR91, ES93, FR92, HK91, JJ90, LH92, MU90, Myk95b, Nai90, Ros93, Sto90, Swe92, BZ97, BRF97, FH96, GW95b, Hos97, Hug96, IR97, Jia99, KN98, LPS99, MM97, MYG96, MvdV97, Rob97, Wef96]. Inferences [BR94, Eat92, BR96]. Infinite [Hes90, MGKA95, PS95, WS91, Fre99, KT96b, VW98]. Infinite-Dimensional [PS95, Fre99, VW98]. Infinite-Variance [Hes90]. infinitely [BCZ^{+}97]. infinity [SY99]. Inflation [FG94]. Influence [Che90b, She95]. Influential [Vos91]. Information [AB90, BH98, BG90c, BL91, CH94b, DL90, EJ90, GMW98, Ish99, LY94, Maj92, RB90, SZ93a, YB99, YS97, HO97, KMZ98, LL97b, SA96, VA99]. Information-theoretic [YB99]. Informative [LB93]. inhomogeneous [LMS97]. Initial [HP92, PS91]. Innovation [Khm93, STZ98]. Innovations [MGKA95]. input [GG99a]. Inputs [Li95]. Insertions [Neu94]. instability [Bre96]. Integrable [Sun94]. Integral [BM95, EL96a, Lau96, Stu96]. Integrals [BC92, Kon92, Niu98, Owe97, Ste95c]. integrands [Hes96]. Integrated [HP95, Low93, MW92, Mar92, Wu95b, KM96]. Integration [GJ90, ES96]. interacting [WS97]. Interaction [Che91b, Fry90]. Interactions [DT98, Str90, Str99]. intercept [BFH96]. Interface [BR90]. interpoint [MPB96]. Interpolation [HT97, Mee92]. Interpretation [Lo92]. Interval [GC92, HP94, SF94, GG99b, Hua96, LZ98a, WG96]. interval-censored [WG96]. Interval-Valued [SF94]. Intervals [BD95, CH93, Det94, FK91, GC91, Hal92b, Hal92a, HJ95, Hey92, Kem91, Kon91, LY95a, Sho90, Low97, Neu95, Qin99]. Intervention [SZ93a]. Intrinsic [OC95]. Invalidity [Sri94]. Invariance [HWG65, HWG95, KE95, Was92, BNO96, DG96, NO96, Yak98]. Invariant [Che95e, GC91, LR92, RT92b, YsC91, YP92, Xu97]. Invariants [ES93, Wij85, Wij93]. Inverse [Csi91, Dud94, EL95, HC92, HL90c, JS90, Koo93, GG97, Hsi99, JvEvZ98, KC98, LM98, ZF96]. invited [Res97]. Involving [TS92b]. Ion [DMR90, DM92]. irregularly [HT97]. Isotonic [GJ95, JvEvZ98]. isotropic [Ste95b, Ste99]. Item [Jun93]. Iterated [LY95a]. Iterating [Per92]. Iterative [CS93a, GC93, Rus95].

Jackknife [BH93b, KG91, Sha93, Wu90, Stu96]. Jackknifing [Shi91]. Jacobian [DGJ97]. James [BS91c, LY91b, SS94]. Johnson [MW99b]. Joint [HL90b, Ste97]. Jordan [MN98]. Jump [KL92, WC93].

Kalman [De 91]. Kaplan [BG97, GV91, Pru91, Stu96, Yan92, Zho91]. Kendall [VMT95]. Kernel [BG90a, Chi91a, CGH91, Dev94, FH92, GH91a, HHM95, HR95, Mes91, NL95, RT92a, RC94, TS92a, WC93, Xia95, Che95d, DL96, DL97, HKP98, LMS97, ZF96]. Kernel-Type [WC93]. Kernels [Dat92, MG93]. Kiefer [Arc96]. Known [BRW91, BTZ93, Fel91]. Kolmogorov [MS92, Ano90, CC94, Cab96, CC97, Che90b, DCLP90, HI90, IL90, MS83, NY97, Ruk90, YP92]. Kullback [CMS94].

lack [MNS98b]. Lancaster [Str90]. landmark [PM95]. Laplace [KS96a]. Large [AA90, BR91b, Con99, Din93, Hes90, HL90b, KG91, LY91b, MBL91, Pfa90, PR94, Por85, Por91, Rob97, Sto90, TC85, TC90, VZ92, Yin93a, Kor96, Myk99, Rub96]. large-deviations [Kor96]. Large-Sample [MBL91, Sto90, Rob97]. Last [HF92]. latent [JE97]. Latin [Loh96b]. Lattice [AP93, Com92, KM90, Owe94, RSW97, AM98, MN98, Ste95b, Ste95c, Ste99]. Lattice-based [RSW97]. Law [AGW90, SW93c, HY98b]. Laws [DL93, DL95, CL97]. Learning [GW92]. Least [BP91, Buj90, Cha93, CS93a, Csi91, KK96, Lai94, O’S95, Pat93, Pos92, RW94, Wei92, YZ93, vE92, Gou96, Wan96a, WvdG96]. least-squares [Wan96a]. Lecture [Fre99, SHKT97, WWG^{+}95]. Left [LY91c, GKL98, GKL99]. left-spherically [GKL98, GKL99]. Left-Truncated [LY91c]. Leibler [CMS94]. Lemma [Jon92]. Length [BH90b, Yak95, BW99, GP97, Yak98]. Level [Cos93, Jac94, CMT98, Tsy97]. levels [HSS97b]. Life [Hjo90a, Hjo90b, LD93, LD95, CZ96]. lifetables [WMC98]. lifetime [BM97]. Like [NW92]. Likelihood [AP95b, BG90b, CD91, CH94a, CH93, CG91, CO90, Cox91, CL91, DHR91, EL95, Hal90, Hor93, HW94, KS94, Mac95, MS91, MRS92, Myk95a, O’S95, Owe90, Owe91b, Pfa90, Qin93, QL94, RS95, SW92a, SV95, Ste95a, Ste90d, Tho96, TC85, TC90, WS91, WS95, Yin93b, ZK90, vdG93, vdG95, vdV94, BRR98, BL97, CHY99, CM99, DE98, EL99, EM99, FGK97, Har98, Hes96, JP99, JT99, JW96, Kit97, Lan96, LHMY96, LL98, Loa96b, MvdG97b, MS95, MvdV97, MR96, Myk99, Phe96, Qin99, Rub96, Sta96, Van96, VZ97, Wef96, vdV96]. Limit [BH92a, CGH91, EW95, HS96b, HT82, HT93, Ste95a, Stu93, Stu95, Bis98, CL97, GEGM99, Hos97, Loh96a, Yan97, ZBCL97]. Limited [Kem91]. Limiting [Aba93, Cha93, Kni98, LL98, Swa99]. Limits [Nob99]. lineage [Hug96]. Linear [Ber90, BRW91, Bro90a, Bro90b, Cas90, CM93, CS94, Che95b, Che90a, Che92b, Cop90, Csi91, DS91, DG91, Dav90, Dav93, DHS95, Efr90, EKS91, FG92, Fan93b, FH91, FW95, FR90, FZ94, Gle90, Har94, HS93, Hes90, Hil90, HT82, HT93, HW94, Joh90, Kne94b, Kre90, KM92, Lah92, LD95, Li95, LZ98a, LB95, Loa92, Lu90, Mam93, Mee92, Mor90, NF91, Owe91b, PS92, PK89, PK90, Pos92, Rit90c, Seo90, Sho96, Ste90c, Ste90a, Ste90b, SL94, Tsi90, Zha94, ZL96, BZ97, CGWW98, CHY99, Che97, DM98c, GG96a, GG99a, GKL98, GKL99, HS96a, Hsi97, Hua99, Jia99, Kol99, KC98, Kus97, Kus98, Lan96, Lee96, LHC99, Mam96, MvdG97b, MC96, MM99, Mül98, SY98, SHKT97, Str99, TRYV96, Zha97, ZP96]. Linearity [HL93]. linearizable [Dav98]. Linearization [KG91]. Linearly [PT91]. Link [DL91c, WW94, HSS97a]. Link-Free [DL91c]. linkage [DS99b, Tho96]. Liu [LPS99]. Local [BM92, CGWW98, Cha91, CHST99, CMS94, Fal98, FG92, Fan93a, Fan93b, FGK97, GW95b, Gus96, Hal93, LD95, Loa96b, Nyc95, Por97, SWZ98, SZ93b, Tak94, Tra93, CW99, Che97, GH97b, OR97, Spo98]. Locally [Har96, HS93, HJ96a, MvdG97a, RW94, Ste95b, Ste99, DCG99, Dav98, MPZ98]. Location [BD95, Cha92, Dah95, DL90, DG92, KT91, Kim92, LR91, Lop92, MZ93b, Mat93, SW95, Tyl94, Xia94, BH99, Dav98, Eri96, Grü96a, Hsi96, Ish96a, KT96a, LZ98b, Lop99, Roc96, Xu97]. location-scale [Hsi96]. Log [KC98, Loa92, Rob95b, Sto90, Bro98, DE98, Lan96, MS99a, Str99]. log-concave [Bro98]. log-convex [DE98]. Log-density [KC98]. Log-Linear [Loa92, Lan96, Str99]. Log-Periodogram [Rob95b, MS99a]. Log-Spline [Sto90]. Logical [GP93a]. Logistic [BL92, cICM92]. Logspline [Sto91a]. Long [CT95, CM95a, CM95b, Dah95, KBH93, Rob94b, Rob95a, Rob95b, Yaj91, CP98, CHST99, GT99, HLP95, HH96, Hos97, KS97, MS99a, RH97, Wan96b].

Long-Memory [CT95, Rob94b, Yaj91, CP98, CHST99, HH96, KS97, Wan96b]. Long-Range [CM95a, CM95b, KBH93, HLP95, Hos97, MS99a, RH97]. longitudinal [SZ98]. Loss [FW95, Lel93, Tak94, Yin90, YP92]. Low [HL93, FHM98, HMNT97]. low-dimensional [FHM98]. Lower [DB90, GK94, HS93, Low93, Yak96b]. Lowner [CD91]. LS [HP92]. LSE [Yaj91]. Lukacs [CL96a]. Lynden [CCL95].

MAD [GH97a]. Main [Gho80, Gho91, MU96]. Manifold [Hen90]. Manifolds [CMS93, Ric95]. many [BCZ^{+}97, HMT96]. mapping [YS97]. Mardia [BH92a]. margin [BFLS98]. Marginal [BRW91, JR94, ST92, BM97]. Marginalization [Fry90]. margins [HJ96b]. Marker [NL95]. Markov [DL95, AMP97, ADS96, Bes94, BRR98, BW99, CG94, Com92, DL93, Dos94b, Eat92, GW95a, HR99, KG90, KL92, Kos96, LL97a, Phe90, Rob94a, Ryd94, Ste95a, Tie94a, Tie94b]. Martingale [Kon94, Myk92, Myk95b, MM97]. martingale-residual [MM97]. Martingales [Khm93, Myk94, vdG95, Nis99]. Mass [Pol95, Pru91]. match [KLW97]. Matching [AGW90, RT92b]. Mathematical [Ruk90]. Matrices [CD91, Che95e, Kur93, Loh91a, LR91]. Matrix [ET91, JV93, LH92, YB94, Adr98, DGJ97, KMZ98, KS99c, KK96]. Maxima [SZ93b]. Maximal [She94, Wij85, Wij93, BH99, Nis99, Swa99]. Maximin [MNS98b]. Maximum [CD91, CH94a, Che90a, CG91, CL91, Csi91, DE98, EL95, Hor93, KS94, Lan96, ND94, O’S95, Pfa90, RS95, Ste95a, Ste90d, TC85, TC90, WS91, Yin93b, vdG93, vdG95, vdV94, vdV96, BRR98, BL97, CHY99, CM98, EL99, GG97, GG96b, Har98, Hes96, HS96b, JP99, JT99, LL98, MR96, Van96]. maximum-likelihood [BRR98]. MCMC [Bro98]. Mean [APRY93, BH90a, BR90, CZ96, FH94, Fel91, Geo91, HP95, HW94, Joh94, Kim92, Loh91b, MW92, Mar92, Mil92, PS91, SW93b, Tak92, BF98, CF99, Cho98, CG95, KP97, RT97, TB97]. Means [BR91a, CR90, CR94, Owe94, WW90, BS96, CAGM97, GEGM99]. Measure [BH92a, BRW91, DG90, DS95a, Cab96, CCR96]. Measurement [SC91]. Measurements [HZ90, HZ92, Kus97, Kus98]. Measures [BD93, KZ95, PS95]. Measuring [Pol95]. Median [HRC94, Koc96, MW98]. Medians [ACG94, Sma97]. Medical [Wor95]. Meier [BG97, GV91, Pru91, Stu96, Yan92, Zho91]. Memorial [WWG^{+}95, SHKT97]. Memory [BN91, CT95, Rob94b, Yaj91, CP98, CHST99, GT99, HH96, KS97, Wan96b]. Method [DGZ95, DL91c, Hor93, HG90, Lo93b, SW92b, ADS96, Fal95, Jia97, Kle96]. Methods [BGL95, Buj90, HPS92b, KG91, Koy90, LY91c, Lin94, RT92b, Tho93, Zha90, BFLS98, CS97, GG97, HMT96, HT97, HKP98, Kit97, Lah99, Lan96, LS97, LN96, Pol99, She97]. metric [HO97]. Metrically [Kim92]. Metropolis [MT96]. Minimal [MS93, PR94, Ste95a]. Minimax [Adr98, AB90, BL91, DHS95, DLM90, DJ98, Fan93b, Fel91, GK94, HS93, HW91, JM92, Joh94, MT95a, Rai98, Tak94, Tan93, YZ93, Zha97, Bei96, Don99, FSW98, Li96, YB99]. Minimax-Bias [YZ93]. Minimaxity [YsC91, YP92, YG95, Kor96]. minimisation [GH95b]. Minimization [Nie92]. Minimizing [Das91, FH94]. Minimum [BM94, BPKN92, BDJ93, CW91, Che92a, Dav92, DT99, Dha91, Lin94, Sko90, Yan91, Yin92, CH96, CM98, SCW97, TW96]. Mises [Ko92, EL92, Fre99, IL90]. Mismodeling [BL92]. Missing [LY94, WW94, CGWW98, vdLM98]. misspecified [DW96a]. Mixed [Cos93, FH91, Jac94, BZ96, BSV99, Jia99]. Mixed-Level [Cos93]. Mixing [BF90, Ler92, PR92, Zha90, Zha95, Ste97, Woo99]. Mixture [Bos94, Che95c, Roe92, Hen97, Ish96a, Ish96b, LR97, Qin99, Van96]. Mixtures [Dat91b, Kem91, DCG99, GGR99, Ish98, Ish99, JT99, Woo99]. mixtures-of-experts [JT99]. ML [Pol98]. ML-density [Pol98]. MLES [WS95]. Möbius [McC96]. Modal [Kem91]. Mode [CH99, GH95b]. Model [BGL95, CKW95, Cha93, CE90, cICM92, Che91a, CS94, CG91, Dav90, DHY95, Dha91, EM90, FW95, GK94, GL92, KE95, Loa92, Low93, MU90, MZ95, Mur94, Mur95, NL95, O’S93, Rit90c, RG91, STZ98, VZ92, Wu95a, Yaj91, Yan91, Zha93, BZ96, BR96, BZ97, BGY97, Bre96, DCG99, DHS97, HP96, HY98a, Hsi96, Hua96, Hua99, Ish96b, KS96b, KS99a, KK96, LHC99, LN99, NLB98, Nis99, OR97, Par98, SY99, Stu97, Sun96, vdL96]. modeling [CP98, Kle96, Res97, SHKT97]. Modelling [Lav92, Lav94, FH99, HSS97a]. Models [AA90, AP93, Bai94, BLEL94, BNB93, BH92b, BL92, BM94, BR91b, BF90, BH93a, BTZ93, CD91, CH94b, Che91b, Che95b, Che95c, CS93b, Cuz92, Dah95, DS91, DG91, DL93, DL95, DHS95, EKS91, ES93, FH91, Fry90, FZ94, GP93a, HHI93, Hjo90a, Hjo90b, HPS92b, HG90, Jun93, KG90, KS93, KM92, Lai94, LD93, LY95b, Mam93, Mas94, MS93, MGKA95, NF91, Owe91b, PK89, PK90, RB90, Ros95, Ryd94, SF94, Sch93, Sch95, SW92a, She95, Sri92, SC91, Sto90, Sto91a, Tan93, Zha95, BB98, BRR98, BSV99, Bis98, BS98, BM98, BL97, CGWW98, Cha96, Che98, CHY99, CHS96, CHST99, CS98, Con99, DW96a, Dah97, DW95, Det96, DW96b, DM98c, DHI99, DKW97, DE98, FHM98, FZ99, HJ99, HSS96, Hen97, Hos97, Hsi97, Hua98, Hug96, Huw96, Ish96a, JP99]. models [Jia99, JT99, JE97, KMZ98, Kol99, Kos96, Lan96, LW99, LR97, Lin98, Mam96, MvdG97b, MM97, MN98, Mat96, MM99, Mül98, Peñ98, SZ98, SQ96, Str96, Str99, VA99, Van96, Wef96, YD97, ZP96]. modes [Min97]. modification [GH97a]. Modified [LY91b, MS83, MS92, Ste90d, CC97, Swa99]. Modulation [BD98]. Molecular [KDK90]. Moment [Che95e, LR97]. Moment-based [LR97]. Moments [AW93, BP91, CW91, Gas95]. Monotone [FR94, HZ94, Jun93, KL91, KST95, Mam91a, Muk93a, CCR96, JE97, SW99]. Monotonicity [HW94, Kha91]. Monte [Asm90, CS97, KL95, LL97a, Owe94, Wal97b]. mortality [WMC98]. most [CHS96]. Motion [CK95, SW94b]. Moving [Lah95, AM97a, HH96, LL98]. moving-average [LL98]. MR [Kus98]. multi [HP96]. multi-type [HP96]. Multidimensional [BLRS94, Sma97, Cho98, LL97a, MPB96]. multifactor [DR97, DR98]. Multifold [Zha93]. Multilevel [Li95]. Multinomial [DB90, DF90]. Multiple [Ber90, Bro90a, Bro90b, Cas90, Cop90, DM92, Efr90, FG94, FR90, Gle90, Hil90, Joh90, Lah92, Lu90, Mor90, Nak97, Ruk94, Ste90a, EM99, FR98, Fin99, Hua98, MNS98a]. Multiple-comparison [Nak97]. multiple-use [MNS98a]. Multiplicative [LY95b, VZ92]. Multivariate [AP93, ACG94, Bar91, BH92a, BX91, BR90, BG93, Bre91, BDHT91, Buj94, CD91, Cha92, Cho91, Deg90, Fri91a, Fri91b, Gas95, Geo91, GKL98, GKL99, GH91b, GM93, GW91, Has94, Hor91, KT91, LPS99, Loh91b, Loh90, LR91, Lop92, MK94, O’S91, Owe91a, RW94, Sch91, Sto91b, Sto94a, Sto94b, TT90, Tyl94, Uhl94, Adr98, AM98, AM97b, BH99, Bro98, CC97, CS98, DGJ97, GEGM99, Grü96a, HW97b, HP99, KT96a, KM97, KS99b, Lop99, MZ96, MNS98a, MOT97, Roc96, Sun96, TB97]. Mutual [HO97].

N [DCLP90, Kim98]. Natural [LM90, WH91, Cas96]. near [BRF97, MNS98b]. near-replicate [MNS98b]. Nearest [Akr94, BG90a, DGKL94, Hsi99, Muk93b, GMW96, SV98]. Nearest-Neighbor [BG90a, GMW96]. Nearly [Cox91, CL91, Don99, MW95]. Necessary [MW98]. Negative [Pru91, Tak92]. Neighbor [Akr94, BG90a, DGKL94, Muk93b, GMW96, Hsi99, SV98]. Neighborhoods [Nai90]. Nested [GL92, MU96]. net [Owe97]. Network [Jon92]. Networks [Tju91]. Neumann [Per92]. Neural [Jon92]. neutral [MW99b]. Neyman [WWG^{+}95, IL96]. No [DG90, Ano95, Kus98]. Noise [KL92, Low92b, Low93, RS95, SM94, BL96a, Lal99, Nis99, Nus96, Zha97]. noisy [STS96]. Non [Che90a, Che92b, Low92a, RB90, Xia94, GT99]. Non-Existence [Low92a]. Non-Gaussian [Che90a, Che92b, GT99]. Non-I.D. [Xia94]. Nonasymptotic [Dat92, DL97]. Noncausal [Gas93]. noncompliance [IR97]. nonequispaced [BC98]. nonequivalence [BZ98]. Nonexistence [LB93]. Nonidentically [Zho91]. noninformative [DG96]. Noniterative [Kan91]. Nonlinear [DN98, HP95, KE95, Lai94, Li97, MU90, MZ94, NF91, Sal98, Yin90, Det96, HT97, Kou96, Wan96a]. nonnegative [KS99b]. Nonnormality [KS94]. Nonparametric [BF98, BNW92, BF90, BL91, CA91, CH90a, Cha91, CM99, Chi91b, Cox93, CM95b, DM98b, DF93, DS95a, DHS97, DH93, Dos94a, Dum91, Efr95, EL92, Fan91b, FT93, Fer95, GM92, GP94, GP93b, HJ90, Hal92b, HFH94, HM91, HM93, HW91, Hen90, Hjo90b, HG95, HPS92b, HT96, JC90, Kim99, Kne94a, Koo93, KS99a, LD95, LHMY96, Liu96, Low92b, Mam91b, MG93, Mil92, Min97, MSY94, MYG96, Mul92, NL95, O’S93, Pfa90, Pit94, SM90, Stu97, Tra93, TS92b, YD97, Zha91, ZK90, vdG93, BSW99, BFH96, Bos97, BL96a, BL96b, BLZ97, BZ98, DHL97, Düm98, Efr96, EL99, EHSS95, Gir98, Gru96b, HLP95, HJ96a, JvEvZ98, Koc96, Kor96, KN98, LS97, LK99, Loa96a, Low97, Neu95, PV99, PH96, Rob97, Tsy97, WvdG96]. nonparametrically [GW99]. nonquadratic [KP96]. nonrecursive [Kos96]. nonregular [DT99]. Nonsmooth [vE92]. Nonstandard [Fer95, VZ97]. Nonstationary [Bel84, Bel91, Cox91, CL91, HG90, Li95, TT90, Büh98, Dah97, Lin98]. nonstochastic [HS96a]. nonzero [STS96]. Norm [Yu93, Sil96, Tsy98]. Normal [AP93, APRY93, BR90, CS93b, CM95b, DS91, Fel91, GC91, Hor93, HW94, IK91, Joh94, Loh91b, PS91, Por85, Por91, Roe92, Ryd94, ST92, Sho90, Tak92, AM98, BS96, GG96b, Hes96, Sun96, TB97]. Normality [ACG94, BD95, RT92a, Zho92, BRR98, CHST99, Fal98, GHL99, JP99]. Normalized [Lah95]. Normalizing [Kon91, CS97]. Norms [Dud94, Hor91]. Note [AP95a, Det93b, Dev92, Gre93, HKT97, KG91, Pro91, Ste94, XY95, Yak95, vZ91, Bei96, Det96, GH97a, Yak97]. Notion [Liu90]. Novel [LK93]. NPMLE [Hua97, vdL96]. Nuisance [AP95b, Pfa93, Str96]. Number [HF92, Kem91, SZ93b, Wor95, SY99]. numbers [CMT98].

O’Brien [Bet96]. Observation [SM94]. Observations [Hor93, HZ94, Lar90, NNR94, Efr96, GK96, Neu98, Ste95c, UM97]. Observed [KL92, LL97b]. Obtained [KZ92]. Odd [Det92]. Odds [Wu95a]. old [WMC98]. oldest [WMC98]. oldest-old [WMC98]. Olkin [CL96a]. Omnibus [BH90a, MNS95]. One [Alb91, Fan93a, Fin94, Gho80, Gho91, Kur93, LB95, PS95, Sar91, CS98, Kar96, MPB96, Muk99, SY98]. One-Armed [Sar91]. One-Dimensional [Fan93a, MPB96]. one-parameter [Kar96]. One-Sample [Alb91]. One-Sided [Fin94, Kur93, LB95, CS98]. one-step [SY98]. Only [AP95b]. operating [HT96]. Operators [Dud94, FS97]. Optimal [BN91, Bur91, CH90a, CW91, Che95c, Det90, Det92, Det93b, Det95, DW96b, DR97, DR98, DH98, DHI99, DL92, Don94, Dre98, EL99, EKS91, FZS96, FM93, Fan91b, GH91a, GM92, HJ99, HHI93, HSS96, HZ90, HZ92, Hei94, Kan92, Kar96, Koo93, Kus97, Kus98, LS97, LMS97, Maj92, MU96, PT91, PS93, She95, Wu97, CC99, DW96a, DS99a, Det96, EL96b, GG99b, Maj96b, UM97, Wef96, Yak97]. Optimality [CB91, Dat91a, Det93a, DS93, Gir91, Rit90b, Ste90b, IL96, KPY99, Kus97, Kus98, LL97b, Maj96a, Muk99, Rit96]. Optimization [She94]. Optimum [EFM95, KM92, Mül98, RSW97]. oracle [Cai99]. Orbit [CE90]. Order [Alb91, BMM^{+}91, BP91, CD91, CS93b, CKS94a, DHY95, Dud94, EH92, GK94, GSM94, HP94, KL91, Mar83, Mar86, Mar91b, MZ94, Myk95b, ND94, RG91, Ste90b, Ste93, BPR99, CS96, DCG99, DH98, GK96, Kol96, KPY99, LL97b, Mat96, Pol98, PH96]. Order-Restricted [RG91, CS96]. Ordered [GP95, Kne94b, RG91, Sar98, SSK95]. Ordering [DKR91, LS92b, Roj92, CS96]. Orderings [BCFS90, LR92]. Orders [BCFS90, Cos93, Jac94]. Ordinal [DS95b]. ortho [Don97]. Orthogeodesic [BNB93]. Orthogonal [Che95f, HPS92a, Owe94, Zam92, HSS97b, Loh96a, MW95, Muk99]. orthogonality [BSV99]. Orthogonalization [Grü96a]. orthomedian [Grü96a]. oscillation [LR97]. other [BM97]. Outliers [Dha91, DT98]. Outlyingness [DG92]. Outperforming [GMW96]. overtraining [Kle96]. overtraining-resistant [Kle96]. Overweight [Loc91].

paired [SY99]. pairwise [HT98]. Panel [GH95a]. paper [Res97]. Paradox [Ber90, Bro90a, Bro90b, Cas90, Cop90, Efr90, FR90, Gle90, Hil90, Joh90, Lu90, Mor90, Ste90a, Wan99]. Parameter [AP95b, BB93, DHY95, HT82, HT93, Jin92, Ko92, KT96b, Lah92, MGKA95, Ryd94, Sri92, Ste93, Sun94, TM93, WS91, Wu95a, Yan91, BB96, GH97b, Kar96, McC96]. Parameters [BD95, cICM92, Hey92, Hor93, KS93, NF91, Pfa93, Por85, Por91, Ste90d, Tsi90, Vel92, Xia94, Yin93b, BZ96, Fre99, Ish96a, SY99, Str96, Xu97]. Parametric [Bos97, HM93, HG95, SW92a, CHS96, HJ96a, Mat96]. parametrization [DCG99]. Pareto [FH99]. Part [SS94]. Partial [BCFS90, CH91, Deg90, DL90, Gou96, ZK90, BZ97, FGK97, Gir98, MvdG97b]. Partially [CS94, CB91, vdV94, DW96b, GP97, LHC99, SSK95]. Partition [BH92b]. Partitions [BPKN92, Nob96]. partly [Hua99]. partner [HR99]. parts [ES96]. pattern [Kle96]. Peak [MP92]. Peaks [Wor95]. Pearson [LD93]. penalization [She97]. Penalized [CO90, HBT95, Mac95, MvdG97b, ZK90]. penalty [Poo98]. Performance [CH90a, HM94, HMT96, HR98, Koc96]. Period [HZ90, HZ92]. Periodic [AM97a]. Periodogram [CHP95, Rob95b, KM96, MS99a]. Periodogram-Based [CHP95]. Permutation [Che95e, Tho93, Sho96, ZBCL97]. Permutation-Invariant [Che95e]. Permutational [HMM92]. Permutations [BCFS90]. perturbations [Ami96]. Phylogenetic [ES93]. Pickands [Dre95]. Piecewise [Rie97, TK97]. pilot [Rie97]. Pitman [GSM94, HMM92]. Plan [SW94b]. Plane [Ell95, DS96, PM95]. Plane-Fitters [Ell95]. Planes [Ell91]. Plans [Gho80, Gho91, Yeh94, MU96, Muk99]. plot [BS99]. plots [EM99, GG96b]. Plug [GM92, Hal93, CF97, Loa99]. Plug-in [GM92, Hal93, CF97, Loa99]. Plus [Gho80, Gho91, Muk99]. Point [BH92b, CH90b, CMS93, CMS94, Dum91, Fer95, FG93, Har94, HK91, HKL92, Loa92, Lop92, Rit90a, SV95, Yak95, BG97, Loa96a]. Points [LR91, Mul92, PT91, TLF95, Tyl94, WC93, MM99, Rai98]. Pointwise [DL92, Tsy98, LS97]. Poisson [BPR99, FR92, Hen97, JM92, Loa92, Neu94]. Policy [Yak95, Yak96a]. Pólya [Ish98, Lav92, Lav94, MSW92, WM97]. Pólya-urn [WM97]. Polyhedra [Nai90]. Polynomial [Buj94, Det90, Det93b, Det94, Det95, Has94, Hei94, PS93, Sto94a, Sto94b, SHKT97, Che98, DW95, DHI99, OR97, Spo98]. Polynomials [Det92]. Polytopes [Tho93]. Pooling [GC93, FS97]. Population [BNW92, DH93, Mee92, Che98, DCG99]. Populations [CC94, DKR91, WS97]. posed [GG97]. Poset [Ros91]. Positive [BLRS94, SS94, Sal98]. Positive-Definite [BLRS94]. Positive-Part [SS94]. Positron [CH94a, JS90, O’S95]. possess [Yak98]. Possible [FM92]. Possibly [Hes90, BM97, BM98]. Posterior [Bes94, CG94, CG95, Dat91b, Dos94b, GGR99, KS94, Men94b, Nyc90, Rob94a, Sun94, Tie94a, Tie94b, BSW99, GGS95, Gus96, Ish98]. postulate [MW99b]. pot [Fal95]. pot-method [Fal95]. Potential [LM94]. Power [DS95a, GH91a, HL90a, HW94, SM90]. Powerful [LB95, CHS96]. pre [GP97]. pre-change [GP97]. Precision [AP95a]. Predicting [Ste95c, MGR98]. Prediction [Ber92, Ber93, DG91, DW93, Fos91, GH95a, CBL99]. Predictions [Ste90c, Ste90b]. Predictive [Men94b, Wei92]. predictors [Bos97]. Preface [Ano90]. preferences [SSK95]. Preferred [CMS93, CMS94]. prescribed [BB96]. Presence [AB90, SM94, SS97]. Present [AP95b]. Principal [DBS94a, DBS94b, Flu94, TLF95, DS96, Sal98, Sil96]. Principle [LY94, NY97]. Principles [Wei92]. Prior [GD92, Maj92, SZ93a, Was90, YB94, Har96, Har98, MW99b]. Priority [NT94]. Priors [BLEL94, BR94, Bos94, CE90, Was92, BS96, DG96]. Probabilistic [Pin94]. Probabilities [DF90, FK91, IL90, JR94, Nau92, Pfa90, SW93a, WK92, DE98, HW97a, Kor96, RS98, STS96]. Probability [APRY93, Ber92, Ber93, BRW91, DG90, ES93, Hal92a, HHM95, Hen90, KZ95, Low92a, PS95, Ric95, SF94, VMT95, WS95, HY98a, Sar98, SdH99, SQ96, Yak96a, Yak96b]. Probability-Centered [Ber93]. Problem [Alb91, BLRS94, Fan93a, Geo91, GJ95, Khm93, Lo92, MP90, Mas93, MBL91, Neu93, PR92, WH90, BHM97, ET98, ES96, GL98, Zha97]. Problems [BH92b, BM92, Csi91, Das91, EL95, Fan91b, Fer95, GD92, Har94, HL90c, JS90, Koo93, LB93, Low92b, Low95, Mac95, NW92, Pin94, Qin93, Sar91, Sha94, BSW99, BCZ^{+}97, GG97, KC98, Yak98]. Procedure [BTZ93, KL92, Mar83, Mar86, Mar91b, Per92, Rit90b, Rus95, Sri92, SW94b, Tak92, Bei96]. Procedures [DL90, MP90, SW93b, Yin90, FR98, Fin99, Nak97]. Process [BP91, Buh94, CR90, CR94, FR92, KL92, Lar90, Loa92, MU90, Phe90, Ste90d, WW90, HH96, Hsi96, Hsi97, KP97, Mam96, STZ98]. Processes [ACG94, Bai94, BMM^{+}91, BC92, BF90, CT95, CH94b, Che90a, Che92b, Cox91, CH95, EM92, Gas93, Hes90, Hey92, Hjo90b, HT82, HT93, KG90, KS95, LY95b, Mil92, NNR94, PS92, Pos92, She94, SBH91, Sri94, Ste95a, Swe92, TT90, Wel89, Wel90, vZ91, vdG95, BG97, Bos97, CC97, CP98, CZ96, DCG99, Dah97, DM98a, Fal98, Kim99, Kit97, KM96, Lee96, LW99, Lin98, MPZ98, MM97, Niu98, Nob99, WM97, HT93]. Product [BH92b, GJ90, Stu93, CL97, Yan97]. Product-Integration [GJ90]. Product-Limit [Stu93, CL97, Yan97]. Products [Buj94, Has94, Sto94a, Sto94b, SHKT97]. Profile [SW92a, Sta96]. Projected [CKS94a, DG92]. Projection [CH91, Che91a, Che95f, Hua98, Jon92, Tyl94, MLN99]. Projection-Pursuit [Che91a]. Projections [HL93, MY93]. projective [MW99a]. Proof [DGJ97, DS97, Sar98]. Proper [Wij85, Wij93]. Properties [CHP95, Che95f, DMR90, DG92, GP93a, GZ93, Hal92c, HL90b, HW94, JC90, KG91, Lai94, MBL91, SW92b, TT90, Was92, Wu90, Yaj91, Zho91, Ami96, Cha96, DS96, Hua97, KLW97, Kos96, LZ98b, LR97, MLN99, Roc96, SS97, ST96]. Property [CH91, HH94, YZ93, Bei96]. prophet [AGSC98]. Proportional [BGL95, CH94b, Mur94, Rus95, Wu95a, BB96, Hua96]. proportions [Qin99].

Pseudo [Hal90]. Pseudo-Likelihood [Hal90]. Pure [Roj92]. Pure-Tail [Roj92]. Pursuit [Che91a, Jon92].

Quadratic [Fan91a, LM98, WH91, Cas96, EL96b]. Qualitative [Mam91b]. Quality [CH90a]. Quantal [MS93]. Quantile [BG90a, CH95, EM92, FF91, Muk93a, Roj92, Shi91, Xia95, YZ93, CDS97, Che95d, EM99, LHMY96, Por97]. Quantiles [Arc95, Cha91, CH93, FK91, GJ92, Hes90, KS95, Muk92, SW92b, Arc96, Kol97, Sin98, ZP96]. quantizers [CAGM97]. Quasi [Wef96, CHY99, CM99, MvdG97b]. Quasi-likelihood [Wef96, CHY99, CM99, MvdG97b]. Queue [Pit94]. queueing [Con99]. Queues [Asm90]. Quickest [Poo98].

Random [Akr94, Asm90, BH92c, BM94, Com92, CGH91, ET91, JK95, KG90, Lah95, Liu90, LR93, MSW92, Mur94, NT94, Neu93, Per92, Pfa93, PR92, SZ93b, SW95, Ste90b, SW93c, Stu95, VZ92, Wor95, Yan91, Zam91, Zho91, AM97a, BFH96, BL97, Csö96, GG99a, GMW96, HY98a, HY98b, HM98, KS96b, MW98, MGR98, Sal98, Sar98, Ste97, Ste95b, Ste95c, Ste99, vdLM98]. Random-Multiplicative [VZ92]. Randomized [Gir91, Owe94, Gir98, IR97, Loh96a]. Randomly [AW93, Jan91, Kim93, KO94]. Range [CM95a, CM95b, Dah95, HL90a, HH94, KBH93, Rob95a, Rob95b, Seo90, HLP95, Hos97, MS99a, RH97]. Rank [Alb91, Cuz88, Cuz90, GL91, GJ92, Jan91, KS95, LY91c, MN92, Neu93, Seo90, Tsi90, Yin93a, And99, HJ99, MOT97, Sen96, Sho96]. Rank-Scores [KS95]. Ranked [Tho93, Hua97]. Rankings [AC91, VMT95]. rare [SdH99]. Rate [BG93, CA91, Che95c, EJ90, HP92, KL91, Men94a, Pat93, SW94a, CL97]. Rates [Bur91, Che91b, DL91a, DL91b, DL92, Fan91b, GJ95, Hjo90a, Jon92, Koo93, MT96, NY97, Ros95, WS95, Bos97, Dre98, EL99, HKT97, Ish96a, Ish96b, Kar96, WS97, YB99]. Ratio [AP95b, BG90b, HW94, MS91, MRS92, Owe90, Roj92, SV95, Was92, DJ96, LHMY96, MS95, MvdV97, Qin99, Rub96, VZ97]. rational [DHI99, HSS96]. Ratios [Wij85, Wij93, WS95, CS97]. RC [RG91]. Reaching [GC93]. Real [LM90, Wu95a]. receiver [HT96]. recognition [Kle96]. Reconstruction [O’S95]. record [FH96]. record-value [FH96]. Recovery [Don94, MT95a]. Recurrence [Eat92]. Recursive [RT92a, Ruk94, SM94]. Redescending [KT91]. reduced [And99]. Reducing [MPB96]. Reduction [CKW95, RC94, CF99, CLW99, NY97]. redundancy [DHL97]. Reference [CE90, YB94]. Refined [Dre95]. Region [MK94]. Regions [Ber93, BH93a, cICM92, HL90c, Owe90, PR94, SWZ98, Zha94, ET98, MZ96, MNS98a]. Regression [Asm90, BX91, Bar93, BL92, BM94, Ber90, BR91b, Bre91, BTZ93, Bro90a, Bro90b, BL91, BDHT91, CA91, CKW95, Cas90, cICM92, Cha91, Che91a, CV92, Chi91b, CKS94b, Cop90, Cox93, CM95b, Cuz88, Cuz90, Cuz92, Dah95, Dav90, Dav93, Det90, Det93b, Det94, DHS95, Det95, DGKL94, DF93, DS95a, DH93, DL91c, Efr90, EKS91, EH92, FG92, Fan93b, FT93, FG94, FR90, Fri91a, Fri91b, FZ94, GD92, Gir91, Gle90, GL92, GN90, GH91b, GW91, GJ92, HJ90, Hal92b, HM90, HM91, HM93, HW91, Hei94, Hil90, HC92, HL90c, JV93, Joh90, Jon92, KE95, KN98, Lah92, LY91b, LY91c, Lai94, LY94, LD95, Lu90, Mam91a, Mam91b, MY93, MU90, Mil92, Mor90, Muk92, Muk93b, Mul92, NT94, O’S91, Owe91a, Por85, Por91, PS93, RG97, Rit90c, Rob95b, RT92a]. Regression [RW94, Sas93, Sch93, Sch95, Sch91, Sen96, SWZ98, Sri92, Ste90a, Sto91b, SL94, Tan93, Tsi90, WC93, Yaj91, Yin93a, Zam92, Zha91, Zho92, vdG90, And99, BH99, BFH96, Bis98, BL96a, CGWW98, CDS97, CLW99, CHST99, Dab97, Det96, DM98b, DH98, DM98c, Det99, DHI99, Efr96, EHSS95, FGK97, Gir98, HLP95, HT97, HS96a, Hsi97, Hsi99, Hua98, JT99, KMZ98, Kni98, Koc96, Kor96, KPY99, KK96, LW99, Li97, LZ98a, Loa96a, MvdG97a, MM97, MC96, MNS98b, MS99a, MS99b, Neu95, NY97, Nob96, Nob99, OR97, Qiu98, RSW97, Rob97, RH97, SZ98, SY98, Stu97, STZ98, TRYV96, VA99, Wan96a, WvdG96, Yan97, ZL96, ZP96, ZF96]. Regression-type [KN98]. Regressions [BH92c, LM98]. Regressors [Kne94a]. Regret [Tak92]. Regular [CM98, MW99a, Rit96, Wei95]. Regularization [Gir91]. regularly [AM97a]. Rejoinder [Bro90b, DBS94b, Fri91b, Sto94a, Tie94b, Bre98, LPS99]. Related [ACG94, CO90, JS90, KS95, Lin94, Mas94, Hos97, Jia96, MU96]. Relationship [HWG65, HWG95, RT97]. Relationships [BH93a, Zha94]. Relative [Hal91, HO97, RS98]. Relaxed [Oeh92]. Remainder [Dud94]. Remarks [Buj90, NO96]. REML [Jia96, Jia97]. Remote [AB90]. Renewal [DMR90, GP93b, Phe90]. Renormalization [DL92, Low92b]. Renormalizing [Low93]. Rényi [AGW90]. Repair [HPS92b, DHS97]. repairing [vdL96]. Repeated [HZ90, HZ92, HRC94, KG91, SW92b, Kus97, Kus98]. replicate [MNS98b]. Replication [KG91, SW92b]. Representation [Cha91, Che92b, GV91, Hes90, Arc96, Bos98, HS96a, SSK95]. Representations [Hal91, Stu93]. Representing [KG92]. Resampling [BHW93, Hal91, HM94, LS92a, PR92, PvZ96]. Residual [KO94, CZ96, Kou96, LW99, MM97]. Residuals [Bai94, Lin98, Mam96]. Resistant [Jac94, Kle96]. Respect [Xia95]. Response [Jun93, MS93, MSY94, Ros93, SM90, Sto91a]. Response-Adaptive [Ros93]. Restricted [CS93b, CKS94a, HS95, MS91, RG91, CS96, MS95]. Restrictions [CD91, HP94, NF91, Pol98, PH96]. Results [BB93, CW91, Che92a, Che95e, Lel93, Shi91, Ste95a, Vel92, BS99]. Retinopathy [WWG^{+}95]. Reverse [HH94]. Revisited [Gen92, Owe94, Str90, WMC98]. Reweighted [HP92, Lop99]. Riccati [DS97]. Ridge [Gir91]. Riemannian [Hen90, LK93]. Right [FZ94, LY91c, LvE93, MBL91, BM97, MW99b, vdLM98]. Right-Censored [LY91c, vdLM98]. Rigid [CK95]. Risk [BG90c, BL91, DLM90, Efr95, FG94, GK94, GN90, Low93, Mar83, Mar86, Mar91b, Tak94, BL96b, DS97, HO97, SA96, SV98]. Rissanen [HG90]. Ritov [Bei96, EL96a]. Robbins [Liu97]. Roberts [SW93b, SW94b]. Robust [BB98, BZ96, BF90, DS91, Dav93, DL90, Eri96, GP95, HS92, Hug96, Kim92, Ko92, KS99c, KM92, MY93, MZ93a, MZ93b, MC96, SM94, She95, Zam92, Adr98, MM99, Mül98, Sta96]. robustify [CAGM97]. Robustness [BD95, Bos94, Lin94, LS92a, Roc96]. Root [Aba93, CT95, JMP91, KP90a, Muk92, Wu95b]. roots [BRF97]. Rosenblatt [GH91a]. Rousseeuw [Dav92]. rows [MU96]. Rubin [CL96a]. Rule [KZ95, Ruk94, Sam92]. Rules [Hal93, Mar92, HKP98, Maj96b]. Run [BH90b, Cos93, Jac94, Yak95, GP97, Muk99, Yak98]. runs [HP99].

Saddlepoint [JR94, RS98]. same [HP92]. Sample [Alb91, BR91a, BR91b, HS95, HL90b, JK95, Kan92, KG92, KG91, LY91b, LvE93, MBL91, Neu93, PR94, Qin93, SW92b, Shi91, Sto90, TC85, TC90, Tyl94, VZ92, Wu95a, Xia95, XY95, Yin93a, Con99, DM98a, GW99, Hsi96, MPB96, Rob97, RT97, Rub96]. Sampled [BGL95]. Sampler [Yu93, Ami96, GMW96]. samplers [GMW98]. Samples [KG91, BC98, Hua97]. Sampling [APRY93, BLEL94, BC92, BR91b, BNW92, BG93, Dos94a, GL92, Lo93b, Loh90, Owe94, Ros95, SW94b, Yeh94, Yin93b, Che98, DS98, DHL97, Kol99, KLW97, Liu97, Loh96a, Loh96b, NW97, Sho96, Ste95b, Ste99, Wal97b, Yak96a, vdV96]. saturated [MW95]. Scale [GK94, MZ93a, SW95, Vel92, Dav98, Hsi96, Ish96a]. Scaling [BLRS94]. Scanning [Khm93]. Scatter [KT91, Tyl94, KT96a, Lop99]. Scheme [GP94, GP95, PR92, Yin93b, Liu97, WM97, vdV96]. schemes [Che98, GP97, Ish98]. Score [CH91, SC91, VMT95, DS99b]. Scores [GJ92, KS95, Seo90, GKL98, GKL99, HJ99, Sen96]. Scoring [KDK90]. Scrambled [Owe97]. Search [Che92a, FZS96, STS96]. Searching [STS96]. Second [Alb91, BPR99, GK94, GSM94, GK96, Mar83, Mar86, Mar91b, Myk95b, Ste90b, KS96a, LL97b]. Second-Order [GK94, GSM94, Ste90b, BPR99, GK96, LL97b]. seemingly [KK96]. Segments [KDK90]. Selected [ND94, Vel92]. Selection [Chi91a, EH92, FM92, GD92, Jin92, Zha91, Zha93, BB98, Bre96, Loa99, LN99]. Selector [JMP91, Che97]. Selectors [Chi91b, CM91, LMS97]. Self [Cha90b, GZ93, TLF95, TC85, TC90, MR96]. Self-Consistency [TC85, TC90]. Self-Consistent [Cha90b, GZ93, TLF95, MR96]. Semi [DW93]. Semi-Bayesian [DW93]. Semiparametric [BZ97, CKW95, Che95b, Cuz92, HM90, HSS97a, LY95b, Mag93, MZ95, MvdV97, RB90, Rob94b, Rob95a, Roe92, She95, CHS96, HKT97, HT96, Ish96b, Ish98, Ish99, LHC99, NLB98, Van96]. Sense [DG90]. Sensitive [Mar91a, MS83, MS92]. sensitivity [GW95b, Gus96]. Sequence [AGW90, Hsi93, Neu94, Rit96]. Sequences [BS91a, BS91b, Buh94, Dev94, KDK90, CBL99, Hes96]. Sequential [Bai94, BBW94, BB93, BG92, CH90b, Efr95, GP94, GL91, HWG65, HWG95, HL90c, Lee96, Mar83, Mar86, Mar91b, OAT91, PS91, Sie93, SV95, SBH91, Sri92, Tak92, WH90, XM95, Bet96, FZS96, Gre96, HM98, Kou96, Lin98, Liu96, Liu97]. Serial [HMM92, TM93]. Series [Bel84, Bel91, CL91, Deg90, GH94, KM90, MT95b, MZ94, NF91, Rob94b, Rob95b, Tra93, TS92b, Büh98, DW96a, DJ96, Dah97, DKW97, GT99, KMZ98, Kou96, KS99a, LW99, LL98, MYG96, MS99a, RH97, TRYV96]. Set [CD91, Hua97]. Sets [Che95e, Das91, DGZ95, Ell95, FH91, MT95a, Nau92, SW93a, Wor95, And96, BD98, Cho98, Düm98, Huw96, TB97, Tsy97, ZP96]. setting [Efr96]. Several [DKR91, JV93, DM98b]. Shape [GM93, HS95, KE95, LK93, PM95, CW99, Roc96]. Shape-Restricted [HS95]. sharp [Efr98, Rai98, Tsy98]. Shift [SW93b, Maj96b]. Shiryayev [SW93b, SW94b]. short [HLP95]. short- [HLP95]. Shrinkage [ES96, Geo91, Sun96, BC98, DJ98, Gou96, Wan96b]. sibships [DS99b]. Side [AB90]. Sided [BG92, Fin94, Kur93, LB95, CS98]. Sieve [Büh98, SW94a, WS95]. sieves [Jia97, She97]. sign [MOT97, Xu97]. sign-invariant [Xu97]. Signal [Bel84, Bel91, RS95, SW95]. Signed [Seo90]. silhouette [Pol98]. Simes [Sar98]. Similar [Kne94a]. Simple [BBW94, BTZ93, JMP91, Jon92, Cas96]. Simplices [Liu90]. Simulation [KL95, ADS96]. simulations [Nak97]. Simultaneous [Cho91, Efr98, HM91, JJ90, SL94, Xu97]. Single [HHI93, HSS97a, Hug96]. Single-Index [HHI93]. Singular [Ell95, LB93, Uhl94, DGJ97]. Singularities [Ell91]. Size [cICM92, FZS96]. size-dependent [FZS96]. Skewness [AM97b, BH92a]. Skorokhod [ES96]. Sliced [HC92, ZF96]. Slicing [DL91c]. Slope [HRC94, BFH96]. Small [DG91, RS98]. Smallest [DBS94a, DBS94b, Flu94]. Smirnov [MS92, CC94, Cab96, CC97, IL90, MS83, YP92]. Smooth [KL95, Mac95, Mam91a, MT95a, MT99, Peñ98, GG99b, HR98, IKL97, MS99b, Owe97, TK97]. Smoothed [CH93, CH95, Dab97, EL95, RC94, Sil96]. Smoothers [FG92, Fan93b, Kne94b, Nyc95, Koc96]. Smoothing [Dev94, Far92, GQ93, Hal93, HHI93, Jin92, Nyc90, Oeh92, Ste93, SL94, WWG^{+}95, Büh98, DL96, DL97, Luo98, Por97, Wal97a]. Smoothness [Mam91b, Bir97, BZ98, LMS97]. Sobolev [GN90, Tsy98]. Solution [Mac95]. Some [And96, AG91, BNO96, BS99, CH94b, CW91, Che92a, CB91, Che95f, Chi91b, Dat92, Dum91, ES93, JJ90, KS94, Lav92, Mas94, Mat93, NO96, Ros91, Sar98, Sas93, SS97, Shi91, Ste95a, SCW97, Yan92, Zho91, KL92, Kou96, Ric95]. Space [CK95, HW94, Jon92, PS95, CP98, JP99, Niu98, Wal97b]. space-time [Niu98]. spaced [HT97]. Spaces [JK95, LK93, WS91]. spacing [Swa99]. Spacings [Kan92]. Sparse [DS95b, Joh94, SQ96]. Spatial [MGR98, Yin93b, BG97, BRF97, LMS97, MOT97, vdV96]. Special [Szk92, Res97]. specially [ET96]. Species [Lo92]. specified [DW96b, GP97]. spectra [NvS97]. Spectral [And93, FH92, Ioa94, MP92, PR92]. Speed [GG99a, JS90]. Sphere [CK95, HK96]. Spheres [LS92b]. Spherical [Bar91, BS91c, Nai90]. Spherically [FW95, GKL98, GKL99]. Spline [Che91b, GQ93, Mes91, Nyc90, Ste93, Sto90, TA91, Luo98, WWG^{+}95]. Splines [BX91, Bre91, BDHT91, Buj94, Fri91a, Fri91b, GH91b, GW91, Has94, Nyc95, O’S91, Oeh92, Owe91a, Sch91, SWZ98, Sto91b, Sto94a, Sto94b, MvdG97a, Por97, SHKT97]. split [BS99, KLW97]. split-plot [BS99]. Square [AA90, IL90, Kim93, LD93, CF99, EL96a]. Squared [DB90, HP95, MW92, Mar92, Nyc90, Wu95b]. Squares [BP91, Buj90, Cha93, CS93a, Csi91, Lai94, O’S95, Pat93, Pos92, RW94, Wei92, vE92, Gou96, KK96, Wan96a, WvdG96]. stability [SY98]. stabilization [Bre96]. Stabilized [Chi91b]. stable [KM96]. Stacy [WM97]. Stage [Alb91]. staggered [BGY97]. Stahel [GH97a]. Standard [BD95]. Standardized [Che90a, HS96b]. Start [HG95]. State [CP98, JP99, Nak97]. Stationarity [MT95b, Mil92]. Stationary [Buh94, Din95, Gas93, Hes90, HT82, HT93, Hsi93, NNR94, Pit94, SF94, SW95, Yaj91, DCG99, GK96, MPZ98, MYG96]. Statist [Ano95, Kus98]. Statistic [BG90b, Hel91, SV95, Rub96]. Statistical [BRR91, CMS93, DM92, DL93, DL95, Don94, DKR91, Eat92, FR92, FZ99, KDK90, KG92, Koo93, Lav92, Lav94, Nai90, OC95, Sha93, AGSC98, Hos97]. Statistics [AG92, BG93, Cha92, DCLP90, GH94, GL91, HJ93, IL90, LK93, MN92, ND94, Ros91, Ruk90, Seo90, Sha94, Ste95a, Stu94, Tyl94, Xia94, BGvZ97, DJ96, FR98, HS96b, Hös96, LPS99, Mat96, PvZ98, Sho96, ZBCL97]. Stay [Sam92]. Stay-with-a-Winner [Sam92]. steady [Nak97]. steady-state [Nak97]. Stein [BS91c, SS94, Szk92]. Step [MP90, FR98, SY98]. Step-Down [MP90, FR98]. step-up [FR98]. Steps [Wij85, Wij93]. Stepwise [Fin99]. Stochastic [BC92, DKR91, Hey92, Kon92, Lai94, Ste90d, Swe92, TA91, Yin90, CS96, DLM99, ES96, Kle96, LW99, Niu98]. Stochastically [GP95]. Stopped [AW93]. Stopping [Mar92, FZS96]. Stratification [FZ94]. Stratified [KG91]. Strong [CCL95, CHY99, DGKL94, Neu98, SW93c, CL97, HY98b, Hes96]. structural [GW98, Huw96]. Structure [LK93, PS95, Ste90b, DW96b, Yak98]. structured [MC96]. structures [BS98]. Student [GC92]. Studentised [HL90a]. Studentized [CS93b, Hel91]. Study [Asm90, O’S95, TC85, TC90, VZ92, WWG^{+}95, Yin93a, Jam97]. Sturdy [Mar91a]. Subject [HP94, NF91, SM90, DE98]. Subjective [BR90]. Submartingale [HH94]. Submodels [KZ92]. Subproblems [Fan93a]. Subsamples [PR94]. Substitution [Dos94a, PvZ96]. Successive [Dos94a]. Sufficiency [HWG65, HWG95, Lar90, MS93, BNO96, NO96]. Sufficient [Lar90, Ste95a, MW98]. sufficiently [Ste97]. sufficientness [MW99b]. Sums [BH90b, Lah95]. sup [Tsy98]. sup-norm [Tsy98]. Superefficiency [BLZ97, BH98]. Superkernels [Dev92]. Support [PT91, CF97]. Suprema [Zam91]. Sure [GV91, Stu93]. Surface [MSY94]. surfaces [Qiu98]. Surveillance [GP95, GP97, Yak98]. Survey [FF91, GJ90, Sha94]. Survival [Cha90b, GJ90, JC90, Mag93, MNS95, Zho92, ZK90, LHMY96, LK99, NLB98, PH96, WG96]. Symmetric [ET91, Eat92, FW95, KW96, Mas94, Sri92, WK92, And96, BGvZ97, CL96a, PvZ98, Swa99]. Symmetry [AM98, AG91, Bar91, BS91c, Pin94]. Synchronizing [GW99]. Synthetic [Zho92]. system [GG99a]. Systems [Li95, DHL97, GG96a].

Tables [DS95b, RG91, HJ96b, LL97a, SQ96, Str99]. Tail [Hsi91, Lah95, MS83, MS92, Roj92, SF94, VMT95, vZ91, FH99, HW97a, Res97, RS98, VA99]. Tailed [Muk93b, SM94, DM98a]. Tails [Loc91, AM97a]. tangent [Eri96]. Target [CF99, HF92]. Techniques [EL92]. teletraffic [Res97]. tends [SY99]. Tensor [Buj94, Has94, Sto94a, Sto94b, SHKT97]. Terms [EJ90, TW96]. Terry [SY99]. Test [AP95b, BH90a, BBW94, GH91a, HMM92, HL90a, Kur93, Maj92, MS83, MS92, MNS95, MS91, MZ95, NT94, Pin94, BHM97, Det99, DS99b, FR98, Fin99, HP99, HJ96b, MS95]. Testing [Bar91, Bar93, BBW94, BM92, BN91, Cha90a, CKS94b, DCG99, DHY95, DP94, EL92, EH92, EHSS95, GL91, HKL92, Hor93, HCR^{+}92, Kre90, MT95b, MS91, Mil92, RG91, Ruk94, SW95, SM90, SW99, CH99, CS96, Fal95, MN98, MS95, Min97, Mül98, Spo96, VW98]. Tests [Alb91, AA90, And93, AG91, BL92, BG92, CC94, CL92, Che95a, CS93b, CKS94a, CAMRRdB99, DB90, FM93, Fer95, Fin94, Hjo90a, HL90b, HW94, Jan91, JC90, KL95, Kim93, Kon92, LD93, LB95, MRS92, Neu93, SC91, TM93, Tsi90, Cab96, CC97, CHS96, CS98, DN98, Düm98, GKL98, GKL99, HJ99, IL96, IKL97, Kar96, LL97a, LK99, MNS98b, MOT97, Peñ98, SZ98, Sho96, VZ97]. Their [Cha91, Che91b, Fan93b, Sto94b, Buj94, Düm98, Efr98, HM94, Has94, Sto94a, SHKT97, TW96]. Theorem [BRR91, Det93a, HT82, HT93, JJ90, Mas94, Stu95, Szk92, Bis98, DS97, Fre99, GEGM99, HS96b, Loh96a, Rub96, ZBCL97, WK90]. Theorems [CGH91, EW95, HJ95]. Theoretic [Rit90b, YB99]. Theoretical [Lah99, BS99]. Theory [BR91b, CC99, DMR90, GL92, GM93, GP93b, GQ93, Hal90, HI90, HC92, LY91b, Mur95, NT94, OAT91, BH98, BGY97, DM98b, Hos97, Myk99, Par98, PH96, Sin98]. Three [Sie93, Bet96, HSS97b]. Threshold [Cha90a, Cha93, HKP98]. thresholding [Cai99, DJKP96, NvS97]. Tied [VMT95]. Time [Bel84, Bel91, CH90b, CL91, Deg90, GH94, GL91, HF92, JK95, MT95b, MZ94, MNS95, NF91, Pit94, Rob94b, Rob95b, RH97, TM93, Tra93, TS92b, ZK90, Bos97, Büh98, Con99, DW96a, DJ96, Dah97, DKW97, GT99, KMZ98, Kou96, KS99a, LW99, LL98, MYG96, MS99a, Niu98, TRYV96, WG97]. Time-Dependent [ZK90]. Time-Sequential [CH90b, GL91]. time-series [DKW97]. Times [HF92]. Tolerance [MNS98a]. Tomography [CH94a, JS90, O’S95]. Tools [KG92]. topics [Jia96]. Topological [DP94]. Tossing [AGW90]. Total [DG90, Mee92]. Tradeoffs [Low95]. Training [Jon92]. Transfer [VA99]. transform [KS96a]. Transformation [HR95, Mas94, Szk92, DGJ97, McC96, YD97]. Transformations [Bur91, Cab96, CL92, Che95a, Kon91, RC94]. Transformed [CC97, JR94]. Transforms [Gol91]. Transitivity [Gre93]. Treatment [MBL91, Ros93, Maj96a]. treatment-control [Maj96a]. Treatments [Maj92, Sie93, Bet96]. Tree [Lav92, Lav94]. Trees [MSW92, Hug96]. Trend [CM93, Cos93, Jac94, LD91, Mil92, FH96]. Trend-Free [CM93, Cos93, LD91]. Trial [Sie93, Bet96]. Trials [GL91, WS97]. Triangular [PR92]. Trimmed [CAGM97, Kim92, GEGM99]. trimming [KM97]. Triple [Loh90]. Truncated [CCL95, LY91a, LY91c, LY94, Stu93, Fal98]. Truncation [KG90, HY98a, HY98b]. Tube [NW92]. Tube-Like [NW92]. Tubes [JJ90, SZ93b, EM99, NW97]. Tubular [Nai90]. Two [Alb91, BB93, BG92, Cha94, CB91, CMT98, CM91, Fin94, HK91, HZ90, HZ92, HL90b, Kur93, Loh91b, MBL91, MU91, Neu93, Sun94, Wu95a, AM97b, BB96, DT98, HJ96b, Hsi96, MPB96, RT97, Sun96]. Two-Armed [Cha94]. Two-Associate-Class [CB91]. Two-Dimensional [MU91]. Two-level [CMT98]. Two-Parameter [BB93, Sun94, BB96]. Two-Period [HZ90, HZ92]. Two-Sample [MBL91, Neu93, Wu95a, Hsi96, MPB96]. Two-Sided [BG92, Fin94]. Two-Stage [Alb91]. two-way [DT98, HJ96b, Sun96]. Tyler [GH97a]. Type [CC94, Cha92, Che91a, CL91, Hes90, Ioa94, Koy90, LvE93, Myk94, Seo90, WC93, XY95, Yin92, HP96, KN98, MM97, Yeh94].

Uhlig [DGJ97]. UMVU [YG95]. Unbalanced [FH91, BZ96]. Unbiased [CKS94b, LB93, BHM97]. Unbiasedness [BM92]. Unbounded [AP95a]. Unfathomable [Che90b]. Unidimensional [Jun93, JE97]. Unified [BBW94, Kub94, Mas94]. Uniform [CCL95, Che95d, DF90, DKR91, Hal91, HJ95, Ish96b, Ste90b, Tan93, CL97, Har96, HJ96b]. uniformity [IL96, SW99]. Uniformly [LB95, CHS96, Dav98]. unimodal [Bir97]. Unique [Che92b]. Unit [Aba93, CT95, BRF97]. Univariate [AG91]. Universal [Csö96, DGKL94, DL97]. Universally [Stu94, UM97, DL96]. Unknown [BP91, Cuz92, HKL92, Low92a, PS91, Rit90a, SW95, TM93]. unrelated [KK96]. Unstable [BMM^{+}91, CT95, LL98]. Upper [Low93, WK92, KK96]. urn [Ish98, WM97]. Use [Buj94, Has94, Sto94a, Sto94b, MNS98a, ZP96]. Used [ES93]. Usefulness [Dev92]. Using [Cha92, DBS94a, DBS94b, ET96, Flu94, FZ94, GC93, Hen90, Hsi91, LD95, Loh90, Mar92, Pos92, SZ93a, SV95, Ste90c, Ste90b, Tsi90, Wor95, YB94, DCG99, Dat92, DH93, DS99b, HKP98, HR98, KP96, Loa96a, LN99, Niu98, Nob96, Spo96, Ste95c, VW98]. Utilizing [Maj92].

vaccine [WS97]. Validation [DM98c, Pat93, Ste90d, Zha93, vE92, Gir98]. Validity [HB91, NNR94]. Value [HF92, Low92a, PS91, Dre95, Dre98, Fal95, FH96]. Valued [SF94]. Values [BD93, DHY95, HMM92, Mar91a, Men94b, WC93]. Variable [BW99, FG92, GD92, Hal92c, HHM95, Kan92, TS92a, Yeh94, Zha91, JE97, LMS97]. Variable-Selection [GD92]. Variables [BH93a, Bur91, CV92, FT93, JR94, Lah95, Myk92, PR92, ST92, Zho91, AM97a, BM97, Cha96, Huw96, LHC99, MW98, Sal98, Sar98]. Variance [CD91, Fel91, FH91, GC91, Hes90, KZ92, LM90, Low95, MGKA95, MZ95, Nyc90, Owe94, Ros95, Sho90, WH91, DT98, Det99, GT99, Has99, Hug96, KT96b, KS99b, Owe97, Stu96]. Variates [Buj90]. Variation [DG90, Dud94, MM99, Wei95]. Variational [Haf91, Mac95]. Varietal [LD91]. varying [AM97a, FZ99]. Vector [JM92, Joh94, Pos92]. Vectors [JV93]. Version [KL95, AGSC98, DLM99]. Versions [Gir91]. Versus [HM93, HS93, Lin94, Pfa93, HSS97a, MS99b, SW99]. vertical [Eri96]. Via [Dos94a, EH92, Eva91, KZ92, TC85, Zha93, Bro98, CW99, DJ98, DHL97, DE98, EM99, Fal95, HM98, Ish98, Liu96, Lo92, MW99b, NW92, ST92, Spo98, TC90, Wan96b]. View [GJ90, HI90, Mas94]. Vincentization [Gen92]. vol [Ano95]. Volume [Ano95, Das91, Dav92, JJ90, SZ93b, HJ96b]. Volumes [Nai90]. voting [BFLS98].

Waiting [Pit94]. Wald [Fre99, Jia97, SHKT97]. Walk [NT94]. Walks [Asm90]. warping [WG97]. Wasserstein [CAMRRdB99]. Wavelet [BC98, HP95, NvS97, Cai99, DJKP96, DJ98, HMT96, HT97, HKP98, PV99, Wan96b]. Wavelet-Based [HP95]. wavelets [KP96, Spo96]. way [DT98, HJ96b, Sun96]. Weak [Bai94, Cha90b, GL91, KO94, Lin98, MP92, CH96, CL97, MLN99, MW98]. Weakly [Hsi93, KP90b, Kit97, Neu98]. Wedgelets [Don99]. Weibull [EM90, Ish96b]. Weighted [BH90b, CS93a, Che98, Det93b, Hei94, KO94, Lo93b, RW94, vZ91, DHI99, Jam97]. Weighting [JV93]. Weights [PT91, TK97]. Weiss [CG91]. Wellner [CY96]. Which [Bro90a, BLEL94, BM97, Ber90, BR96, Bro90b, Cas90, Cop90, Efr90, FR90, Gle90, Hil90, Joh90, Lu90, Mor90, Ste90a, Yak98]. White [KL92, Low92b, Low93, RS95, BL96a, Nis99, Nus96, Zha97]. Whittle [GT99]. Wicksell [GL98, GJ95]. Wielandt [ET91]. Wild [CA91, Mam93]. Window [HHM95, Spo98]. Winner [Sam92]. Winsorizing [Sho96]. Wisconsin [WWG^{+}95]. Wishart [CL96a, LM98, Uhl94]. without [Bir97]. Worst [Fos91].

Yatracos [DL97]. Yeh [CM93]. yields [Gou96].

Zero [KL91]. Zero-Crossing [KL91]. Zonoid [KM97].

References

Amemiya:1990:ACS

[AA90] Yasuo Amemiya and T. W. Anderson. Asymptotic chi-square tests for a large class of factor analysis models. Annals of Statistics, 18(3):1453–1463, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

              //projecteuclid.org/euclid.aos/1176347760.

Ahlswede:1990:MEP

[AB90] R. Ahlswede and M. V. Burnashev. On minimax estimation in the presence of side information about remote data. Annals of Statistics, 18(1):141–171, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347496.

Abadir:1993:LDA

[Aba93] Karim M. Abadir. The limiting distribution of the autocorrelation coefficient under a unit root. Annals of Statistics, 21 (2):1058–1070, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349164.

Alvo:1991:BIB

[AC91] M. Alvo and P. Cabilio. On the balanced incomplete block design for rankings. Annals of Statistics, 19(3):1597–1613, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1176348264.

Arcones:1994:ERP

[ACG94] Miguel A. Arcones, Zhiqiang Chen, and Evarist Gine. Estimators related to U-processes with applications to multivariate medians: Asymptotic normality. Annals of Statistics, 22(3):1460–1477, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325637.

Adrover:1998:MBR

[Adr98] Jorge G. Adrover. Minimax bias-robust estimation of the dispersion matrix of a multivariate distribution. Annals of Statistics, 26(6):2301–2320, December 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691472.

Athreya:1996:CMC

[ADS96] Krishna B. Athreya, Hani Doss, and Jayaram Sethuraman. On the convergence of the Markov chain simulation method. Annals of Statistics, 24(1):69–100, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1033066200.

Arcones:1991:SBT

[AG91] Miguel A. Arcones and Evarist Gine. Some bootstrap tests of symmetry for univariate continuous distributions. Annals of Statistics, 19(3):1496–1511, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348258.

Arcones:1992:BS

[AG92] Miguel A. Arcones and Evarist Gine. On the bootstrap of U and V statistics. Annals of Statistics, 20(2):655–674, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348650.

Assaf:1998:SVP

[AGSC98] David Assaf, Larry Goldstein, and Ester Samuel-Cahn. A statistical version of prophet inequalities. Annals of Statistics, 26 (3):1190–1197, June 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1024691094.

Arratia:1990:ERL

[AGW90] R. Arratia, L. Gordon, and M. S. Waterman. The Erdős–Rényi law in distribution, for coin tossing and sequence matching. Annals of Statistics, 18(2):539–570, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347615.

Akritas:1994:NNE

[Akr94] Michael G. Akritas. Nearest neighbor estimation of a bivariate distribution under random censoring. Annals of Statistics, 22(3):1299–1327, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325630.

Albers:1991:SOA

[Alb91] Willem Albers. Second order analysis of two-stage rank tests for the one-sample problem. Annals of Statistics, 19(2):1042–1052, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176348136.

Anderson:1997:PMA

[AM97a] Paul L. Anderson and Mark M. Meerschaert. Periodic moving averages of random variables with regularly varying tails. Annals of Statistics, 25(2):771–785, April 1997. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1031833673.

Averous:1997:SMD

[AM97b] Jean Avérous and Michel Meste. Skewness for multivariate distributions: two approaches. Annals of Statistics, 25(5): 1984–1997, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362381.

Andersson:1998:SLC

[AM98] Steen Andersson and Jesper Madsen. Symmetry and lattice conditional independence in a multivariate normal distribution. Annals of Statistics, 26(2):525–572, April 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1028144848.

Amit:1996:CPG

[Ami96] Yali Amit. Convergence properties of the Gibbs sampler for perturbations of Gaussians. Annals of Statistics, 24(1):122– 140, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1033066202.

Andersson:1997:CME

[AMP97] Steen A. Andersson, David Madigan, and Michael D. Perlman. A characterization of Markov equivalence classes for acyclic digraphs. Annals of Statistics, 25(2):505–541, April 1997. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1031833662.

Anderson:1993:GFT

[And93] T. W. Anderson. Goodness of fit tests for spectral distributions. Annals of Statistics, 21(2):830–847, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349153.

Anderson:1996:SIS

[And96] T. W. Anderson. Some inequalities for symmetric convex sets with applications. Annals of Statistics, 24(2):753–762, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032894463.

Anderson:1999:ADR

[And99] T. W. Anderson. Asymptotic distribution of the reduced rank regression estimator under general conditions. Annals of Statistics, 27(4):1141–1154, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1017938918.

Anonymous:1990:PK

[Ano90] Anonymous. Preface to the Kolmogorov. Annals of Statistics, 18(3):985–986, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347734.

Anonymous:1995:VCA

[Ano95] Anonymous. Volume contents, Ann. Statist., vol. 23, no. 6 (1995). Annals of Statistics, 23(6):iii–xiv, December 1995. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1034713637.

Andersson:1993:LMC

[AP93] Steen Arne Andersson and Michael D. Perlman. Lattice models for conditional independence in a multivariate normal distribution. Annals of Statistics, 21(3):1318–1358, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349261.

Anderson:1995:NAW

[AP95a] Charles Anderson and Nabendu Pal. A note on admissibility when precision is unbounded. Annals of Statistics, 23 (2):593–597, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324537.

Andrews:1995:ALR

[AP95b] Donald W. K. Andrews and Werner Ploberger. Admissibility of the likelihood ratio test when a nuisance parameter is present only under the alternative. Annals of Statistics, 23(5): 1609–1629, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324316.

Assaf:1993:DCN

[APRY93] David Assaf, Moshe Pollak, Ya’acov Ritov, and Benjamin Yakir. Detecting a change of a normal mean by dynamic sampling with a probability bound on a false alarm. Annals of Statistics, 21(3):1155–1165, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349255.

Arcones:1995:AAB

[Arc95] Miguel A. Arcones. The asymptotic accuracy of the bootstrap of U-quantiles. Annals of Statistics, 23(5):1802–1822, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176324324.

Arcones:1996:BKR

[Arc96] Miguel A. Arcones. The Bahadur–Kiefer representation for U-quantiles. Annals of Statistics, 24(3):1400–1422, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032526976.

Asmussen:1990:EFR

[Asm90] Soren Asmussen. Exponential families and regression in the Monte Carlo study of queues and random walks. Annals of Statistics, 18(4):1851–1867, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347883.

Aras:1993:AEM

[AW93] Girish Aras and Michael Woodroofe. Asymptotic expansions for the moments of a randomly stopped average. Annals of Statistics, 21(1):503–519, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349039.

Bai:1994:WCS

[Bai94] Jushan Bai. Weak convergence of the sequential empirical processes of residuals in ARMA models. Annals of Statistics, 22(4): 2051–2061, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325771.

Baringhaus:1991:TSS

[Bar91] Ludwig Baringhaus. Testing for spherical symmetry of a multivariate distribution. Annals of Statistics, 19(2):899–917, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348127.

Barry:1993:TAR

[Bar93] Daniel Barry. Testing for additivity of a regression function. Annals of Statistics, 21(1):235–254, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349024.

Bose:1993:SER

[BB93] Arup Bose and Benzion Boukai. Sequential estimation results for a two-parameter exponential family of distributions. Annals of Statistics, 21(1):484–502, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349038.

Bose:1996:EPP

[BB96] Arup Bose and Benzion Boukai. Estimation with prescribed proportional accuracy for a two-parameter exponential family of distributions. Annals of Statistics, 24(4):1792–1803, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032298295.

Bayarri:1998:RBA

[BB98] M. J. Bayarri and James Berger. Robust Bayesian analysis of selection models. Annals of Statistics, 26(2):645–659, April 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1028144852.

Berger:1994:UCF

[BBW94] James O. Berger, Lawrence D. Brown, and Robert L. Wolpert. A unified conditional frequentist and Bayesian test for fixed and sequential simple hypothesis testing. Annals of Statistics, 22(4): 1787–1807, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325757.

Benhenni:1992:SDE

[BC92] Karim Benhenni and Stamatis Cambanis. Sampling designs for estimating integrals of stochastic processes. Annals of Statistics, 20(1):161–194, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348517.

Brown:1998:WSN

[BC98] Lawrence D. Brown and T. Tony Cai. Wavelet shrinkage for nonequispaced samples. Annals of Statistics, 26(5):1783– 1799, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1024691357.

Block:1990:POP

[BCFS90] H. W. Block, D. Chhetry, Z. Fang, and A. R. Sampson. Partial orders on permutations and dependence orderings on bivariate empirical distributions. Annals of Statistics, 18(4):1840– 1850, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347882.

Berry:1997:BPI

[BCZ^{+}97] Donald A. Berry, Robert W. Chen, Alan Zame, David C. Heath, and Larry A. Shepp. Bandit problems with infinitely many arms. Annals of Statistics, 25(5):2103–2116, October 1997. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1069362389.

Bjerve:1993:CCM

[BD93] Steinar Bjerve and Kjell Doksum. Correlation curves: Measures of association as functions of covariate values. Annals of Statistics, 21(2):890–902, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176349156.

Basu:1995:RSC

[BD95] Sanjib Basu and Anirban DasGupta. Robustness of standard confidence intervals for location parameters under departure from normality. Annals of Statistics, 23(4):1433–1442, August 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176324716.

Beran:1998:MEC

[BD98] Rudolf Beran and Lutz Dümbgen. Modulation of estimators and confidence sets. Annals of Statistics, 26(5):1826–1856, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1024691359.

Buja:1991:DMA

[BDHT91] Andreas Buja, Diane Duffy, Trevor Hastie, and Robert Tibshirani. Discussion: Multivariate adaptive regression splines. Annals of Statistics, 19(1):93–99, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347967.

Butler:1993:AMC

[BDJ93] R. W. Butler, P. L. Davies, and M. Jhun. Asymptotics for the minimum covariance determinant estimator. Annals of Statistics, 21(3):1385–1400, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176349264.

Beibel:1996:NRB

[Bei96] M. Beibel. A note on Ritov’s Bayes approach to the minimax property of the cusum procedure. Annals of Statistics, 24(4): 1804–1812, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032298296.

Bell:1984:SEN

[Bel84] William Bell. Signal extraction for nonstationary time series. Annals of Statistics, 12(2):646–664, June 1984. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346512. See

correction [Bel91].

Bell:1991:CSE

[Bel91] William Bell. Correction: “Signal Extraction for Nonstationary Time Series”. Annals of Statistics, 19(4):2280, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/

1176348401. See [Bel84].

Berger:1990:DAP

[Ber90] James Berger. Discussion: An ancillarity paradox which appears in multiple linear regression. Annals of Statistics, 18 (2):493–496, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347603.

Beran:1992:CCC

[Ber92] Rudolf Beran. Controlling conditional coverage probability in prediction. Annals of Statistics, 20(2):1110–1119, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348673.

Beran:1993:PCP

[Ber93] Rudolf Beran. Probability-centered prediction regions. Annals of Statistics, 21(4):1967–1981, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349405.

Besag:1994:DMC

[Bes94] Julian Besag. Discussion: Markov chains for exploring posterior distributions. Annals of Statistics, 22(4):1734–1741, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325752.

Betensky:1996:OFS

[Bet96] Rebecca A. Betensky. An O’Brien–Fleming sequential trial for comparing three treatments. Annals of Statistics, 24(4): 1765–1791, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032298294.

Boente:1990:ADR

[BF90] Graciela Boente and Ricardo Fraiman. Asymptotic distribution of robust estimators for nonparametric models from mixing processes. Annals of Statistics, 18(2):891–906, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347631.

Beran:1998:NCM

[BF98] Rudolf Beran and Nicholas I. Fisher. Nonparametric comparison of mean directions or mean axes. Annals of Statistics, 26 (2):472–493, April 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1028144845.

Beran:1996:NEI

[BFH96] Rudolf Beran, Andrey Feuerverger, and Peter Hall. On nonparametric estimation of intercept and slope distributions in random coefficient regression. Annals of Statistics, 24(6):2569– 2592, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032181170.

Bartlett:1998:BMN

[BFLS98] Peter Bartlett, Yoav Freund, Wee Sun Lee, and Robert E. Schapire. Boosting the margin: a new explanation for the effectiveness of voting methods. Annals of Statistics, 26(5): 1651–1686, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1024691352.

Bhattacharya:1990:KNN

[BG90a] P. K. Bhattacharya and Ashis K. Gangopadhyay. Kernel and nearest-neighbor estimation of a conditional quantile. Annals of Statistics, 18(3):1400–1415, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347757.

Bickel:1990:DLR

[BG90b] Peter J. Bickel and J. K. Ghosh. A decomposition for the likelihood ratio statistic and the Bartlett correction — a Bayesian argument. Annals of Statistics, 18(3):1070–1090, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347740.

Brown:1990:IIB

[BG90c] Lawrence D. Brown and Leslaw Gajek. Information inequalities for the Bayes risk. Annals of Statistics, 18(4):1578–1594, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347867.

Brown:1992:TSS

[BG92] Lawrence D. Brown and Eitan Greenshtein. Two-sided sequential tests. Annals of Statistics, 20(1):555–561, March 1992. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348539.

Bolthausen:1993:RCM

[BG93] Erwin Bolthausen and Friedrich Gotze. The rate of convergence for multivariate sampling statistics. Annals of Statistics, 21(4): 1692–1710, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349393.

Baddeley:1997:KME

[BG97] Adrian Baddeley and Richard D. Gill. Kaplan–Meier estimators of distance distributions for spatial point processes. Annals of Statistics, 25(1):263–292, February 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1034276629.

Borgan:1995:MAS

[BGL95] O. Borgan, L. Goldstein, and B. Langholz. Methods for the analysis of sampled cohort data in the Cox proportional hazards model. Annals of Statistics, 23(5):1749–1778, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176324322.

Bentkus:1997:EES

[BGvZ97] V. Bentkus, F. Götze, and W. R. van Zwet. An Edgeworth expansion for symmetric statistics. Annals of Statistics, 25 (2):851–896, April 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1031833676.

Bilias:1997:TGA

[BGY97] Yannis Bilias, Minggao Gu, and Zhiliang Ying. Towards a general asymptotic theory for Cox model with staggered entry. Annals of Statistics, 25(2):662–682, April 1997. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1031833668.

Barry:1990:OTD

[BH90a] Daniel Barry and J. A. Hartigan. An omnibus test for departures from constant mean. Annals of Statistics, 18(3):1340– 1357, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347753.

Bohm:1990:IBA

[BH90b] Walter Bohm and Peter Hackl. Improved bounds for the average run length of control charts based on finite weighted sums. Annals of Statistics, 18(4):1895–1899, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347887.

Baringhaus:1992:LDM

[BH92a] L. Baringhaus and N. Henze. Limit distributions for Mardia’s measure of multivariate skewness. Annals of Statistics, 20(4): 1889–1902, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348894.

Barry:1992:PPM

[BH92b] Daniel Barry and J. A. Hartigan. Product partition models for change point problems. Annals of Statistics, 20(1):260–279, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176348521.

Beran:1992:ECD

[BH92c] Rudolf Beran and Peter Hall. Estimating coefficient distributions in random coefficient regressions. Annals of Statistics, 20(4):1970–1984, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348898.

Booth:1993:BCR

[BH93a] James G. Booth and Peter Hall. Bootstrap confidence regions for functional relationships in errors-in-variables models. Annals of Statistics, 21(4):1780–1791, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349397.

Booth:1993:IJD

[BH93b] James G. Booth and Peter Hall. An improvement of the jackknife distribution function estimator. Annals of Statistics, 21(3):1476–1485, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176349268.

Barron:1998:ITS

[BH98] Andrew Barron and Nicolas Hengartner. Information theory and superefficiency. Annals of Statistics, 26(5):1800–1825, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1024691358.

Bai:1999:ADM

[BH99] Zhi-Dong Bai and Xuming He. Asymptotic distributions of the maximal depth estimators for regression and multivariate location. Annals of Statistics, 27(5):1616–1637, October 1999. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1017939144.

Brown:1997:UTB

[BHM97] Lawrence D. Brown, J. T. Gene Hwang, and Axel Munk. An unbiased test for the bioequivalence problem. Annals of Statistics, 25(6):2345–2367, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1030741076.

Booth:1992:BEC

[BHW92] James Booth, Peter Hall, and Andrew Wood. Bootstrap estimation of conditional distributions. Annals of Statistics, 20(3):1594–1610, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348786.

Booth:1993:BIR

[BHW93] James G. Booth, Peter Hall, and Andrew T. A. Wood. Balanced importance resampling for the bootstrap. Annals of Statistics, 21(1):286–298, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349026.

Birge:1997:EUD

[Bir97] Lucien Birgé. Estimation of unimodal densities without smoothness assumptions. Annals of Statistics, 25(3):970–981, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1069362733.

Bischoff:1998:FCL

[Bis98] Wolfgang Bischoff. A functional central limit theorem for regression models. Annals of Statistics, 26(4):1398–1410, August 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1024691248.

Brown:1991:IIB

[BL91] Lawrence D. Brown and Mark G. Low. Information inequality bounds on the minimax risk (with an application to nonparametric regression). Annals of Statistics, 19(1):329–337, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347985.

Begg:1992:EMT

[BL92] Melissa Dowd Begg and Stephen Lagakos. Effects of mismodeling on tests of association based on logistic regression models. Annals of Statistics, 20(4):1929–1952, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348896.

Brown:1996:AEN

[BL96a] Lawrence D. Brown and Mark G. Low. Asymptotic equivalence of nonparametric regression and white noise. Annals of Statistics, 24(6):2384–2398, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1032181159.

Brown:1996:CRI

[BL96b] Lawrence D. Brown and Mark G. Low. A constrained risk inequality with applications to nonparametric functional estimation. Annals of Statistics, 24(6):2524–2535, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032181166.

Butler:1997:CML

[BL97] Steven M. Butler and Thomas A. Louis. Consistency of maximum likelihood estimators in general random effects models for binary data. Annals of Statistics, 25(1):351–377, February 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1034276633.

Bar-Lev:1994:SMW

[BLEL94] Shaul K. Bar-Lev, Peter Enis, and Gerard Letac. Sampling models which admit a given general exponential family as a conjugate family of priors. Annals of Statistics, 22(3):1555– 1586, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325643.

Buja:1994:IPD

[BLRS94] Andreas Buja, B. F. Logan, J. A. Reeds, and L. A. Shepp. Inequalities and positive-definite functions arising from a problem in multidimensional scaling. Annals of Statistics, 22(1): 406–438, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325376.

Brown:1997:SNF

[BLZ97] Lawrence D. Brown, Mark G. Low, and Linda H. Zhao. Superefficiency in nonparametric function estimation. Annals of Statistics, 25(6):2607–2625, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1030741087.

Brown:1992:LAL

[BM92] Lawrence D. Brown and John I. Marden. Local admissibility and local unbiasedness in hypothesis testing problems. Annals of Statistics, 20(2):832–852, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348659.

Beran:1994:MDE

[BM94] R. Beran and P. W. Millar. Minimum distance estimation in random coefficient regression models. Annals of Statistics, 22(4):1976–1992, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325767.

Birge:1995:EIF

[BM95] Lucien Birge and Pascal Massart. Estimation of integral functionals of a density. Annals of Statistics, 23(1):11–29, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176324452.

Bedford:1997:CMD

[BM97] Tim Bedford and Isaac Meilijson. A characterization of marginal distributions of (possibly dependent) lifetime variables which right censor each other. Annals of Statistics, 25(4): 1622–1645, August 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1031594734.

Bunke:1998:ABB

[BM98] Olaf Bunke and Xavier Milhaud. Asymptotic behavior of Bayes estimates under possibly incorrect models. Annals of Statistics, 26(2):617–644, April 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1028144851.

Basawa:1991:BUF

[BMM^{+}91] I. V. Basawa, A. K. Mallik, W. P. McCormick, J. H. Reeves, and R. L. Taylor. Bootstrapping unstable first-order autoregressive processes. Annals of Statistics, 19(2):1098–1101, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348142.

Bucklew:1991:AOH

[BN91] J. A. Bucklew and P. E. Ney. Asymptotically optimal hypothesis testing with memory constraints. Annals of Statistics, 19 (2):982–998, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348132.

Barndorff-Nielsen:1993:OM

[BNB93] O. E. Barndorff-Nielsen and P. Blaesild. Orthogeodesic models. Annals of Statistics, 21(2):1018–1039, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349162.

Berk:1996:SCC

[BNO96] R. H. Berk, A. G. Nogales, and J. A. Oyola. Some counterexamples concerning sufficiency and invariance. Annals of Statistics, 24(2):902–905, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032894472.

Bickel:1992:NIU

[BNW92] Peter J. Bickel, Vijayan N. Nair, and Paul C. C. Wang. Nonparametric inference under biased sampling from a finite population. Annals of Statistics, 20(2):853–878, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348660.

Bose:1994:BRM

[Bos94] Sudip Bose. Bayesian robustness with mixture classes of priors. Annals of Statistics, 22(2):652–667, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325489.

Bosq:1997:PRN

[Bos97] Denis Bosq. Parametric rates of nonparametric estimators and predictors for continuous time processes. Annals of Statistics, 25(3):982–1000, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362734.

Bose:1998:BRE

[Bos98] Arup Bose. Bahadur representation of M_{m} estimates. Annals of Statistics, 26(2):771–777, April 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1028144859.

Bhansali:1991:CML

[BP91] R. J. Bhansali and F. Papangelou. Convergence of moments of least squares estimators for the coefficients of an autoregressive process of unknown order. Annals of Statistics, 19(3):1155– 1162, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348243.

Burshtein:1992:MIP

[BPKN92] David Burshtein, Vincent Della Pietra, Dimitri Kanevsky, and Arthur Nadas. Minimum impurity partitions. Annals of Statistics, 20(3):1637–1646, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348789.

Babu:1999:SOC

[BPR99] G. Jogesh Babu, P. K. Pathak, and C. R. Rao. Second-order correctness of the Poisson bootstrap. Annals of Statistics, 27(5): 1666–1683, October 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1017939146.

Berger:1990:SHB

[BR90] James O. Berger and Christian Robert. Subjective hierarchical Bayes estimation of a multivariate normal mean: On the frequentist interface. Annals of Statistics, 18(2):617–651, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347619.

Bai:1991:EEF

[BR91a] Z. D. Bai and C. Radhakrishna Rao. Edgeworth expansion of a function of sample means. Annals of Statistics, 19(3):1295– 1315, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348250.

Bickel:1991:LST

[BR91b] Peter J. Bickel and J. Ritov. Large sample theory of estimation in biased sampling regression models. I. Annals of Statistics, 19(2):797–816, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348121.

Berti:1994:CII

[BR94] Patrizia Berti and Pietro Rigo. Coherent inferences and improper priors. Annals of Statistics, 22(3):1177–1194, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325624.

Berti:1996:EIW

[BR96] Patrizia Berti and Pietro Rigo. On the existence of inferences which are consistent with a given model. Annals of Statistics, 24 (3):1235–1249, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032526966.

Breiman:1991:DMA

[Bre91] Leo Breiman. Discussion: Multivariate adaptive regression splines. Annals of Statistics, 19(1):82–91, March 1991. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347965.

Breiman:1996:HIS

[Bre96] Leo Breiman. Heuristics of instability and stabilization in model selection. Annals of Statistics, 24(6):2350–2383, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032181158.

Breiman:1998:ACD

[Bre98] Leo Breiman. Arcing classifier (with discussion and a rejoinder by the author). Annals of Statistics, 26(3):801–849, June 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1024691079.

Bhattacharyya:1997:AIN

[BRF97] B. B. Bhattacharyya, G. D. Richardson, and L. A. Franklin. Asymptotic inference for near unit roots in spatial autoregression. Annals of Statistics, 25(4):1709–1724, August 1997. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1031594738.

Brown:1990:APW

[Bro90a] Lawrence D. Brown. An ancillarity paradox which appears in multiple linear regression. Annals of Statistics, 18(2):471–493, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347602.

Brown:1990:RAP

[Bro90b] Lawrence D. Brown. Rejoinder: An ancillarity paradox which appears in multiple linear regression. Annals of Statistics, 18 (2):533–538, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347614.

Brooks:1998:MCD

[Bro98] Stephen P. Brooks. MCMC convergence diagnosis via multivariate bounds on log-concave densities. Annals of Statistics, 26(1): 398–433, February 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1030563991.

Berti:1991:CSI

[BRR91] Patrizia Berti, Eugenio Regazzini, and Pietro Rigo. Coherent statistical inference and Bayes theorem. Annals of Statistics, 19 (1):366–381, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347988.

Bickel:1998:ANM

[BRR98] Peter J. Bickel, Ya’acov Ritov, and Tobias Rydén. Asymptotic normality of the maximum-likelihood estimator for general hidden Markov models. Annals of Statistics, 26(4):1614–1635, August 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1024691255.

Brunk:1991:FCI

[Bru91] H. D. Brunk. Fully coherent inference. Annals of Statistics, 19(2):830–849, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348123.

Bickel:1991:EEL

[BRW91] Peter J. Bickel, Ya’acov Ritov, and Jon A. Wellner. Efficient estimation of linear functionals of a probability measure P with known marginal distributions. Annals of Statistics, 19(3):1316– 1346, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348251.

Barron:1991:ADF

[BS91a] Andrew R. Barron and Chyong-Hwa Sheu. Approximation of density functions by sequences of exponential families. Annals of Statistics, 19(3):1347–1369, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348252. See

correction [BS91b].

Barron:1991:CAD

[BS91b] Andrew R. Barron and Chyong-Hwa Sheu. Correction: “Approximation of Density Functions by Sequences of Exponential Families”. Annals of Statistics, 19(4):2284, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/

1176348405. See [BS91a].

Brandwein:1991:GJS

[BS91c] Ann Cohen Brandwein and William E. Strawderman. Generalizations of James–Stein estimators under spherical symmetry. Annals of Statistics, 19(3):1639–1650, September 1991. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348267.

Berger:1996:CHP

[BS96] James O. Berger and William E. Strawderman. Choice of hierarchical priors: admissibility in estimation of normal means. Annals of Statistics, 24(3):931–951, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1032526950.

Bouckaert:1998:CGM

[BS98] Remco R. Bouckaert and Milan Studený. On chain graph models for description of conditional independence structures. Annals of Statistics, 26(4):1434–1495, August 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1024691250.

Bingham:1999:STR

[BS99] D. R. Bingham and R. R. Sitter. Some theoretical results for fractional factorial split-plot designs. Annals of Statistics, 27(4):1240–1255, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1017938924.

Birkes:1999:BOD

[BSV99] David S. Birkes, Justus F. Seely, and Dawn M. Vanleeuwen. Balance and orthogonality in designs for mixed classification models. Annals of Statistics, 27(6):1927–1947, December 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1017939245.

Barron:1999:CPD

[BSW99] Andrew Barron, Mark J. Schervish, and Larry Wasserman. The consistency of posterior distributions in nonparametric problems. Annals of Statistics, 27(2):536–561, April 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1018031206.

Breiman:1993:SEP

[BTZ93] Leo Breiman, Yacov Tsur, and Amos Zemel. On a simple estimation procedure for censored regression models with known error distributions. Annals of Statistics, 21(4):1711–1720, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349394.

Buhlmann:1994:BBE

[Buh94] Peter Buhlmann. Blockwise bootstrapped empirical process for stationary sequences. Annals of Statistics, 22(2):995–1012, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176325508.

Buhlmann:1998:SBS

[Büh98] Peter Bühlmann. Sieve bootstrap for smoothing in nonstationary time series. Annals of Statistics, 26(1):48–83, February 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1030563978.

Buja:1990:RFC

[Buj90] Andreas Buja. Remarks on functional canonical variates, alternating least squares methods and ACE. Annals of Statistics, 18(3):1032–1069, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347739.

Buja:1994:DUP

[Buj94] Andreas Buja. Discussion: The use of polynomial splines and their tensor products in multivariate function estimation. Annals of Statistics, 22(1):171–177, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325362.

Burman:1991:RCE

[Bur91] Prabir Burman. Rates of convergence for the estimates of the optimal transformations of variables. Annals of Statistics, 19 (2):702–723, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348116.

Buhlmann:1999:VLM

[BW99] Peter Bühlmann and Abraham J. Wyner. Variable length Markov chains. Annals of Statistics, 27(2):480–513, April 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1018031204.

Barron:1991:DMA

[BX91] Andrew R. Barron and Xiangyu Xiao. Discussion: Multivariate adaptive regression splines. Annals of Statistics, 19(1):67–82, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347964.

Bednarski:1996:REP

[BZ96] Tadeusz Bednarski and Stefan Zontek. Robust estimation of parameters in a mixed unbalanced model. Annals of Statistics, 24(4):1493–1510, August 1996. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1032298279.

Bhattacharya:1997:SIP

[BZ97] P. K. Bhattacharya and Peng-Liang Zhao. Semiparametric inference in a partial linear model. Annals of Statistics, 25(1): 244–262, February 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1034276628.

Brown:1998:ANN

[BZ98] Lawrence D. Brown and Cun-Hui Zhang. Asymptotic nonequivalence of nonparametric experiments when the smoothness index is 1/2. Annals of Statistics, 26(1):279–287, February 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1030563986.

Cao-Abad:1991:RCW

[CA91] R. Cao-Abad. Rate of convergence for the wild bootstrap in nonparametric regression. Annals of Statistics, 19(4):2226– 2231, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348394.

Cabana:1996:TEM

[Cab96] Alejandra Cabaña. Transformations of the empirical measure and Kolmogorov–Smirnov tests. Annals of Statistics, 24(5): 2020–2035, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362307.

Cuesta-Albertos:1997:TMA

[CAGM97] J. A. Cuesta-Albertos, A. Gordaliza, and C. Matrán. Trimmed k-means: an attempt to robustify quantizers. Annals of Statistics, 25(2):553–576, April 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1031833664.

Cai:1999:AWE

[Cai99] T. Tony Cai. Adaptive wavelet estimation: a block thresholding and oracle inequality approach. Annals of Statistics, 27 (3):898–924, June 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1018031262.

Cuesta-Albertos:1999:TGF

[CAMRRdB99] Juan A. Cuesta-Albertos, Carlos Matrán, Jesús M. Rodríguez-Rodríguez, and Eustasio del Barrio. Tests of goodness of fit based on the L_{2}-Wasserstein distance. Annals of Statistics, 27(4):1230–1239, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1017938923.

Casella:1990:DAP

[Cas90] George Casella. Discussion: An ancillarity paradox which appears in multiple linear regression. Annals of Statistics, 18 (2):497–500, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347604.

Casalis:1996:SQN

[Cas96] M. Casalis. The 2d + 4 simple quadratic natural exponential families on R^{d}. Annals of Statistics, 24(4):1828–1854, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032298298.

Cheng:1991:OST

[CB91] C.-S. Cheng and R. A. Bailey. Optimality of some twoassociate-class partially balanced incomplete-block designs. Annals of Statistics, 19(3):1667–1671, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348270.

Cesa-Bianchi:1999:PIS

[CBL99] Nicolò Cesa-Bianchi and Gábor Lugosi. On prediction of individual sequences. Annals of Statistics, 27(6):1865–1895, December 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1017939242.

Cabana:1994:GFC

[CC94] Alejandra Cabana and Enrique M. Cabana. Goodness-of-fit and comparison tests of the Kolmogorov–Smirnov type for bivariate populations. Annals of Statistics, 22(3):1447–1459, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325636.

Cabana:1997:TEP

[CC97] A. Cabaña and E. M. Cabaña. Transformed empirical processes and modified Kolmogorov–Smirnov tests for multivariate distributions. Annals of Statistics, 25(6):2388–2409, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1030741078.

Chen:1999:TOB

[CC99] Hegang Chen and Ching-Shui Cheng. Theory of optimal blocking of 2^{n−m} designs. Annals of Statistics, 27(6):1948–1973, December 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1017939246.

Chen:1995:SUC

[CCL95] Kani Chen, Min-Te Chao, and Shaw-Hwa Lo. On strong uniform consistency of the Lynden–Bell estimator for truncated data. Annals of Statistics, 23(2):440–449, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176324529.

Cifarelli:1996:ADG

[CCR96] Donato Michele Cifarelli, Pier Luigi Conti, and Eugenio Regazzini. On the asymptotic distribution of a general measure of monotone dependence. Annals of Statistics, 24(3):1386–1399, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1032526975.

Calvin:1991:MLE

[CD91] James A. Calvin and Richard L. Dykstra. Maximum likelihood estimation of a set of covariance matrices under lowner order restrictions with applications to balanced multivariate variance components models. Annals of Statistics, 19(2):850–869, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348124.

Chaudhuri:1997:ADQ

[CDS97] Probal Chaudhuri, Kjell Doksum, and Alexander Samarov. On average derivative quantile regression. Annals of Statistics, 25 (2):715–744, April 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1031833670.

Chang:1990:RPO

[CE90] Ted Chang and David Eaves. Reference priors for the orbit in a group model. Annals of Statistics, 18(4):1595–1614, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347868.

Cuevas:1997:PAS

[CF97] Antonio Cuevas and Ricardo Fraiman. A plug-in approach to support estimation. Annals of Statistics, 25(6):2300–2312, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1030741073.

Cabrera:1999:TEB

[CF99] Javier Cabrera and Luisa Turrin Fernholz. Target estimation for bias and mean square error reduction. Annals of Statistics, 27(3):1080–1104, June 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1018031269.

Cheng:1997:ABC

[CFM97] Ming-Yen Cheng, Jianqing Fan, and J. S. Marron. On automatic boundary corrections. Annals of Statistics, 25(4): 1691–1708, August 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1031594737.

Comets:1991:AML

[CG91] Francis Comets and Basilis Gidas. Asymptotics of maximum likelihood estimators for the Curie–Weiss model. Annals of Statistics, 19(2):557–578, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348111.

Chan:1994:DMC

[CG94] Kung Sik Chan and Charles J. Geyer. Discussion: Markov chains for exploring posterior distributions. Annals of Statistics, 22(4):1747–1758, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325754.

Clarke:1995:PCG

[CG95] B. Clarke and J. K. Ghosh. Posterior convergence given the mean. Annals of Statistics, 23(6):2116–2144, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1034713650.

Csorgo:1991:CLT

[CGH91] Miklós Csörgő, Edit Gombay, and Lajos Horváth. Central limit theorems for L_{p} distances of kernel estimators of densities under random censorship. Annals of Statistics, 19(4):1813–1831, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176348372.

Carroll:1998:LLR

[CGWW98] R. J. Carroll, Roberto G. Gutierrez, C. Y. Wang, and Suojin Wang. Local linear regression for generalized linear models with missing data. Annals of Statistics, 26(3):1028–1050, June 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1024691087.

Carroll:1990:NEO

[CH90a] R. J. Carroll and Peter Hall. Nonparametric estimation of optimal performance criteria in quality engineering. Annals of Statistics, 18(1):281–302, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347501.

Chang:1990:TSP

[CH90b] I-Shou Chang and Chao A. Hsiung. Time-sequential point estimation through estimating equations. Annals of Statistics, 18(3):1378–1388, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347755.

Chang:1991:APP

[CH91] I.-Shou Chang and Chao A. Hsiung. An E-ancillarity projection property of Cox’s partial score function. Annals of Statistics, 19(3):1651–1660, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348268.

Chen:1993:SEL

[CH93] Song Xi Chen and Peter Hall. Smoothed empirical likelihood confidence intervals for quantiles. Annals of Statistics, 21(3):1166–1181, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176349256.

Chang:1994:ACM

[CH94a] I-Shou Chang and Chao A. Hsiung. Asymptotic consistency of the maximum likelihood estimate in positron emission tomography and applications. Annals of Statistics, 22(4):1871– 1883, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325761.

Chang:1994:IAE

[CH94b] I-Shou Chang and Chao A. Hsiung. Information and asymptotic efficiency in some generalized proportional hazards models for counting processes. Annals of Statistics, 22(3):1275– 1298, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325629.

Csorgo:1995:DBS

[CH95] Miklós Csörgő and Lajos Horváth. On the distance between smoothed empirical and quantile processes. Annals of Statistics, 23(1):113–131, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324458.

Chen:1996:DWM

[CH96] Hegang Chen and A. S. Hedayat. 2^{n−l} designs with weak minimum aberration. Annals of Statistics, 24(6):2536–2548, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032181167.

Cheng:1999:MTD

[CH99] Ming-Yen Cheng and Peter Hall. Mode testing in difficult cases. Annals of Statistics, 27(4):1294–1315, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1017938927.

Chan:1990:TTA

[Cha90a] K. S. Chan. Testing for threshold autoregression. Annals of Statistics, 18(4):1886–1894, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347886.

Chang:1990:WCS

[Cha90b] Myron N. Chang. Weak convergence of a self-consistent estimator of the survival function with doubly censored data. Annals of Statistics, 18(1):391–404, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347506.

Chaudhuri:1991:NER

[Cha91] Probal Chaudhuri. Nonparametric estimates of regression quantiles and their local Bahadur representation. Annals of Statistics, 19(2):760–777, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348119.

Chaudhuri:1992:MLE

[Cha92] Probal Chaudhuri. Multivariate location estimation using extension of R-estimates through U-statistics type approach. Annals of Statistics, 20(2):897–916, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348662.

Chan:1993:CLD

[Cha93] K. S. Chan. Consistency and limiting distribution of the least squares estimator of a threshold autoregressive model. Annals of Statistics, 21(1):520–533, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349040.

Chattopadhyay:1994:TAD

[Cha94] Manas K. Chattopadhyay. Two-armed Dirichlet bandits with discounting. Annals of Statistics, 22(3):1212–1221, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325626.

Chanda:1996:APE

[Cha96] Kamal C. Chanda. Asymptotic properties of estimators for autoregressive models with errors in variables. Annals of Statistics, 24(1):423–430, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1033066218.

Cheng:1990:MSC

[Che90a] Qiansheng Cheng. Maximum standardized cumulant deconvolution of non-Gaussian linear processes. Annals of Statistics, 18(4):1774–1783, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347877.

Chentsov:1990:UIK

[Che90b] Nikolai N. Chentsov. The unfathomable influence of Kolmogorov. Annals of Statistics, 18(3):987–998, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347735.

Chen:1991:EPP

[Che91a] Hung Chen. Estimation of a projection-pursuit type regression model. Annals of Statistics, 19(1):142–157, March 1991. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347974.

Chen:1991:ISM

[Che91b] Zehua Chen. Interaction spline models and their convergence rates. Annals of Statistics, 19(4):1855–1868, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348374.

Chen:1992:SRF

[Che92a] Jiahua Chen. Some results on 2^{n−k} fractional factorial designs and search for minimum aberration designs. Annals of Statistics, 20(4):2124–2141, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348907.

Cheng:1992:URN

[Che92b] Qiansheng Cheng. On the unique representation of non-Gaussian linear processes. Annals of Statistics, 20(2):1143– 1145, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1176348677.

Chen:1995:TFT

[Che95a] Hanfeng Chen. Tests following transformations. Annals of Statistics, 23(5):1587–1593, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324314.

Chen:1995:AEE

[Che95b] Hung Chen. Asymptotically efficient estimation in semiparametric generalized linear models. Annals of Statistics, 23(4): 1102–1129, August 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324700.

Chen:1995:ORC

[Che95c] Jiahua Chen. Optimal rate of convergence for finite mixture models. Annals of Statistics, 23(1):221–233, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176324464.

Cheng:1995:UCG

[Che95d] C. Cheng. Uniform consistency of generalized kernel estimators of quantile density. Annals of Statistics, 23(6):2285–2291, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1034713657.

Cheng:1995:CCR

[Che95e] Ching-Shui Cheng. Complete class results for the moment matrices of designs over permutation-invariant sets. Annals of Statistics, 23(1):41–54, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324454.

Cheng:1995:SPP

[Che95f] Ching-Shui Cheng. Some projection properties of orthogonal arrays. Annals of Statistics, 23(4):1223–1233, August 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176324706.

Cheng:1997:BSL

[Che97] Ming-Yen Cheng. A bandwidth selector for local linear density estimators. Annals of Statistics, 25(3):1001–1013, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1069362735.

Chen:1998:WPM

[Che98] Sean X. Chen. Weighted polynomial models and weighted sampling schemes for finite population. Annals of Statistics, 26(5): 1894–1915, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1024691362.

Chiu:1990:ADB

[Chi90] Shean-Tsong Chiu. On the asymptotic distributions of bandwidth estimates. Annals of Statistics, 18(4):1696–1711, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347873.

Chiu:1991:BSK

[Chi91a] Shean-Tsong Chiu. Bandwidth selection for kernel density estimation. Annals of Statistics, 19(4):1883–1905, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348376.

Chiu:1991:SSB

[Chi91b] Shean-Tsong Chiu. Some stabilized bandwidth selectors for nonparametric regression. Annals of Statistics, 19(3):1528– 1546, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348260.

Chou:1991:SED

[Cho91] Jine-Phone Chou. Simultaneous estimation in discrete multivariate exponential families. Annals of Statistics, 19(1): 314–328, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347984.

Choudhuri:1998:BBC

[Cho98] Nidhan Choudhuri. Bayesian bootstrap credible sets for multidimensional mean functional. Annals of Statistics, 26(6):2104– 2127, December 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1024691463.

Chan:1995:PBE

[CHP95] Grace Chan, Peter Hall, and D. S. Poskitt. Periodogram-based estimators of fractal properties. Annals of Statistics, 23(5): 1684–1711, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324319.

Choi:1996:AUM

[CHS96] Sungsub Choi, W. J. Hall, and Anton Schick. Asymptotically uniformly most powerful tests in parametric and semiparametric models. Annals of Statistics, 24(2):841–861, April 1996. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1032894469.

Choy:1999:LAN

[CHST99] Kokyo Choy, Marc Hallin, Abdeslam Serroukh, and Masanobu Taniguchi. Local asymptotic normality for regression models with long-memory disturbance. Annals of Statistics, 27(6): 2054–2080, December 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1017939250.

Chen:1999:SCM

[CHY99] Kani Chen, Inchi Hu, and Zhiliang Ying. Strong consistency of maximum quasi-likelihood estimators in generalized linear models with fixed and adaptive designs. Annals of Statistics, 27(4):1155–1163, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1017938919.

Chang:1992:FSC

[cICM92] Yuan chin Ivan Chang and Adam T. Martinsek. Fixed size confidence regions for parameters of a logistic regression model. Annals of Statistics, 20(4):1953–1969, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348897.

Chang:1995:ERB

[CK95] Ted Chang and Daijin Ko. M-estimates of rigid body motion on the sphere and in Euclidean space. Annals of Statistics, 23(5): 1823–1847, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324325.

Cohen:1994:PTO

[CKS94a] Arthur Cohen, J. H. B. Kemperman, and H. B. Sackrowitz. Projected tests for order restricted alternatives. Annals of Statistics, 22(3):1539–1546, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325641.

Cohen:1994:UTE

[CKS94b] Arthur Cohen, J. H. B. Kemperman, and H. B. Sackrowitz. Unbiased testing in exponential family regression. Annals of Statistics, 22(4):1931–1946, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176325765.

Carroll:1995:DRS

[CKW95] R. J. Carroll, R. K. Knickerbocker, and C. Y. Wang. Dimension reduction in a semiparametric regression model with errors in covariates. Annals of Statistics, 23(1):161–181, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176324461.

Cox:1991:MLT

[CL91] Dennis D. Cox and Isabel Llatas. Maximum likelihood type estimation for nearly nonstationary autoregressive time series. Annals of Statistics, 19(3):1109–1128, September 1991. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348240.

Chen:1992:BAT

[CL92] Hanfeng Chen and Wei-Yin Loh. Bounds on AREs of tests following Box–Cox transformations. Annals of Statistics, 20 (3):1485–1500, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348780.

Casalis:1996:LOR

[CL96a] M. Casalis and G. Letac. The Lukacs–Olkin–Rubin characterization of Wishart distributions on symmetric cones. Annals of Statistics, 24(2):763–786, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1032894464.

Chen:1996:BAC

[CL96b] Kani Chen and Shaw-Hwa Lo. On bootstrap accuracy with censored data. Annals of Statistics, 24(2):569–595, April 1996. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1032894453.

Chen:1997:RUC

[CL97] Kani Chen and Shaw-Hwa Lo. On the rate of uniform convergence of the product-limit estimator: strong and weak laws. Annals of Statistics, 25(3):1050–1087, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1069362738.

Chen:1999:DRC

[CLW99] Chun-Houh Chen, Ker-Chau Li, and Jane-Ling Wang. Dimension reduction for censored regression data. Annals of Statistics, 27(1):1–23, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1018031098.

Chu:1991:CTB

[CM91] C.-K. Chu and J. S. Marron. Comparison of two bandwidth selectors with dependent errors. Annals of Statistics, 19(4): 1906–1918, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348377.

Chai:1993:YBC

[CM93] Feng-Shun Chai and Dibyen Majumdar. On the Yeh–Bradley conjecture on linear trend-free block designs. Annals of Statistics, 21(4):2087–2097, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176349411.

Csorgo:1995:DEU

[CM95a] Sandor Csorgo and Jan Mielniczuk. Density estimation under long-range dependence. Annals of Statistics, 23(3):990–999, June 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176324632.

Csorgo:1995:NRU

[CM95b] Sandor Csorgo and Jan Mielniczuk. Nonparametric regression under long-range dependent normal errors. Annals of Statistics, 23(3):1000–1014, June 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176324633.

Cheng:1998:RFF

[CM98] Ching-Shui Cheng and Rahul Mukerjee. Regular fractional factorial designs with minimum aberration and maximum estimation capacity. Annals of Statistics, 26(6):2289–2300, December 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1024691471.

Chiou:1999:NQL

[CM99] Jeng-Min Chiou and Hans-Georg Müller. Nonparametric quasi-likelihood. Annals of Statistics, 27(1):36–64, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1018031100.

Critchley:1993:PPG

[CMS93] Frank Critchley, Paul Marriott, and Mark Salmon. Preferred point geometry and statistical manifolds. Annals of Statistics, 21(3):1197–1224, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176349258.

Critchley:1994:PPG

[CMS94] Frank Critchley, Paul Marriott, and Mark Salmon. Preferred point geometry and the local differential geometry of the Kullback–Leibler divergence. Annals of Statistics, 22(3):1587– 1602, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325644.

Cheng:1998:TLF

[CMT98] Ching-Shui Cheng, R. J. Martin, and Boxin Tang. Two-level factorial designs with extreme numbers of level changes. Annals of Statistics, 26(4):1522–1539, August 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691252.

Cox:1990:AAP

[CO90] Dennis D. Cox and Finbarr O’Sullivan. Asymptotic analysis of penalized likelihood and related estimators. Annals of Statistics, 18(4):1676–1695, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347872.

Comets:1992:CCE

[Com92] Francis Comets. On consistency of a class of estimators for exponential families of Markov random fields on the lattice. Annals of Statistics, 20(1):455–468, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348532.

Conti:1999:LSB

[Con99] Pier Luigi Conti. Large sample Bayesian analysis for Geo/G/1 discrete-time queueing models. Annals of Statistics, 27(6): 1785–1807, December 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1017939239.

Copas:1990:DAP

[Cop90] J. B. Copas. Discussion: An ancillarity paradox which appears in multiple linear regression. Annals of Statistics, 18 (2):500–502, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347605.

Coster:1993:TFR

[Cos93] Daniel C. Coster. Trend-free run orders of mixed-level fractional factorial designs. Annals of Statistics, 21(4):2072–2086, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349410.

Cox:1991:GLE

[Cox91] Dennis D. Cox. Gaussian likelihood estimation for nearly nonstationary AR(1) processes. Annals of Statistics, 19(3):1129– 1142, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348241.

Cox:1993:ABI

[Cox93] Dennis D. Cox. An analysis of Bayesian inference for nonparametric regression. Annals of Statistics, 21(2):903–923, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349157.

Chan:1998:SSM

[CP98] Ngai Hang Chan and Wilfredo Palma. State space modeling of long-memory processes. Annals of Statistics, 26(2):719–740, April 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1028144856.

Cifarelli:1990:DFM

[CR90] Donato Michele Cifarelli and Eugenio Regazzini. Distribution functions of means of a Dirichlet process. Annals of Statistics, 18(1):429–442, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347509. See correction [CR94].

Cifarelli:1994:CDF

[CR94] Donato Michele Cifarelli and Eugenio Regazzini. Correction: “Distribution Functions of Means of a Dirichlet Process”. Annals of Statistics, 22(3):1633–1634, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325647. See

[CR90].

Chen:1993:IWL

[CS93a] Jiahua Chen and Jun Shao. Iterative weighted least squares estimators. Annals of Statistics, 21(2):1071–1092, June 1993. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349165.

Cohen:1993:IST

[CS93b] Arthur Cohen and H. B. Sackrowitz. Inadmissibility of studentized tests for normal order restricted models. Annals of Statistics, 21(2):746–752, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176349148.

Chen:1994:DDE

[CS94] Hung Chen and Jyh-Jen Horng Shiau. Data-driven efficient estimators for a partially linear model. Annals of Statistics, 22 (1):211–237, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325366.

Cohen:1996:COA

[CS96] Arthur Cohen and H. B. Sackrowitz. Cone order association and stochastic cone ordering with applications to orderrestricted testing. Annals of Statistics, 24(5):2036–2048, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1069362308.

Chen:1997:MCM

[CS97] Ming-Hui Chen and Qi-Man Shao. On Monte Carlo methods for estimating ratios of normalizing constants. Annals of Statistics, 25(4):1563–1594, August 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1031594732.

Cohen:1998:DTO

[CS98] Arthur Cohen and Harold B. Sackrowitz. Directional tests for one-sided alternatives in multivariate models. Annals of Statistics, 26(6):2321–2338, December 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691473.

Csiszar:1991:WLS

[Csi91] Imre Csiszar. Why least squares and maximum entropy? An axiomatic approach to inference for linear inverse problems. Annals of Statistics, 19(4):2032–2066, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348385.

Csorgo:1996:UGA

[Csö96] Sándor Csörgő. Universal Gaussian approximations under random censorship. Annals of Statistics, 24(6):2744–2778, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032181178.

Chan:1995:IUL

[CT95] Ngai Hang Chan and Norma Terrin. Inference for unstable long-memory processes with applications to fractional unit root autoregressions. Annals of Statistics, 23(5):1662–1683, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176324318.

Cuzick:1988:RR

[Cuz88] Jack Cuzick. Rank regression. Annals of Statistics, 16(4): 1369–1389, December 1988. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176351044. See correction [Cuz90].

Cuzick:1990:CRR

[Cuz90] Jack Cuzick. Correction: “Rank Regression”. Annals of Statistics, 18(1):469, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347513. See [Cuz88].

Cuzick:1992:EES

[Cuz92] Jack Cuzick. Efficient estimates in semiparametric additive regression models with unknown error distribution. Annals of Statistics, 20(2):1129–1136, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348675.

Cheng:1992:GEE

[CV92] Chi-Lun Cheng and John W. Van Ness. Generalized Mestimators for errors-in-variables regression. Annals of Statistics, 20(1):385–397, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348528.

Chen:1991:SRF

[CW91] Jiahua Chen and C. F. J. Wu. Some results on s^{n−k} fractional factorial designs with minimum aberration or optimal moments. Annals of Statistics, 19(2):1028–1041, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348135.

Cao:1999:DLS

[CW99] Jin Cao and Keith J. Worsley. The detection of local shape changes via the geometry of Hotelling’s T^{2} fields. Annals of Statistics, 27(3):925–942, June 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1018031263.

Chen:1996:CCC

[CY96] Kani Chen and Zhiliang Ying. A counterexample to a conjecture concerning the Hall–Wellner band. Annals of Statistics, 24(2):641–646, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032894456.

Csorgo:1996:MRL

[CZ96] Miklós Csörgő and Ričardas Zitikis. Mean residual life processes. Annals of Statistics, 24(4):1717–1739, August 1996. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1032298292.

Dabrowska:1997:SCR

[Dab97] Dorota M. Dabrowska. Smoothed Cox regression. Annals of Statistics, 25(4):1510–1540, August 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1031594730.

Dahlhaus:1995:ELR

[Dah95] Rainer Dahlhaus. Efficient location and regression estimation for long range dependent regression models. Annals of Statistics, 23(3):1029–1047, June 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176324635.

Dahlhaus:1997:FTS

[Dah97] R. Dahlhaus. Fitting time series models to nonstationary processes. Annals of Statistics, 25(1):1–37, February 1997. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1034276620.

DasGupta:1991:DVM

[Das91] Anirban DasGupta. Diameter and volume minimizing confidence sets in Bayes and classical problems. Annals of Statistics, 19(3):1225–1243, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348246.

Datta:1991:AOB

[Dat91a] Somnath Datta. Asymptotic optimality of Bayes compound estimators in compact exponential families. Annals of Statistics, 19(1):354–365, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347987.

Datta:1991:CPM

[Dat91b] Somnath Datta. On the consistency of posterior mixtures and its applications. Annals of Statistics, 19(1):338–353, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347986.

Datta:1992:SNB

[Dat92] Somnath Datta. Some nonasymptotic bounds for L_{1} density estimation using kernels. Annals of Statistics, 20(3):1658– 1667, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348791.

Davies:1990:AEL

[Dav90] Laurie Davies. The asymptotics of S-estimators in the linear regression model. Annals of Statistics, 18(4):1651–1675, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347871.

Davies:1992:ARM

[Dav92] Laurie Davies. The asymptotics of Rousseeuw’s minimum volume ellipsoid estimator. Annals of Statistics, 20(4):1828– 1843, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348891.

Davies:1993:ARL

[Dav93] P. L. Davies. Aspects of robust linear regression. Annals of Statistics, 21(4):1843–1899, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349401.

Davies:1998:LUL

[Dav98] P. L. Davies. On locally uniformly linearizable high breakdown location and scale functionals. Annals of Statistics, 26 (3):1103–1125, June 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1024691090.

Delampady:1990:LBB

[DB90] Mohan Delampady and James O. Berger. Lower bounds on Bayes factors for multinomial distributions, with application to chi-squared tests of fit. Annals of Statistics, 18(3):1295– 1316, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347750.

Donnell:1994:AAD

[DBS94a] Deborah J. Donnell, Andreas Buja, and Werner Stuetzle. Analysis of additive dependencies and concurvities using smallest additive principal components. Annals of Statistics, 22(4):

1635–1668, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325746.

Donnell:1994:RAA

[DBS94b] Deborah J. Donnell, Andreas Buja, and Werner Stuetzle. Rejoinder: Analysis of additive dependencies and concurvities using smallest additive principal components. Annals of Statistics, 22(4):1671–1673, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325748.

Dacunha-Castelle:1999:TOM

[DCG99] D. Dacunha-Castelle and E. Gassiat. Testing the order of a model using locally conic parametrization: population mixtures and stationary ARMA processes. Annals of Statistics, 27(4): 1178–1209, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1017938921.

Dudley:1990:KSC

[DCLP90] R. M. Dudley, Stephanie L. Cook, Jimena Llopis, and N. P. Peng. A. N. Kolmogorov and statistics: A citation bibliography. Annals of Statistics, 18(3):1017–1031, September 1990. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347738.

DeJong:1991:DKF

[De 91] Piet De Jong. The diffuse Kalman filter. Annals of Statistics, 19 (2):1073–1083, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348139.

Dykstra:1998:MLE

[DE98] Richard Dykstra and Hammou El Barmi. Maximum likelihood estimates via duality for log-convex models when cell probabilities are subject to convex constraints. Annals of Statistics, 26(5):1878–1893, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691361.

Degerine:1990:CPA

[Deg90] Serge Degerine. Canonical partial autocorrelation function of a multivariate time series. Annals of Statistics, 18(2):961–971, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347635.

Dette:1990:GOD

[Det90] H. Dette. A generalization of D- and D_{1}-optimal designs in polynomial regression. Annals of Statistics, 18(4):1784– 1804, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347878.

Dette:1992:ODC

[Det92] H. Dette. Optimal designs for a class of polynomials of odd or even degree. Annals of Statistics, 20(1):238–259, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348520.

Dette:1993:ETO

[Det93a] Holger Dette. Elfving’s theorem for D-optimality. Annals of Statistics, 21(2):753–766, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176349149.

Dette:1993:NOD

[Det93b] Holger Dette. A note on E-optimal designs for weighted polynomial regression. Annals of Statistics, 21(2):767–771, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349150.

Dette:1994:DDP

[Det94] Holger Dette. Discrimination designs for polynomial regression on compact intervals. Annals of Statistics, 22(2):890–903, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325501.

Dette:1995:ODI

[Det95] Holger Dette. Optimal designs for identifying the degree of a polynomial regression. Annals of Statistics, 23(4):1248–1266, August 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176324708.

Dette:1996:NBC

[Det96] Holger Dette. A note on Bayesian c- and D-optimal designs in nonlinear regression models. Annals of Statistics, 24(3): 1225–1234, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032526965.

Dette:1999:CTF

[Det99] Holger Dette. A consistent test for the functional form of a regression based on a difference of variance estimators. Annals of Statistics, 27(3):1012–1040, June 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1018031266.

Devroye:1992:NUS

[Dev92] Luc Devroye. A note on the usefulness of superkernels in density estimation. Annals of Statistics, 20(4):2037–2056, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348901.

Devroye:1994:GDS

[Dev94] Luc Devroye. On good deterministic smoothing sequences for kernel density estimates. Annals of Statistics, 22(2):886–889, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176325500.

Diaconis:1990:UCB

[DF90] P. Diaconis and D. Freedman. On the uniform consistency of Bayes estimates for multinomial probabilities. Annals of Statistics, 18(3):1317–1327, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347751.

Diaconis:1993:NBR

[DF93] P. Diaconis and D. A. Freedman. Nonparametric binary regression: A Bayesian approach. Annals of Statistics, 21(4): 2108–2137, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349413.

Devroye:1990:NEP

[DG90] Luc Devroye and László Györfi. No empirical probability measure can converge in the total variation sense for all distributions. Annals of Statistics, 18(3):1496–1499, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347765.

Datta:1991:BPL

[DG91] Gauri Sankar Datta and Malay Ghosh. Bayesian prediction in linear models: Applications to small area estimation. Annals of Statistics, 19(4):1748–1770, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348369.

Donoho:1992:BPL

[DG92] David L. Donoho and Miriam Gasko. Breakdown properties of location estimates based on halfspace depth and projected outlyingness. Annals of Statistics, 20(4):1803–1827, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348890.

Datta:1996:INP

[DG96] Gauri Sankar Datta and Malay Ghosh. On the invariance of noninformative priors. Annals of Statistics, 24(1):141–159, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1033066203.

Diaz-Garcia:1997:PCH

[DGJ97] José A. Díaz-García and Ramón Gutiérrez Jáimez. Proof of the conjectures of H. Uhlig on the singular multivariate beta and the Jacobian of a certain matrix transformation. Annals of Statistics, 25(5):2018–2023, October 1997. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1069362383.

Devroye:1994:SUC

[DGKL94] Luc Devroye, László Györfi, Adam Krzyzak, and Gabor Lugosi. On the strong universal consistency of nearest neighbor regression function estimates. Annals of Statistics, 22(3):1371– 1385, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325633.

DasGupta:1995:NGM

[DGZ95] A. DasGupta, J. K. Ghosh, and M. M. Zen. A new general method for constructing confidence sets in arbitrary dimensions: With applications. Annals of Statistics, 23(4):1408–1432, August 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176324715.

Dorfman:1993:EFP

[DH93] Alan H. Dorfman and Peter Hall. Estimators of the finite population distribution function using nonparametric regression. Annals of Statistics, 21(3):1452–1475, September 1993. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349267.

Dette:1998:ODI

[DH98] Holger Dette and Gerd Haller. Optimal designs for the identification of the order of a Fourier regression. Annals of Statistics, 26(4):1496–1521, August 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691251.

Dhar:1991:MDE

[Dha91] Sunil K. Dhar. Minimum distance estimation in an additive effects outliers model. Annals of Statistics, 19(1):205–228, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347977.

Dette:1999:ODR

[DHI99] Holger Dette, Linda M. Haines, and Lorens Imhof. Optimal designs for rational models and weighted polynomial regression.

Annals of Statistics, 27(4):1272–1293, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1017938926.

Doss:1997:BNE

[DHL97] Hani Doss, Fred W. Huffer, and Kevin L. Lawson. Bayesian nonparametric estimation via Gibbs sampling for coherent systems with redundancy. Annals of Statistics, 25(3):1109–1139, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1069362740.

DiCiccio:1991:ELB

[DHR91] Thomas DiCiccio, Peter Hall, and Joseph Romano. Empirical likelihood is Bartlett-correctable. Annals of Statistics, 19(2): 1053–1061, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348137.

Dette:1995:MDL

[DHS95] H. Dette, B. Heiligers, and W. J. Studden. Minimax designs in linear regression models. Annals of Statistics, 23(1): 30–40, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324453.

Dorado:1997:NEG

[DHS97] Crisanto Dorado, Myles Hollander, and Jayaram Sethuraman. Nonparametric estimation for a general repair model. Annals of Statistics, 25(3):1140–1160, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1069362741.

Davis:1995:TCP

[DHY95] Richard A. Davis, Dawei Huang, and Yi-Ching Yao. Testing for a change in the parameter values and order of an autoregressive model. Annals of Statistics, 23(1):282–304, February 1995. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176324468.

Dinwoodie:1993:LDC

[Din93] I. H. Dinwoodie. Large deviations for censored data. Annals of Statistics, 21(3):1608–1620, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349274.

Dinwoodie:1995:SEF

[Din95] I. H. Dinwoodie. Stationary exponential families. Annals of Statistics, 23(1):327–337, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324470.

Dahlhaus:1996:FDB

[DJ96] R. Dahlhaus and D. Janas. A frequency domain bootstrap for ratio statistics in time series analysis. Annals of Statistics, 24(5):1934–1963, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1069362304.

Donoho:1998:MEW

[DJ98] David L. Donoho and Iain M. Johnstone. Minimax estimation via wavelet shrinkage. Annals of Statistics, 26(3):879–921, June 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1024691081.

Donoho:1996:DEW

[DJKP96] David L. Donoho, Iain M. Johnstone, Gérard Kerkyacharian, and Dominique Picard. Density estimation by wavelet thresholding. Annals of Statistics, 24(2):508–539, April 1996. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1032894451.

Dykstra:1991:SIU

[DKR91] Richard Dykstra, Subhash Kochar, and Tim Robertson. Statistical inference for uniform stochastic ordering in several populations. Annals of Statistics, 19(2):870–888, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348125.

Drost:1997:AET

[DKW97] Feike C. Drost, Chris A. J. Klaassen, and Bas J. M. Werker. Adaptive estimation in time-series models. Annals of Statistics, 25(2):786–817, April 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1031833674.

Doksum:1990:CRB

[DL90] Kjell A. Doksum and Albert Y. Lo. Consistent and robust Bayes procedures for location based on partial information. Annals of Statistics, 18(1):443–453, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347510.

Donoho:1991:GRCa

[DL91a] David L. Donoho and Richard C. Liu. Geometrizing rates of convergence, II. Annals of Statistics, 19(2):633–667, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348114.

Donoho:1991:GRCb

[DL91b] David L. Donoho and Richard C. Liu. Geometrizing rates of convergence, III. Annals of Statistics, 19(2):668–701, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348115.

Duan:1991:SRL

[DL91c] Naihua Duan and Ker-Chau Li. Slicing regression: A link-free regression method. Annals of Statistics, 19(2):505–530, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348109.

Donoho:1992:REO

[DL92] David L. Donoho and Mark G. Low. Renormalization exponents and optimal pointwise rates of convergence. Annals of Statistics, 20(2):944–970, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348665.

Dawid:1993:HML

[DL93] A. P. Dawid and S. L. Lauritzen. Hyper Markov laws in the statistical analysis of decomposable graphical models. Annals of Statistics, 21(3):1272–1317, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349260. See

correction [DL95].

Dawid:1995:CHM

[DL95] A. P. Dawid and S. L. Lauritzen. Correction: “Hyper Markov Laws in the Statistical Analysis of Decomposable Graphical Models”. Annals of Statistics, 23(5):1864, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/

1176324328. See [DL93].

Devroye:1996:UAS

[DL96] Luc Devroye and Gábor Lugosi. A universally acceptable smoothing factor for kernel density estimates. Annals of Statistics, 24(6):2499–2512, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1032181164.

Devroye:1997:NUS

[DL97] Luc Devroye and Gábor Lugosi. Nonasymptotic universal smoothing factors, kernel complexity and Yatracos classes. Annals of Statistics, 25(6):2626–2637, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1030741088.

Donoho:1990:MRH

[DLM90] David L. Donoho, Richard C. Liu, and Brenda MacGibbon. Minimax risk over hyperrectangles, and implications. Annals of Statistics, 18(3):1416–1437, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347758.

Delyon:1999:CSA

[DLM99] Bernard Delyon, Marc Lavielle, and Eric Moulines. Convergence of a stochastic approximation version of the EM algorithm. Annals of Statistics, 27(1):94–128, February 1999. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1018031103.

Dabrowski:1992:SAM

[DM92] Andre Robert Dabrowski and David McDonald. Statistical analysis of multiple ion channel data. Annals of Statistics, 20(3):1180–1202, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348765.

Davis:1998:SAH

[DM98a] Richard A. Davis and Thomas Mikosch. The sample autocorrelations of heavy-tailed processes with applications to ARCH. Annals of Statistics, 26(5):2049–2080, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1024691368.

Dette:1998:NCS

[DM98b] Holger Dette and Axel Munk. Nonparametric comparison of several regression functions: exact and asymptotic theory. Annals of Statistics, 26(6):2339–2368, December 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1024691474.

Dette:1998:VLR

[DM98c] Holger Dette and Axel Munk. Validation of linear regression models. Annals of Statistics, 26(2):778–800, April 1998. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1028144860.

Dabrowski:1990:RTP

[DMR90] Andre Robert Dabrowski, David McDonald, and Uwe Rosler. Renewal theory properties of ion channels. Annals of Statistics, 18(3):1091–1115, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347741.

Dauxois:1998:NCA

[DN98] Jacques Dauxois and Guy Martial Nkiet. Nonlinear canonical analysis and independence tests. Annals of Statistics, 26(4): 1254–1278, August 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1024691242.

Donoho:1994:SEO

[Don94] David L. Donoho. Statistical estimation and optimal recovery. Annals of Statistics, 22(1):238–270, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325367.

Donoho:1997:CBO

[Don97] David L. Donoho. CART and best-ortho-basis: a connection. Annals of Statistics, 25(5):1870–1911, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1069362377.

Donoho:1999:WNM

[Don99] David L. Donoho. Wedgelets: nearly minimax estimation of edges. Annals of Statistics, 27(3):859–897, June 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1018031261.

Doss:1994:BNE

[Dos94a] Hani Doss. Bayesian nonparametric estimation for incomplete data via successive substitution sampling. Annals of Statistics, 22(4):1763–1786, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176325756.

Doss:1994:DMC

[Dos94b] Hani Doss. Discussion: Markov chains for exploring posterior distributions. Annals of Statistics, 22(4):1728–1734, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325751.

Dembo:1994:TCH

[DP94] Amir Dembo and Yuval Peres. A topological criterion for hypothesis testing. Annals of Statistics, 22(1):106–117, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325360.

Dette:1997:ODD

[DR97] Holger Dette and Ingo Röder. Optimal discrimination designs for multifactor experiments. Annals of Statistics, 25(3): 1161–1175, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362742. See correction [DR98].

Dette:1998:COD

[DR98] Holger Dette and Ingo Röder. Correction: “Optimal discrimination designs for multifactor experiments”. Annals of Statistics, 26(1):449, February 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1030563993. See [DR97].

Drees:1995:RPE

[Dre95] Holger Drees. Refined Pickands estimators of the extreme value index. Annals of Statistics, 23(6):2059–2080, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1034713647.

Drees:1998:ORC

[Dre98] Holger Drees. Optimal rates of convergence for estimates of the extreme value index. Annals of Statistics, 26(1):434–448, February 1998. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1030563992.

DasGupta:1991:RBE

[DS91] A. DasGupta and W. J. Studden. Robust Bayesian experimental designs in normal linear models. Annals of Statistics, 19(3):1244–1256, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348247.

Dette:1993:GO

[DS93] Holger Dette and William J. Studden. Geometry of Eoptimality. Annals of Statistics, 21(1):416–433, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349034.

Doksum:1995:NEG

[DS95a] Kjell Doksum and Alexander Samarov. Nonparametric estimation of global functionals and a measure of the explanatory power of covariates in regression. Annals of Statistics, 23(5): 1443–1473, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324307.

Dong:1995:GCE

[DS95b] Jianping Dong and Jeffrey S. Simonoff. A geometric combination estimator for d-dimensional ordinal sparse contingency tables. Annals of Statistics, 23(4):1143–1159, August 1995. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176324702.

Duchamp:1996:EPP

[DS96] Tom Duchamp and Werner Stuetzle. Extremal properties of principal curves in the plane. Annals of Statistics, 24(4): 1511–1520, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032298280.

DasGupta:1997:AEG

[DS97] Anirban DasGupta and William E. Strawderman. All estimates with a given risk, Riccati differential equations and a new proof of a theorem of Brown. Annals of Statistics, 25(3):1208–1221, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1069362745.

Diaconis:1998:AAS

[DS98] Persi Diaconis and Bernd Sturmfels. Algebraic algorithms for sampling from conditional distributions. Annals of Statistics, 26 (1):363–397, February 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1030563990.

Dawid:1999:CDC

[DS99a] A. Philip Dawid and Paola Sebastiani. Coherent dispersion criteria for optimal experimental design. Annals of Statistics, 27(1):65–81, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1018031101.

Dudoit:1999:STL

[DS99b] Sandrine Dudoit and Terence P. Speed. A score test for linkage using identity by descent data from sibships. Annals of Statistics, 27(3):943–986, June 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1018031264.

Davies:1998:IOT

[DT98] P. Laurie Davies and Wolfgang Terbeck. Interactions and outliers in the two-way analysis of variance. Annals of Statistics, 26(4):1279–1305, August 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691243.

Deng:1999:MAN

[DT99] Lih-Yuan Deng and Boxin Tang. Minimum G_{2}-aberration for nonregular fractional factorial designs. Annals of Statistics, 27(6):1914–1926, December 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1017939244.

Dudley:1994:ORD

[Dud94] R. M. Dudley. The order of the remainder in derivatives of composition and inverse operators for p-variation norms. Annals of Statistics, 22(1):1–20, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176325354.

Dumbgen:1991:ABS

[Dum91] L. Dumbgen. The asymptotic behavior of some nonparametric change-point estimators. Annals of Statistics, 19(3):1471– 1495, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348257.

Dumbgen:1998:NGF

[Düm98] Lutz Dümbgen. New goodness-of-fit tests and their application to nonparametric confidence sets. Annals of Statistics, 26(1): 288–314, February 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1030563987.

Dawid:1993:FPS

[DW93] A. Philip Dawid and Jinglong Wang. Fiducial prediction and semi-Bayesian inference. Annals of Statistics, 21(3):1119– 1138, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349253.

Dette:1995:GEC

[DW95] Holger Dette and Weng Kee Wong. On G-efficiency calculation for polynomial models. Annals of Statistics, 23(6):2081– 2101, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1034713648.

Dahlhaus:1996:AOE

[DW96a] R. Dahlhaus and W. Wefelmeyer. Asymptotically optimal estimation in misspecified time series models. Annals of Statistics, 24(3):952–974, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032526951.

Dette:1996:OBD

[DW96b] Holger Dette and Weng Kee Wong. Optimal Bayesian designs for models with partially specified heteroscedastic structure. Annals of Statistics, 24(5):2108–2127, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1069362313.

Eaton:1992:SDA

[Eat92] Morris L. Eaton. A statistical diptych: Admissible inferences — recurrence of symmetric Markov chains. Annals of Statistics, 20(3):1147–1179, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348764.

Erickson:1995:ODE

[EFM95] Roy V. Erickson, Vaclav Fabian, and Jan Marik. An optimum design for estimating the first derivative. Annals of Statistics, 23(4):1234–1247, August 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324707.

Efron:1990:DAP

[Efr90] B. Efron. Discussion: An ancillarity paradox which appears in multiple linear regression. Annals of Statistics, 18(2):502–503, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347606.

Efromovich:1995:SNE

[Efr95] Sam Efromovich. Sequential nonparametric estimation with assigned risk. Annals of Statistics, 23(4):1376–1392, August 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176324713.

Efromovich:1996:NRI

[Efr96] Sam Efromovich. On nonparametric regression for IID observations in a general setting. Annals of Statistics, 24(3):1126–1144, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1032526960.

Efromovich:1998:SSE

[Efr98] Sam Efromovich. Simultaneous sharp estimation of functions and their derivatives. Annals of Statistics, 26(1):273–278, February 1998. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1030563985.

Eubank:1992:TGF

[EH92] R. L. Eubank and Jeffrey D. Hart. Testing goodness-of-fit in regression via order selection criteria. Annals of Statistics, 20(3):1412–1425, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348775.

Eubank:1995:TAN

[EHSS95] R. L. Eubank, J. D. Hart, D. G. Simpson, and L. A. Stefanski. Testing for additivity in nonparametric regression. Annals of Statistics, 23(6):1896–1920, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1034713639.

Efron:1990:FIT

[EJ90] Bradley Efron and Iain M. Johnstone. Fisher’s information in terms of the hazard rate. Annals of Statistics, 18(1):38–62, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347492.

El-Krunz:1991:BOD

[EKS91] Sadi M. El-Krunz and W. J. Studden. Bayesian optimal designs for linear regression models. Annals of Statistics, 19(4): 2183–2208, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348392.

Eubank:1992:ACC

[EL92] R. L. Eubank and V. N. LaRiccia. Asymptotic comparison of Cramér–von Mises and nonparametric function estimation techniques for testing goodness-of-fit. Annals of Statistics, 20(4):2071–2086, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348903.

Eggermont:1995:MSL

[EL95] P. P. B. Eggermont and V. N. LaRiccia. Maximum smoothed likelihood density estimation for inverse problems. Annals of Statistics, 23(1):199–220, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324463.

Efromovich:1996:BRC

[EL96a] Sam Efromovich and Mark Low. On Bickel and Ritov’s conjecture about adaptive estimation of the integral of the square of density derivative. Annals of Statistics, 24(2):682–686, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032894459.

Efromovich:1996:OAE

[EL96b] Sam Efromovich and Mark Low. On optimal adaptive estimation of a quadratic functional. Annals of Statistics, 24(3): 1106–1125, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032526959.

Eggermont:1999:OCR

[EL99] P. P. B. Eggermont and V. N. LaRiccia. Optimal convergence rates for Good’s nonparametric maximum likelihood density estimator. Annals of Statistics, 27(5):1600–1615, October 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1017939143.

Ellis:1991:SFP

[Ell91] Steven P. Ellis. The singularities of fitting planes to data. Annals of Statistics, 19(3):1661–1666, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348269.

Ellis:1995:DSS

[Ell95] Steven P. Ellis. Dimension of the singular sets of planefitters. Annals of Statistics, 23(2):490–501, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176324532.

Emoto:1990:WMD

[EM90] Sherrie E. Emoto and Peter C. Matthews. A Weibull model for dependent censoring. Annals of Statistics, 18(4):1556–1577, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347866.

Einmahl:1992:GQP

[EM92] John H. J. Einmahl and David M. Mason. Generalized quantile processes. Annals of Statistics, 20(2):1062–1078, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348670.

Einmahl:1999:CTM

[EM99] John H. J. Einmahl and Ian W. McKeague. Confidence tubes for multiple quantile plots via empirical likelihood. Annals of Statistics, 27(4):1348–1367, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1017938929.

Erickson:1996:REL

[Eri96] R. V. Erickson. Robust estimation of the location of a vertical tangent in distribution. Annals of Statistics, 24(3):1423–1431, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1032526977.

Evans:1993:ISP

[ES93] Steven N. Evans and T. P. Speed. Invariants of some probability models used in phylogenetic inference. Annals of Statistics, 21 (1):355–377, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349030.

Evans:1996:SES

[ES96] Steven N. Evans and Philip B. Stark. Shrinkage estimators, Skorokhod’s problem and stochastic integration by parts. Annals of Statistics, 24(2):809–815, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1032894466.

Eaton:1991:WIA

[ET91] Morris L. Eaton and David E. Tyler. On Wielandt’s inequality and its application to the asymptotic distribution of the eigenvalues of a random symmetric matrix. Annals of Statistics, 19 (1):260–271, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347980.

Efron:1996:USD

[ET96] Bradley Efron and Robert Tibshirani. Using specially designed exponential families for density estimation. Annals of Statistics, 24(6):2431–2461, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1032181161.

Efron:1998:PR

[ET98] Bradley Efron and Robert Tibshirani. The problem of regions. Annals of Statistics, 26(5):1687–1718, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1024691353.

Evans:1991:CA

[Eva91] M. Evans. Chaining via annealing. Annals of Statistics, 19 (1):382–393, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347989.

Eisele:1995:CLT

[EW95] Jeffrey R. Eisele and Michael B. Woodroofe. Central limit theorems for doubly adaptive biased coin designs. Annals of Statistics, 23(1):234–254, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324465.

Falk:1995:TEV

[Fal95] Michael Falk. On testing the extreme value index via the pot-method. Annals of Statistics, 23(6):2013–2035, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1034713645.

Falk:1998:LAN

[Fal98] Michael Falk. Local asymptotic normality of truncated empirical processes. Annals of Statistics, 26(2):692–718, April 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1028144855.

Fan:1991:EQF

[Fan91a] Jianqing Fan. On the estimation of quadratic functionals. Annals of Statistics, 19(3):1273–1294, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348249.

Fan:1991:ORC

[Fan91b] Jianqing Fan. On the optimal rates of convergence for nonparametric deconvolution problems. Annals of Statistics, 19 (3):1257–1272, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348248.

Fan:1993:ALO

[Fan93a] Jianqing Fan. Adaptively local one-dimensional subproblems with application to a deconvolution problem. Annals of Statistics, 21(2):600–610, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176349139.

Fan:1993:LLR

[Fan93b] Jianqing Fan. Local linear regression smoothers and their minimax efficiencies. Annals of Statistics, 21(1):196–216, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349022.

Faraway:1992:SAE

[Far92] Julian J. Faraway. Smoothing in adaptive estimation. Annals of Statistics, 20(1):414–427, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348530.

Feldman:1991:CME

[Fel91] Israel Feldman. Constrained minimax estimation of the mean of the normal distribution with known variance. Annals of Statistics, 19(4):2259–2265, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348398.

Ferger:1995:NTN

[Fer95] D. Ferger. Nonparametric tests for nonstandard change-point problems. Annals of Statistics, 23(5):1848–1861, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176324326.

Francisco:1991:QEC

[FF91] Carol A. Francisco and Wayne A. Fuller. Quantile estimation with a complex survey design. Annals of Statistics, 19 (1):454–469, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347993.

Fan:1992:VBL

[FG92] Jianqing Fan and Irene Gijbels. Variable bandwidth and local linear regression smoothers. Annals of Statistics, 20(4): 2008–2036, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348900.

Foster:1993:ECP

[FG93] Dean P. Foster and Edward I. George. Estimation up to a change-point. Annals of Statistics, 21(2):625–644, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349141.

Foster:1994:RIC

[FG94] Dean P. Foster and Edward I. George. The risk inflation criterion for multiple regression. Annals of Statistics, 22(4):1947– 1975, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325766.

Fan:1997:LLL

[FGK97] Jianqing Fan, Irène Gijbels, and Martin King. Local likelihood and local partial likelihood in hazard regression. Annals of Statistics, 25(4):1661–1690, August 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1031594736.

Fenech:1991:ECS

[FH91] Alan P. Fenech and David A. Harville. Exact confidence sets for variance components in unbalanced mixed linear models. Annals of Statistics, 19(4):1771–1785, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348370.

Franke:1992:BKS

[FH92] J. Franke and W. Hardle. On bootstrapping kernel spectral estimates. Annals of Statistics, 20(1):121–145, March 1992. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348515.

Fan:1994:CEM

[FH94] Jianqing Fan and Peter Hall. On curve estimation by minimizing mean absolute deviation and its implications. Annals of Statistics, 22(2):867–885, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325499.

Feuerverger:1996:DFI

[FH96] Andrey Feuerverger and Peter Hall. On distribution-free inference for record-value data with trend. Annals of Statistics, 24(6):2655–2678, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1032181174.

Feuerverger:1999:ETE

[FH99] Andrey Feuerverger and Peter Hall. Estimating a tail exponent by modelling departure from a Pareto distribution. Annals of Statistics, 27(2):760–781, April 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1018031215.

Fan:1998:DEL

[FHM98] Jianqing Fan, Wolfgang Härdle, and Enno Mammen. Direct estimation of low-dimensional components in additive models. Annals of Statistics, 26(3):943–971, June 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1024691083.

Finner:1994:TST

[Fin94] H. Finner. Two-sided tests and one-sided confidence bounds. Annals of Statistics, 22(3):1502–1516, September 1994. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176325639.

Finner:1999:SMT

[Fin99] H. Finner. Stepwise multiple test procedures and control of directional errors. Annals of Statistics, 27(1):274–289, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1018031111.

Falk:1991:CPB

[FK91] Michael Falk and Edgar Kaufmann. Coverage probabilities of bootstrap-confidence intervals for quantiles. Annals of Statistics, 19(1):485–495, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347995.

Flury:1994:DAA

[Flu94] Bernard D. Flury. Discussion: Analysis of additive dependencies and concurvities using smallest additive principal components. Annals of Statistics, 22(4):1668–1670, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325747.

Fan:1992:BPC

[FM92] Jianqing Fan and James S. Marron. Best possible constant for bandwidth selection. Annals of Statistics, 20(4):2057–2070, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176348902.

Falk:1993:AOT

[FM93] M. Falk and F. Marohn. Asymptotically optimal tests for conditional distributions. Annals of Statistics, 21(1):45–60, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349014.

Foster:1991:PWC

[Fos91] Dean P. Foster. Prediction in the worst case. Annals of Statistics, 19(2):1084–1090, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348140.

Fraser:1990:DAP

[FR90] D. A. S. Fraser and N. Reid. Discussion: An ancillarity paradox which appears in multiple linear regression. Annals of Statistics, 18(2):503–507, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347607.

Falk:1992:SIC

[FR92] M. Falk and R.-D. Reiss. Statistical inference for conditional curves: Poisson process approach. Annals of Statistics, 20 (2):779–796, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348656.

Fygenson:1994:MEE

[FR94] Mendel Fygenson and Ya’acov Ritov. Monotone estimating equations for censored data. Annals of Statistics, 22(2):732– 746, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1176325493.

Finner:1998:ACS

[FR98] H. Finner and M. Roters. Asymptotic comparison of stepdown and step-up multiple test procedures based on exchangeable test statistics. Annals of Statistics, 26(2):505–524, April 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1028144847.

Freedman:1999:WLB

[Fre99] David Freedman. Wald lecture: On the Bernstein–von Mises theorem with infinite-dimensional parameters. Annals of Statistics, 27(4):1119–1141, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1017938917.

Friedman:1991:MAR

[Fri91a] Jerome H. Friedman. Multivariate adaptive regression splines. Annals of Statistics, 19(1):1–67, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347963.

Friedman:1991:RMA

[Fri91b] Jerome H. Friedman. Rejoinder: Multivariate adaptive regression splines. Annals of Statistics, 19(1):123–141, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347973.

Frydenberg:1990:MCG

[Fry90] Morten Frydenberg. Marginalization and collapsibility in graphical interaction models. Annals of Statistics, 18(2):790– 805, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1176347626.

Faria:1997:CEB

[FS97] Álvaro E. Faria and Jim Q. Smith. Conditionally externally Bayesian pooling operators in chain graphs. Annals of Statistics, 25(4):1740–1761, August 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1031594740.

Fourdrinier:1998:CBM

[FSW98] Dominique Fourdrinier, William E. Strawderman, and Martin T. Wells. On the construction of Bayes minimax estimators. Annals of Statistics, 26(2):660–671, April 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1028144853.

Fan:1993:NRE

[FT93] Jianqing Fan and Young K. Truong. Nonparametric regression with errors in variables. Annals of Statistics, 21(4):1900– 1925, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349402.

Fourdrinier:1995:ELF

[FW95] Dominique Fourdrinier and Martin T. Wells. Estimation of a loss function for spherically symmetric distributions in the general linear model. Annals of Statistics, 23(2):571–592, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176324536.

Fygenson:1994:USA

[FZ94] Mendel Fygenson and Mai Zhou. On using stratification in the analysis of linear regression models with right censoring. Annals of Statistics, 22(2):747–762, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176325494.

Fan:1999:SEV

[FZ99] Jianqing Fan and Wenyang Zhang. Statistical estimation in varying coefficient models. Annals of Statistics, 27(5):1491– 1518, October 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1017939139.

Fakhre-Zakeri:1996:OSS

[FZS96] Issa Fakhre-Zakeri and Eric Slud. Optimal stopping of sequential size-dependent search. Annals of Statistics, 24(5): 2215–2232, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362318.

Gassiat:1993:AEN

[Gas93] E. Gassiat. Adaptive estimation in noncausal stationary AR processes. Annals of Statistics, 21(4):2022–2042, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349408.

Gasparini:1995:EMB

[Gas95] Mauro Gasparini. Exact multivariate Bayesian bootstrap distributions of moments. Annals of Statistics, 23(3):762–768, June 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176324620.

Goutis:1991:IIC

[GC91] Constantinos Goutis and George Casella. Improved invariant confidence intervals for a normal variance. Annals of Statistics, 19(4):2015–2031, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348384.

Goutis:1992:ICS

[GC92] Constantinos Goutis and George Casella. Increasing the confidence in Student’s t interval. Annals of Statistics, 20(3):1501– 1513, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348781.

Gilardoni:1993:RCU

[GC93] Gustavo L. Gilardoni and Murray K. Clayton. On reaching a consensus using Degroot’s iterative pooling. Annals of Statistics, 21(1):391–401, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349032.

Garthwaite:1992:EPD

[GD92] Paul H. Garthwaite and James M. Dickey. Elicitation of prior distributions for variable-selection problems in regression. Annals of Statistics, 20(4):1697–1719, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348886.

Garcia-Escudero:1999:CLT

[GEGM99] Luis A. García-Escudero, Alfonso Gordaliza, and Carlos Matrán. A central limit theorem for multivariate generalized trimmed k-means. Annals of Statistics, 27(3):1061–1079, June 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1018031268.

Genest:1992:VR

[Gen92] Christian Genest. Vincentization revisited. Annals of Statistics, 20(2):1137–1142, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348676.

George:1991:SDM

[Geo91] Edward I. George. Shrinkage domination in a multivariate common mean problem. Annals of Statistics, 19(2):952–960, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348130.

Geyer:1994:ACE

[Gey94] Charles J. Geyer. On the asymptotics of constrained Mestimation. Annals of Statistics, 22(4):1993–2010, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325768.

Gamboa:1996:BDD

[GG96a] F. Gamboa and E. Gassiat. Blind deconvolution of discrete linear systems. Annals of Statistics, 24(5):1964–1981, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1069362305.

Goldstein:1996:AMH

[GG96b] Larry Goldstein and Louis Gordon. Asymptotics for the maximum of half-normal plots. Annals of Statistics, 24(5):2250– 2265, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362320.

Gamboa:1997:BMM

[GG97] F. Gamboa and E. Gassiat. Bayesian methods and maximum entropy for ill-posed inverse problems. Annals of Statistics, 25 (1):328–350, February 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1034276632.

Gassiat:1999:SCB

[GG99a] E. Gassiat and E. Gautherat. Speed of convergence for the blind deconvolution of a linear system with discrete random input. Annals of Statistics, 27(5):1684–1705, October 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1017939147.

Geskus:1999:AOE

[GG99b] Ronald Geskus and Piet Groeneboom. Asymptotically optimal estimation of smooth functionals for interval censoring, case 2. Annals of Statistics, 27(2):627–674, April 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1018031211.

Ghosal:1999:PCD

[GGR99] S. Ghosal, J. K. Ghosh, and R. V. Ramamoorthi. Posterior consistency of Dirichlet mixtures in density estimation. Annals of Statistics, 27(1):143–158, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1018031105.

Ghosal:1995:CPD

[GGS95] Subhashis Ghosal, Jayanta K. Ghosh, and Tapas Samanta. On convergence of posterior distributions. Annals of Statistics, 23(6):2145–2152, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1034713651.

Ghosh:1991:POK

[GH91a] B. K. Ghosh and Wei-Min Huang. The power and optimal kernel of the Bickel–Rosenblatt test for goodness of fit. Annals of Statistics, 19(2):999–1009, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348133.

Golubev:1991:DMA

[GH91b] George K. Golubev and Rafael Z. Hasminskii. Discussion: Multivariate adaptive regression splines. Annals of Statistics, 19 (1):91–92, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347966.

Gotze:1994:ADS

[GH94] F. Gotze and C. Hipp. Asymptotic distribution of statistics in time series. Annals of Statistics, 22(4):2062–2088, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325772.

Gilula:1995:PFC

[GH95a] Zvi Gilula and Shelby J. Haberman. Prediction functions for categorical panel data. Annals of Statistics, 23(4):1130–1142, August 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176324701.

Grund:1995:MEM

[GH95b] Birgit Grund and Peter Hall. On the minimisation of L^{p} error in mode estimation. Annals of Statistics, 23(6):2264–2284, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1034713656.

Gather:1997:NTM

[GH97a] Ursula Gather and Torsten Hilker. A note on Tyler’s modification of the MAD for the Stahel–Donoho estimator. Annals of Statistics, 25(5):2024–2026, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1069362384.

Geiger:1997:CDD

[GH97b] Dan Geiger and David Heckerman. A characterization of the Dirichlet distribution through global and local parameter independence. Annals of Statistics, 25(3):1344–1369, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1069362752.

Groeneboom:1999:ANE

[GHL99] Piet Groeneboom, Gerard Hooghiemstra, and Hendrik P. Lopuhaä. Asymptotic normality of the L_{1} error of the Grenander estimator. Annals of Statistics, 27(4):1316–1347, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1017938928.

Ghosh:1980:MEP

[Gho80] Subir Ghosh. On main effect plus one plans for 2^{m} factorials. Annals of Statistics, 8(4):922–930, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345084. See

correction [Gho91].

Ghosh:1991:CME

[Gho91] Subir Ghosh. Correction: “On Main Effect plus One Plans for 2^{m} Factorials”. Annals of Statistics, 19(4):2281, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/

1176348402. See [Gho80].

Girard:1991:AOF

[Gir91] Didier A. Girard. Asymptotic optimality of the fast randomized versions of GCV and C_{L} in ridge regression and regularization. Annals of Statistics, 19(4):1950–1963, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348380.

Girard:1998:ACP

[Gir98] Didier A. Girard. Asymptotic comparison of (partial) crossvalidation, GCV and randomized GCV in nonparametric regression. Annals of Statistics, 26(1):315–334, February 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1030563988.

Gill:1990:SPI

[GJ90] Richard D. Gill and Soren Johansen. A survey of productintegration with a view toward application in survival analysis. Annals of Statistics, 18(4):1501–1555, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347865.

Gutenbrunner:1992:RRS

[GJ92] C. Gutenbrunner and J. Jureckova. Regression rank scores and regression quantiles. Annals of Statistics, 20(1):305–330, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348524.

Groeneboom:1995:IER

[GJ95] Piet Groeneboom and Geurt Jongbloed. Isotonic estimation and rates of convergence in Wicksell’s problem. Annals of Statistics, 23(5):1518–1542, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324310.

Gajek:1994:LBA

[GK94] L. Gajek and M. Kaluszka. Lower bounds for the asymptotic Bayes risk in the scale model (with an application to the second-order minimax estimation). Annals of Statistics, 22(4): 1831–1839, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325759.

Gotze:1996:SOC

[GK96] F. Götze and H. R. Künsch. Second-order correctness of the blockwise bootstrap for stationary observations. Annals of Statistics, 24(5):1914–1933, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1069362303.

Glimm:1998:MTB

[GKL98] Ekkehard Glimm, Siegfried Kropf, and Jürgen Läuter. Multivariate tests based on left-spherically distributed linear scores. Annals of Statistics, 26(5):1972–1988, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1024691365. See

correction [GKL99].

Glimm:1999:CMT

[GKL99] E. Glimm, S. Kropf, and J. Lauter. Correction: “Multivariate tests based on left-spherically distributed linear scores”. Annals of Statistics, 27(4):1441, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1017938934. See [GKL98].

Gu:1991:WCT

[GL91] Ming Gao Gu and Tze Leung Lai. Weak convergence of timesequential censored rank statistics with applications to sequential testing in clinical trials. Annals of Statistics, 19(3):1403– 1433, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348254.

Goldstein:1992:ATN

[GL92] Larry Goldstein and Bryan Langholz. Asymptotic theory for nested case-control sampling in the Cox regression model. Annals of Statistics, 20(4):1903–1928, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348895.

Golubev:1998:AEE

[GL98] G. K. Golubev and B. Y. Levit. Asymptotically efficient estimation in the Wicksell problem. Annals of Statistics, 26(6): 2407–2419, December 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1024691477.

Gleser:1990:DAP

[Gle90] Leon Jay Gleser. Discussion: An ancillarity paradox which appears in multiple linear regression. Annals of Statistics, 18 (2):507–513, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347608.

Goldstein:1992:OPE

[GM92] Larry Goldstein and Karen Messer. Optimal plug-in estimators for nonparametric functional estimation. Annals of Statistics, 20(3):1306–1328, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348770.

Goodall:1993:MAS

[GM93] Colin R. Goodall and Kanti V. Mardia. Multivariate aspects of shape theory. Annals of Statistics, 21(2):848–866, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349154.

Greenwood:1996:OGS

[GMW96] Priscilla E. Greenwood, Ian W. McKeague, and Wolfgang Wefelmeyer. Outperforming the Gibbs sampler empirical estimator for nearest-neighbor random fields. Annals of Statistics, 24(4): 1433–1456, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032298276.

Greenwood:1998:IBG

[GMW98] Priscilla E. Greenwood, Ian W. McKeague, and Wolfgang Wefelmeyer. Information bounds for Gibbs samplers. Annals of Statistics, 26(6):2128–2156, December 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691464.

Golubev:1990:RBS

[GN90] Grigori K. Golubev and Michael Nussbaum. A risk bound in Sobolev class regression. Annals of Statistics, 18(2):758–778, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347624.

Goldstein:1991:BTC

[Gol91] Michael Goldstein. Belief transforms and the comparison of hypotheses. Annals of Statistics, 19(4):2067–2089, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348386.

Goutis:1996:PLS

[Gou96] Constantinos Goutis. Partial least squares algorithm yields shrinkage estimators. Annals of Statistics, 24(2):816–824, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032894467.

Geiger:1993:LAP

[GP93a] Dan Geiger and Judea Pearl. Logical and algorithmic properties of conditional independence and graphical models. Annals of Statistics, 21(4):2001–2021, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349407.

Grubel:1993:NER

[GP93b] Rudolf Grubel and Susan M. Pitts. Nonparametric estimation in renewal theory i: The empirical renewal function. Annals of Statistics, 21(3):1431–1451, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349266.

Gordon:1994:ESN

[GP94] Louis Gordon and Moshe Pollak. An efficient sequential nonparametric scheme for detecting a change of distribution. Annals of Statistics, 22(2):763–804, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176325495.

Gordon:1995:RSS

[GP95] Louis Gordon and Moshe Pollak. A robust surveillance scheme for stochastically ordered alternatives. Annals of Statistics, 23(4):1350–1375, August 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324712.

Gordon:1997:ARL

[GP97] Louis Gordon and Moshe Pollak. Average run length to false alarm for surveillance schemes designed with partially specified pre-change distribution. Annals of Statistics, 25(3):1284–1310, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1069362749.

Gu:1993:SSD

[GQ93] Chong Gu and Chunfu Qiu. Smoothing spline density estimation: Theory. Annals of Statistics, 21(1):217–234, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349023.

Greenshtein:1993:NBT

[Gre93] Eitan Greenshtein. A note on Bahadur’s transitivity. Annals of Statistics, 21(4):2163–2167, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349417.

Greenshtein:1996:CSE

[Gre96] Eitan Greenshtein. Comparison of sequential experiments. Annals of Statistics, 24(1):436–448, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1033066220.

Grubel:1996:OML

[Grü96a] Rudolf Grübel. Orthogonalization of multivariate location estimators: the orthomedian. Annals of Statistics, 24(4):1457– 1473, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1032298277.

Gruet:1996:NCA

[Gru96b] Marie-Anne Gruet. A nonparametric calibration analysis. Annals of Statistics, 24(4):1474–1492, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1032298278.

Ghosh:1994:SOP

[GSM94] Jayanta K. Ghosh, Pranab K. Sen, and Rahul Mukerjee. Second-order Pitman closeness and Pitman admissibility. Annals of Statistics, 22(3):1133–1141, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325621.

Giraitis:1999:WEF

[GT99] Liudas Giraitis and Murad S. Taqqu. Whittle estimator for finite-variance non-Gaussian time series with long memory. Annals of Statistics, 27(1):178–203, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1018031107.

Gustafson:1996:LSP

[Gus96] Paul Gustafson. Local sensitivity of posterior expectations. Annals of Statistics, 24(1):174–195, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1033066205.

Gijbels:1991:ASA

[GV91] Irene Gijbels and Noel Veraverbeke. Almost sure asymptotic representation for a class of functionals of the Kaplan–Meier estimator. Annals of Statistics, 19(3):1457–1470, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348256.

Gu:1991:DMA

[GW91] Chong Gu and Grace Wahba. Discussion: Multivariate adaptive regression splines. Annals of Statistics, 19(1):115–123, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347972.

Gittins:1992:LCD

[GW92] John Gittins and You-Gan Wang. The learning component of dynamic allocation indices. Annals of Statistics, 20(3):1625– 1636, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348788.

Greenwood:1995:EEE

[GW95a] P. E. Greenwood and W. Wefelmeyer. Efficiency of empirical estimators for Markov chains. Annals of Statistics, 23(1): 132–143, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324459.

Gustafson:1995:LSD

[GW95b] Paul Gustafson and Larry Wasserman. Local sensitivity diagnostics for Bayesian inference. Annals of Statistics, 23(6): 2153–2167, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1034713652.

Gasser:1998:ABC

[GW98] Theo Gasser and Kongming Wang. Asymptotic and bootstrap confidence bounds for the structural average of curves. Annals of Statistics, 26(3):972–991, June 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691084.

Gasser:1999:SSC

[GW99] Theo Gasser and Kongming Wang. Synchronizing sample curves nonparametrically. Annals of Statistics, 27(2):439–460, April 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1018031202.

Gu:1993:APS

[GZ93] M. G. Gu and C.-H. Zhang. Asymptotic properties of selfconsistent estimators based on doubly censored data. Annals of Statistics, 21(2):611–624, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176349140.

Haff:1991:VFC

[Haf91] L. R. Haff. The variational form of certain Bayes estimators. Annals of Statistics, 19(3):1163–1190, September 1991. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348244.

Hall:1990:PLT

[Hal90] Peter Hall. Pseudo-likelihood theory for empirical likelihood. Annals of Statistics, 18(1):121–140, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347495.

Hall:1991:BRU

[Hal91] Peter Hall. Bahadur representations for uniform resampling and importance resampling, with applications to asymptotic relative efficiency. Annals of Statistics, 19(2):1062–1072, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348138.

Hall:1992:EBE

[Hal92a] Peter Hall. Effect of bias estimation on coverage accuracy of bootstrap confidence intervals for a probability density. Annals of Statistics, 20(2):675–694, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348651.

Hall:1992:BCI

[Hal92b] Peter Hall. On bootstrap confidence intervals in nonparametric regression. Annals of Statistics, 20(2):695–711, June 1992. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348652.

Hall:1992:GPV

[Hal92c] Peter Hall. On global properties of variable bandwidth density estimators. Annals of Statistics, 20(2):762–778, June 1992. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348655.

Hall:1993:PRL

[Hal93] Peter Hall. On plug-in rules for local smoothing of density estimators. Annals of Statistics, 21(2):694–710, June 1993. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349145.

Hartigan:1994:LEC

[Har94] J. A. Hartigan. Linear estimators in change point problems. Annals of Statistics, 22(2):824–834, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176325497.

Hartigan:1996:LUP

[Har96] J. A. Hartigan. Locally uniform prior distributions. Annals of Statistics, 24(1):160–173, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1033066204.

Hartigan:1998:MLP

[Har98] J. A. Hartigan. The maximum likelihood prior. Annals of Statistics, 26(6):2083–2103, December 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691462.

Hastie:1994:DUP

[Has94] Trevor Hastie. Discussion: The use of polynomial splines and their tensor products in multivariate function estimation. Annals of Statistics, 22(1):177–179, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325363.

Hassairi:1999:GVE

[Has99] Abdelhamid Hassairi. Generalized variance and exponential families. Annals of Statistics, 27(1):374–385, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1018031116.

Hwang:1991:ECU

[HB91] Jiunn T. Hwang and Lawrence D. Brown. Estimated confidence under the validity constraint. Annals of Statistics, 19(4): 1964–1977, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348381.

Hastie:1995:PDA

[HBT95] Trevor Hastie, Andreas Buja, and Robert Tibshirani. Penalized discriminant analysis. Annals of Statistics, 23(1):73–102, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1176324456.

Hsing:1992:ATS

[HC92] Tailen Hsing and Raymond J. Carroll. An asymptotic theory for sliced inverse regression. Annals of Statistics, 20(2):1040–1061, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176348669.

Hwang:1992:EAT

[HCR^{+}92] Jiunn Tzon Hwang, George Casella, Christian Robert, Martin T. Wells, and Roger H. Farrell. Estimation of accuracy in testing. Annals of Statistics, 20(1):490–509, March 1992. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348534.

Heiligers:1994:ODW

[Hei94] Berthold Heiligers. E-optimal designs in weighted polynomial regression. Annals of Statistics, 22(2):917–929, June 1994. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176325503.

Helmers:1991:EEB

[Hel91] R. Helmers. On the Edgeworth expansion and the bootstrap approximation for a studentized U-statistic. Annals of Statistics, 19(1):470–484, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347994.

Hendriks:1990:NEP

[Hen90] Harrie Hendriks. Nonparametric estimation of a probability density on a Riemannian manifold using Fourier expansions. Annals of Statistics, 18(2):832–849, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347628.

Hengartner:1997:ADP

[Hen97] Nicolas W. Hengartner. Adaptive demixing in Poisson mixture models. Annals of Statistics, 25(3):917–928, June 1997. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1069362730.

Hesse:1990:BTR

[Hes90] C. H. Hesse. A Bahadur-type representation for empirical quantiles of a large class of stationary, possibly infinite-variance, linear processes. Annals of Statistics, 18(3):1188–1202, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347746.

Hess:1996:ECS

[Hes96] Christian Hess. Epi-convergence of sequences of normal integrands and strong consistency of the maximum likelihood estimator. Annals of Statistics, 24(3):1298–1315, June 1996. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1032526970.

Heyde:1992:BAC

[Hey92] C. C. Heyde. On best asymptotic confidence intervals for parameters of stochastic processes. Annals of Statistics, 20(1): 603–607, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348545.

Hjort:1992:LTN

[HF92] Nils Lid Hjort and Grete Fenstad. On the last time and the number of times an estimator is more than ε from its target value. Annals of Statistics, 20(1):469–489, March 1992. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348533.

Hall:1994:NEC

[HFH94] Peter Hall, Nicholas I. Fisher, and Branka Hoffmann. On the nonparametric estimation of covariance functions. Annals of Statistics, 22(4):2115–2134, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176325774.

Huang:1990:ENA

[HG90] Dawei Huang and Lei Guo. Estimation of nonstationary AR-MAX models based on the Hannan–Rissanen method. Annals of Statistics, 18(4):1729–1756, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347875.

Hjort:1995:NDE

[HG95] Nils Lid Hjort and Ingrid K. Glad. Nonparametric density estimation with a parametric start. Annals of Statistics, 23 (3):882–904, June 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324627.

Huang:1994:RSP

[HH94] J. S. Huang and W. J. Huang. A reverse submartingale property of the range. Annals of Statistics, 22(1):474–477, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325379.

Ho:1996:AEE

[HH96] Hwai-Chung Ho and Tailen Hsing. On the asymptotic expansion of the empirical process of long-memory moving averages. Annals of Statistics, 24(3):992–1024, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1032526953.

Hardle:1993:OSS

[HHI93] Wolfgang Hardle, Peter Hall, and Hidehiko Ichimura. Optimal smoothing in single-index models. Annals of Statistics, 21 (1):157–178, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349020.

Hall:1995:IVW

[HHM95] Peter Hall, Tien Chung Hu, and J. S. Marron. Improved variable window kernel estimates of probability densities. Annals of Statistics, 23(1):1–10, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324451.

Hasminskii:1990:DEV

[HI90] Rafael Hasminskii and Ildar Ibragimov. On density estimation in the view of Kolmogorov’s ideas in approximation theory. Annals of Statistics, 18(3):999–1010, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347736.

Hill:1990:DAP

[Hil90] B. M. Hill. Discussion: An ancillarity paradox which appears in multiple linear regression. Annals of Statistics, 18(2):513–523, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347609.

Hall:1990:AEN

[HJ90] Peter Hall and M. C. Jones. Adaptive M-estimation in nonparametric regression. Annals of Statistics, 18(4):1712–1728, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1176347874.

Huskova:1993:CGB

[HJ93] Marie Huskova and Paul Janssen. Consistency of the generalized bootstrap for degenerate U-statistics. Annals of Statistics, 21(4):1811–1823, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176349399.

Hall:1995:UCB

[HJ95] Peter Hall and Bing-Yi Jing. Uniform coverage bounds for confidence intervals and Berry–Esseen theorems for Edgeworth expansion. Annals of Statistics, 23(2):363–375, April 1995. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176324525.

Hjort:1996:LPN

[HJ96a] N. L. Hjort and M. C. Jones. Locally parametric nonparametric density estimation. Annals of Statistics, 24(4):1619–1647, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1032298288.

Holmes:1996:UGT

[HJ96b] R. B. Holmes and L. K. Jones. On uniform generation of twoway tables with fixed margins and the conditional volume test of Diaconis and Efron. Annals of Statistics, 24(1):64–68, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1033066199.

Hallin:1999:OTA

[HJ99] Marc Hallin and Jana Jurečková. Optimal tests for autoregressive models based on autoregression rank scores. Annals of Statistics, 27(4):1385–1414, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1017938931.

Hjort:1990:GFT

[Hjo90a] Nils Lid Hjort. Goodness of fit tests in models for life history data based on cumulative hazard rates. Annals of Statistics, 18(3):1221–1258, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347748.

Hjort:1990:NBE

[Hjo90b] Nils Lid Hjort. Nonparametric Bayes estimators based on beta processes in models for life history data. Annals of Statistics, 18(3):1259–1294, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347749.

Hawkins:1991:ICP

[HK91] D. L. Hawkins and Subhash C. Kochar. Inference for the crossing point of two continuous CDF’s. Annals of Statistics, 19(3):1626–1638, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348266.

Healy:1996:EBA

[HK96] Dennis M. Healy and Peter T. Kim. An empirical Bayes approach to directional data and efficient computation on the sphere. Annals of Statistics, 24(1):232–254, February 1996. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1033066208.

Hawkins:1992:TEA

[HKL92] D. L. Hawkins, Subhash Kochar, and Clive Loader. Testing exponentiality against IDMRL distributions with unknown change point. Annals of Statistics, 20(1):280–290, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348522.

Hall:1998:BTR

[HKP98] Peter Hall, Gérard Kerkyacharian, and Dominique Picard. Block threshold rules for curve estimation using kernel and wavelet methods. Annals of Statistics, 26(3):922–942, June 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1024691082.

Hall:1997:NCR

[HKT97] Peter Hall, Hira L. Koul, and Berwin A. Turlach. Note on convergence rates of semiparametric estimators of dependence index. Annals of Statistics, 25(4):1725–1739, August 1997. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1031594739.

Hayter:1990:PFS

[HL90a] A. J. Hayter and W. Liu. The power function of the Studentised range test. Annals of Statistics, 18(1):465–468, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347512.

Hoffman:1990:LSP

[HL90b] Wherly P. Hoffman and Sue E. Leurgans. Large sample properties of two tests for independent joint action of two drugs. Annals of Statistics, 18(4):1634–1650, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347870.

Hwang:1990:SCR

[HL90c] Jiunn T. Hwang and Hung-Kung Liu. Sequential confidence regions in inverse regression problems. Annals of Statistics, 18(3):1389–1399, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347756.

Hall:1993:ALL

[HL93] Peter Hall and Ker-Chau Li. On almost linearity of low dimensional projections from high dimensional data. Annals of Statistics, 21(2):867–889, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176349155.

Hall:1995:BCN

[HLP95] Peter Hall, Soumendra Nath Lahiri, and Jörg Polzehl. On bandwidth choice in nonparametric regression with both shortand long-range dependent errors. Annals of Statistics, 23(6): 1921–1936, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1034713640.

Hall:1995:BCD

[HLT95] Peter Hall, Soumendra Nath Lahiri, and Young K. Truong. On bandwidth choice for density estimation with dependent data. Annals of Statistics, 23(6):2241–2263, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1034713655.

Hardle:1990:SCR

[HM90] W. Hardle and J. S. Marron. Semiparametric comparison of regression curves. Annals of Statistics, 18(1):63–89, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347493.

Hardle:1991:BSE

[HM91] W. Hardle and J. S. Marron. Bootstrap simultaneous error Bars for nonparametric regression. Annals of Statistics, 19 (2):778–796, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348120.

Hardle:1993:CNV

[HM93] W. Hardle and E. Mammen. Comparing nonparametric versus parametric regression fits. Annals of Statistics, 21(4):1926– 1947, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349403.

Hall:1994:GRA

[HM94] Peter Hall and Enno Mammen. On general resampling algorithms and their performance in distribution estimation. Annals of Statistics, 22(4):2011–2030, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325769.

Hill:1998:CRD

[HM98] Theodore Hill and Michael Monticino. Constructions of random distributions via sequential barycenters. Annals of Statistics, 26(4):1242–1253, August 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691241.

Hallin:1992:PEV

[HMM92] Marc Hallin, Guy Melard, and Xavier Milhaud. Permutational extreme values of autocorrelation coefficients and a Pitman test against serial dependence. Annals of Statistics, 20(1):523–534, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176348536.

Hall:1997:CEW

[HMNT97] Peter Hall, J. S. Marron, M. H. Neumann, and D. M. Titterington. Curve estimation when the design density is low. Annals of Statistics, 25(2):756–770, April 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1031833672.

Hall:1996:PWM

[HMT96] Peter Hall, Ian McKay, and Berwin A. Turlach. Performance of wavelet methods for functions with many discontinuities. Annals of Statistics, 24(6):2462–2476, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1032181162.

Haussler:1997:MIM

[HO97] David Haussler and Manfred Opper. Mutual information, metric entropy and cumulative relative entropy risk. Annals of Statistics, 25(6):2451–2492, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1030741081.

Horvath:1991:NMD

[Hor91] Lajos Horváth. On L_{p}-norms of multivariate density estimators. Annals of Statistics, 19(4):1933–1949, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348379.

Horvath:1993:MLM

[Hor93] Lajos Horváth. The maximum likelihood method for testing changes in the parameters of normal observations. Annals of Statistics, 21(2):671–680, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349143.

Hossjer:1996:IGS

[Hös96] Ola Hössjer. Incomplete generalized L-statistics. Annals of Statistics, 24(6):2631–2654, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1032181173.

Hosoya:1997:LTL

[Hos97] Yuzo Hosoya. A limit theory for long-range dependence and statistical inference on related models. Annals of Statistics, 25 (1):105–137, February 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1034276623.

He:1992:RLE

[HP92] Xuming He and Stephen Portnoy. Reweighted LS estimators converge at the same rate as the initial estimator. Annals of Statistics, 20(4):2161–2167, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348910.

Hwang:1994:CIE

[HP94] J. T. Gene Hwang and Shyamal Das Peddada. Confidence interval estimation subject to order restrictions. Annals of Statistics, 22(1):67–93, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325358.

Hall:1995:FMI

[HP95] Peter Hall and Prakash Patil. Formulae for mean integrated squared error of nonlinear wavelet-based density estimators. Annals of Statistics, 23(3):905–928, June 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176324628.

Hall:1996:DBC

[HP96] Peter Hall and Jörg Polzehl. On the dimension of the boundary of clumps in a multi-type Boolean model. Annals of Statistics, 24(4):1521–1535, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1032298281.

Henze:1999:MRT

[HP99] Norbert Henze and Mathew D. Penrose. On the multivariate runs test. Annals of Statistics, 27(1):290–298, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1018031112.

Hedayat:1992:CAO

[HPS92a] A. S. Hedayat, Kewei Pu, and John Stufken. On the construction of asymmetrical orthogonal arrays. Annals of Statistics, 20(4):2142–2152, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348908.

Hollander:1992:NMI

[HPS92b] Myles Hollander, Brett Presnell, and Jayaram Sethuraman. Nonparametric methods for imperfect repair models. Annals of Statistics, 20(2):879–896, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348661.

Heitjan:1991:ICD

[HR91] Daniel F. Heitjan and Donald B. Rubin. Ignorability and coarse data. Annals of Statistics, 19(4):2244–2253, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348396.

Hossjer:1995:ATK

[HR95] Ola Hossjer and David Ruppert. Asymptotics for the transformation kernel density estimator. Annals of Statistics, 23(4): 1198–1222, August 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324705.

Hall:1998:GPA

[HR98] Peter Hall and Marc Raimondo. On global performance of approximations to smooth curves using gridded data. Annals of Statistics, 26(6):2206–2217, December 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691467.

Hobert:1999:EMC

[HR99] James P. Hobert and C. P. Robert. Eaton’s Markov chain, its conjugate partner and P-admissibility. Annals of Statistics, 27 (1):361–373, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1018031115.

Hossjer:1994:ARM

[HRC94] Ola Hossjer, Peter J. Rousseeuw, and Christophe Croux. Asymptotics of the repeated median slope estimator. Annals of Statistics, 22(3):1478–1501, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325638.

He:1992:RDE

[HS92] Xuming He and Douglas G. Simpson. Robust direction estimation. Annals of Statistics, 20(1):351–369, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348526.

He:1993:LBC

[HS93] Xuming He and Douglas G. Simpson. Lower bounds for contamination bias: Globally minimax versus locally linear estimation. Annals of Statistics, 21(1):314–337, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349028.

Hengartner:1995:FSC

[HS95] Nicolas W. Hengartner and Philip B. Stark. Finite-sample confidence envelopes for shape-restricted densities. Annals of Statistics, 23(2):525–550, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176324534.

He:1996:GBR

[HS96a] Xuming He and Qi-Man Shao. A general Bahadur representation of M-estimators and its application to linear regression with nonstochastic designs. Annals of Statistics, 24(6):2608– 2630, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032181172.

Horvath:1996:LTM

[HS96b] Lajos Horváth and Qi-Man Shao. Limit theorem for maximum of standardized U-statistics with an application. Annals of Statistics, 24(5):2266–2279, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1069362321.

Hsing:1991:TIE

[Hsi91] Tailen Hsing. On tail index estimation using dependent data. Annals of Statistics, 19(3):1547–1569, September 1991. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348261.

Hsing:1993:EIE

[Hsi93] Tailen Hsing. Extremal index estimation for a weakly dependent stationary sequence. Annals of Statistics, 21(4):2043– 2071, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349409.

Hsieh:1996:EPA

[Hsi96] Fushing Hsieh. Empirical process approach in a two-sample location-scale model with censored data. Annals of Statistics, 24(6):2705–2719, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1032181176.

Hsieh:1997:EHL

[Hsi97] Fushing Hsieh. Estimations in homoscedastic linear regression models with censored data: an empirical process approach. Annals of Statistics, 25(6):2665–2681, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1030741090.

Hsing:1999:NNI

[Hsi99] Tailen Hsing. Nearest neighbor inverse regression. Annals of Statistics, 27(2):697–731, April 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1018031213.

He:1996:ODR

[HSS96] Zhuoqiong He, William J. Studden, and Dongchu Sun. Optimal designs for rational models. Annals of Statistics, 24(5): 2128–2147, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362314.

Hardle:1997:SSI

[HSS97a] W. Härdle, V. Spokoiny, and S. Sperlich. Semiparametric single index versus fixed link function modelling. Annals of Statistics, 25(1):212–243, February 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1034276627.

Hedayat:1997:CEO

[HSS97b] S. Hedayat, John Stufken, and Guoqin Su. On the construction and existence of orthogonal arrays with three levels and indexes 1 and 2. Annals of Statistics, 25(5):2044–2053, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1069362386.

Hosoya:1982:CLT

[HT82] Yuzo Hosoya and Masanobu Taniguchi. A central limit theorem for stationary processes and the parameter estimation of linear processes. Annals of Statistics, 10(1):132–153, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/

1176345696. See correction [HT93].

Hosoya:1993:CCL

[HT93] Yuzo Hosoya and Masanobu Taniguchi. Correction: “A Central Limit Theorem for Stationary Processes and the Parameter Estimation of Linear Processes. Annals of Statistics, 21 (2):1115–1117, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349167. See [HT82].

Hsieh:1996:NSE

[HT96] Fushing Hsieh and Bruce W. Turnbull. Nonparametric and semiparametric estimation of the receiver operating characteristic curve. Annals of Statistics, 24(1):25–40, February 1996. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1033066197.

Hall:1997:IMN

[HT97] Peter Hall and Berwin A. Turlach. Interpolation methods for nonlinear wavelet regression with irregularly spaced design. Annals of Statistics, 25(5):1912–1925, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1069362378.

Hastie:1998:CPC

[HT98] Trevor Hastie and Robert Tibshirani. Classification by pairwise coupling. Annals of Statistics, 26(2):451–471, April 1998. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1028144844.

Huang:1996:EEP

[Hua96] Jian Huang. Efficient estimation for the proportional hazards model with interval censoring. Annals of Statistics, 24 (2):540–568, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032894452.

Huang:1997:APN

[Hua97] Jian Huang. Asymptotic properties of the NPMLE of a distribution function based on ranked set samples. Annals of Statistics, 25(3):1036–1049, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1069362737.

Huang:1998:PEM

[Hua98] Jianhua Z. Huang. Projection estimation in multiple regression with application to functional ANOVA models. Annals of Statistics, 26(1):242–272, February 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1030563984.

Huang:1999:EEP

[Hua99] Jian Huang. Efficient estimation of the partly linear additive Cox model. Annals of Statistics, 27(5):1536–1563, October 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1017939141.

Huggins:1996:RIV

[Hug96] R. M. Huggins. Robust inference for variance components models for single trees of cell lineage data. Annals of Statistics, 24 (3):1145–1160, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032526961.

Huwang:1996:AHC

[Huw96] Longcheen Huwang. Asymptotically honest confidence sets for structural errors-in-variables models. Annals of Statistics, 24(4):1536–1546, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1032298282.

Hart:1990:DDB

[HV90] Jeffrey D. Hart and Philippe Vieu. Data-driven bandwidth choice for density estimation based on dependent data. Annals of Statistics, 18(2):873–890, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347630.

Heckman:1991:MBE

[HW91] Nancy E. Heckman and Michael Woodroofe. Minimax Bayes estimation in nonparametric regression. Annals of Statistics, 19(4):2003–2014, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348383.

Hu:1994:MPP

[HW94] Xiaomi Hu and F. T. Wright. Monotonicity properties of the power functions of likelihood ratio tests for normal mean hypotheses constrained by a linear space and a cone. Annals of Statistics, 22(3):1547–1554, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325642.

Hall:1997:EET

[HW97a] Peter Hall and Ishay Weissman. On the estimation of extreme tail probabilities. Annals of Statistics, 25(3):1311–1326, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1069362750.

He:1997:CDC

[HW97b] Xuming He and Gang Wang. Convergence of depth contours for multivariate datasets. Annals of Statistics, 25(2):495–504, April 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1031833661.

Hall:1965:RBS

[HWG65] W. J. Hall, R. A. Wijsman, and J. K. Ghosh. The relationship between sufficiency and invariance with applications in sequential analysis. Annals of Mathematical Statistics, 36(2):575–614, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177700169. See correction [HWG95].

Hall:1995:CRB

[HWG95] W. J. Hall, R. A. Wijsman, and J. K. Ghosh. Correction: “The Relationship Between Sufficiency and Invariance with Applications in Sequential Analysis”. Annals of Statistics, 23(2):705, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176324543.

He:1998:ETP

[HY98a] Shuyuan He and Grace L. Yang. Estimation of the truncation probability in the random truncation model. Annals of Statistics, 26(3):1011–1027, June 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691086.

He:1998:SLU

[HY98b] Shuyuan He and Grace L. Yang. The strong law under random truncation. Annals of Statistics, 26(3):992–1010, June 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1024691085.

Hedayat:1990:OTP

[HZ90] A. Hedayat and W. Zhao. Optimal two-period repeated measurements designs. Annals of Statistics, 18(4):1805–1816, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176347879. See correction [HZ92].

Hedayat:1992:COT

[HZ92] A. Hedayat and W. Zhao. Correction: “Optimal Two-Period Repeated Measurements Designs”. Annals of Statistics, 20 (1):619, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348548. See [HZ90].

Huang:1994:EMD

[HZ94] Youping Huang and Cun-Hui Zhang. Estimating a monotone density from censored observations. Annals of Statistics, 22(3):1256–1274, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325628.

Ibragimov:1991:ANF

[IK91] I. A. Ibragimov and R. Z. Khas’minskii. Asymptotically normal families of distributions and efficient estimation. Annals of Statistics, 19(4):1681–1724, December 1991. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348367.

Inglot:1997:DDS

[IKL97] Tadeusz Inglot, Wilbert C. M. Kallenberg, and Teresa Ledwina. Data driven smooth tests for composite hypotheses. Annals of Statistics, 25(3):1222–1250, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1069362746.

Inglot:1990:PED

[IL90] Tadeusz Inglot and Teresa Ledwina. On probabilities of excessive deviations for Kolmogorov–Smirnov, Cramér–von Mises and chi-square statistics. Annals of Statistics, 18(3):1491– 1495, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347764.

Inglot:1996:AOD

[IL96] Tadeusz Inglot and Teresa Ledwina. Asymptotic optimality of data-driven Neyman’s tests for uniformity. Annals of Statistics, 24(5):1982–2019, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1069362306.

Ioannidis:1994:BCT

[Ioa94] Evangelos E. Ioannidis. On the behavior of a Capon-type spectral density estimator. Annals of Statistics, 22(4):2089– 2114, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325773.

Imbens:1997:BIC

[IR97] Guido W. Imbens and Donald B. Rubin. Bayesian inference for causal effects in randomized experiments with noncompliance. Annals of Statistics, 25(1):305–327, February 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1034276631.

Ishwaran:1996:IRE

[Ish96a] Hemant Ishwaran. Identifiability and rates of estimation for scale parameters in location mixture models. Annals of Statistics, 24(4):1560–1571, August 1996. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1032298284.

Ishwaran:1996:URE

[Ish96b] Hemant Ishwaran. Uniform rates of estimation in the semiparametric Weibull mixture model. Annals of Statistics, 24(4): 1572–1585, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032298285.

Ishwaran:1998:EPC

[Ish98] Hemant Ishwaran. Exponential posterior consistency via generalized Pólya urn schemes in finite semiparametric mixtures. Annals of Statistics, 26(6):2157–2178, December 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1024691465.

Ishwaran:1999:ISM

[Ish99] Hemant Ishwaran. Information in semiparametric mixtures of exponential families. Annals of Statistics, 27(1):159–177, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1018031106.

Jacroux:1994:CTR

[Jac94] Mike Jacroux. On the construction of trend resistant mixed level factorial run orders. Annals of Statistics, 22(2):904–916, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176325502.

James:1997:SCW

[Jam97] Lancelot F. James. A study of a class of weighted bootstraps for censored data. Annals of Statistics, 25(4):1595–1621, August 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1031594733.

Janssen:1991:CRT

[Jan91] Arnold Janssen. Conditional rank tests for randomly censored data. Annals of Statistics, 19(3):1434–1456, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348255.

Jones:1990:APG

[JC90] Michael P. Jones and John Crowley. Asymptotic properties of a general class of nonparametric tests for survival analysis. Annals of Statistics, 18(3):1203–1220, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347747.

Junker:1997:CMU

[JE97] Brian W. Junker and Jules L. Ellis. A characterization of monotone unidimensional latent variable models. Annals of Statistics, 25(3):1327–1343, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1069362751.

Jiang:1996:REA

[Jia96] Jiming Jiang. REML estimation: asymptotic behavior and related topics. Annals of Statistics, 24(1):255–286, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1033066209.

Jiang:1997:WCM

[Jia97] Jiming Jiang. Wald consistency and the method of sieves in REML estimation. Annals of Statistics, 25(4):1781–1803, August 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1031594742.

Jiang:1999:CIA

[Jia99] Jiming Jiang. Conditional inference about generalized linear mixed models. Annals of Statistics, 27(6):1974–2007, December 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1017939247.

Jin:1992:ESP

[Jin92] Kun Jin. Empirical smoothing parameter selection in adaptive estimation. Annals of Statistics, 20(4):1844–1874, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348892.

Johansen:1990:HTV

[JJ90] Soren Johansen and Iain M. Johnstone. Hotelling’s theorem on the volume of tubes: Some illustrations in simultaneous inference and data analysis. Annals of Statistics, 18(2):652–684, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347620.

Jacobsen:1995:CRG

[JK95] Martin Jacobsen and Niels Keiding. Coarsening at random in general sample spaces and random censoring in continuous time. Annals of Statistics, 23(3):774–786, June 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176324622.

Johnstone:1992:MEC

[JM92] Iain M. Johnstone and K. Brenda MacGibbon. Minimax estimation of a constrained Poisson vector. Annals of Statistics, 20(2):807–831, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348658.

Jones:1991:SRB

[JMP91] M. C. Jones, J. S. Marron, and B. U. Park. A simple root n bandwidth selector. Annals of Statistics, 19(4):1919–1932, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176348378.

Johnstone:1990:DAP

[Joh90] Iain M. Johnstone. Discussion: An ancillarity paradox which appears in multiple linear regression. Annals of Statistics, 18 (2):523–525, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347610.

Johnstone:1994:MES

[Joh94] Iain M. Johnstone. On minimax estimation of a sparse normal mean vector. Annals of Statistics, 22(1):271–289, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325368.

Jones:1992:SLG

[Jon92] Lee K. Jones. A simple lemma on greedy approximation in Hilbert space and convergence rates for projection pursuit regression and neural network training. Annals of Statistics, 20 (1):608–613, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348546.

Jensen:1999:ANM

[JP99] Jens Ledet Jensen and Niels Væver Petersen. Asymptotic normality of the maximum likelihood estimator in state space models. Annals of Statistics, 27(2):514–535, April 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1018031205.

Jing:1994:SAM

[JR94] Bingyi Jing and John Robinson. Saddlepoint approximations for marginal and conditional probabilities of transformed variables. Annals of Statistics, 22(3):1115–1132, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325620.

Johnstone:1990:SEP

[JS90] Iain M. Johnstone and Bernard W. Silverman. Speed of estimation in positron emission tomography and related inverse problems. Annals of Statistics, 18(1):251–280, March 1990. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347500.

Jiang:1999:HME

[JT99] Wenxin Jiang and Martin A. Tanner. Hierarchical mixtures-of-experts for exponential family regression models: approximation and maximum likelihood estimation. Annals of Statistics, 27(3):987–1011, June 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1018031265.

Junker:1993:CAE

[Jun93] Brian W. Junker. Conditional association, essential independence and monotone unidimensional item response models. Annals of Statistics, 21(3):1359–1378, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349262.

James:1993:MWS

[JV93] Alan T. James and William N. Venables. Matrix weighting of several regression coefficient vectors. Annals of Statistics, 21 (2):1093–1114, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349166.

Jongbloed:1998:IIE

[JvEvZ98] Geurt Jongbloed, Bert van Es, and Martien van Zuijlen. Isotonic inverse estimators for nonparametric deconvolution. Annals of Statistics, 26(6):2395–2406, December 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1024691476.

Jing:1996:EEL

[JW96] Bing-Yi Jing and Andrew T. A. Wood. Exponential empirical likelihood is not Bartlett correctable. Annals of Statistics, 24 (1):365–369, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1033066214.

Kano:1991:ADN

[Kan91] Yutaka Kano. The asymptotic distribution of a noniterative estimator in exploratory factor analysis. Annals of Statistics, 19(1):272–282, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347981.

Kanazawa:1992:OVC

[Kan92] Yuichiro Kanazawa. An optimal variable Cell histogram based on the sample spacings. Annals of Statistics, 20(1):291–304, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176348523.

Karunamuni:1996:ORC

[Kar96] Rohana J. Karunamuni. Optimal rates of convergence of empirical Bayes tests for the continuous one-parameter exponential family. Annals of Statistics, 24(1):212–231, February 1996. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1033066207.

Kunsch:1993:CUL

[KBH93] H. Kunsch, J. Beran, and F. Hampel. Contrasts under longrange correlations. Annals of Statistics, 21(2):943–964, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349159.

Koo:1998:LDE

[KC98] Ja-Yong Koo and Han-Yeong Chung. Log-density estimation in linear inverse problems. Annals of Statistics, 26(1): 335–362, February 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1030563989.

Karlin:1990:SCH

[KDK90] Samuel Karlin, Amir Dembo, and Tsutomu Kawabata. Statistical composition of high-scoring segments from molecular sequences. Annals of Statistics, 18(2):571–581, June 1990. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347616.

Kneip:1995:MEN

[KE95] Alois Kneip and Joachim Engel. Model estimation in nonlinear regression under shape invariance. Annals of Statistics, 23 (2):551–570, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324535.

Kemperman:1991:MLN

[Kem91] J. H. B. Kemperman. Mixtures with a limited number of modal intervals. Annals of Statistics, 19(4):2120–2144, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348389.

Keiding:1990:RTM

[KG90] Niels Keiding and Richard D. Gill. Random truncation models and Markov processes. Annals of Statistics, 18(2):582–602, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347617.

Korn:1991:NLS

[KG91] Edward L. Korn and Barry I. Graubard. A note on the large sample properties of linearization, jackknife and balanced repeated replication methods for stratified samples. Annals of Statistics, 19(4):2275–2279, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348400.

Kneip:1992:STA

[KG92] Alois Kneip and Theo Gasser. Statistical tools to analyze data representing a sample of curves. Annals of Statistics, 20(3):1266–1305, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348769.

Khan:1991:MCE

[Kha91] Rasul A. Khan. On the monotonicity of a certain expectation. Annals of Statistics, 19(3):1677–1680, September 1991. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348272.

Khmaladze:1993:GFP

[Khm93] E. V. Khmaladze. Goodness of fit problem and scanning innovation martingales. Annals of Statistics, 21(2):798–829, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349152.

Kim:1992:MTM

[Kim92] Seong-Ju Kim. The metrically trimmed mean as a robust estimator of location. Annals of Statistics, 20(3):1534–1547, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1176348783.

Kim:1993:CSG

[Kim93] Joo Han Kim. Chi-square goodness-of-fit tests for randomly censored data. Annals of Statistics, 21(3):1621–1639, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349275.

Kim:1998:DDE

[Kim98] Peter T. Kim. Deconvolution density estimation on SO(N). Annals of Statistics, 26(3):1083–1102, June 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1024691089.

Kim:1999:NBE

[Kim99] Yongdai Kim. Nonparametric Bayesian estimators for counting processes. Annals of Statistics, 27(2):562–588, April 1999. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1018031207.

Kitamura:1997:ELM

[Kit97] Yuichi Kitamura. Empirical likelihood methods with weakly dependent processes. Annals of Statistics, 25(5):2084–2102, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1069362388.

Kurata:1996:LUB

[KK96] Hiroshi Kurata and Takeaki Kariya. Least upper bound for the covariance matrix of a generalized least squares estimator in regression with applications to a seemingly unrelated regression model and a heteroscedastic model. Annals of Statistics, 24(4): 1547–1559, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032298283.

Kedem:1991:MGF

[KL91] Benjamin Kedem and Ta-Hsin Li. Monotone gain, firstorder autocorrelation and zero-crossing rate. Annals of Statistics, 19(3):1672–1676, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348271.

Khasminskii:1992:SFP

[KL92] Rafail Z. Khas’minskii and Betty V. Lazareva. On some filtration procedure for jump Markov process observed in white Gaussian noise. Annals of Statistics, 20(4):2153–2160, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348909.

Kallenberg:1995:CMC

[KL95] Wilbert C. M. Kallenberg and Teresa Ledwina. Consistency and Monte Carlo simulation of a data driven version of smooth goodness-of-fit tests. Annals of Statistics, 23(5):1594–1608, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176324315.

Kleinberg:1996:ORS

[Kle96] E. M. Kleinberg. An overtraining-resistant stochastic modeling method for pattern recognition. Annals of Statistics, 24(6): 2319–2349, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032181157.

Kong:1997:PCM

[KLW97] Augustine Kong, Jun S. Liu, and Wing Hung Wong. The properties of the cross-match estimate and split sampling. Annals of Statistics, 25(6):2410–2432, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1030741079.

Kolassa:1990:ESL

[KM90] John E. Kolassa and Peter McCullagh. Edgeworth series for lattice distributions. Annals of Statistics, 18(2):981–985, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347637.

Kurotschka:1992:ORE

[KM92] Viktor Kurotschka and Christine Muller. Optimum robust estimation of linear aspects in conditionally contaminated linear models. Annals of Statistics, 20(1):331–350, March 1992. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348525.

Kluppelberg:1996:IPS

[KM96] Claudia Klüppelberg and Thomas Mikosch. The integrated periodogram for stable processes. Annals of Statistics, 24(5): 1855–1879, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362301.

Koshevoy:1997:ZTM

[KM97] Gleb Koshevoy and Karl Mosler. Zonoid trimming for multivariate distributions. Annals of Statistics, 25(5):1998–2017, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1069362382.

Klein:1998:CEI

[KMZ98] André Klein, Guy Mélard, and Toufik Zahaf. Computation of the exact information matrix of Gaussian dynamic regression time series models. Annals of Statistics, 26(4):1636–1650, August 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1024691256.

Kreiss:1998:RTI

[KN98] Jens-Peter Kreiss and Michael H. Neumann. Regressiontype inference in nonparametric autoregression. Annals of Statistics, 26(4):1570–1613, August 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691254.

Kneip:1994:NEC

[Kne94a] Alois Kneip. Nonparametric estimation of common regressors for similar curve data. Annals of Statistics, 22(3):1386– 1427, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325634.

Kneip:1994:OLS

[Kne94b] Alois Kneip. Ordered linear smoothers. Annals of Statistics, 22(2):835–866, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325498.

Knight:1998:LDR

[Kni98] Keith Knight. Limiting distributions for L_{1} regression estimators under general conditions. Annals of Statistics, 26 (2):755–770, April 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1028144858.

Ko:1992:REC

[Ko92] Daijin Ko. Robust estimation of the concentration parameter of the von Mises–Fisher distribution. Annals of Statistics, 20 (2):917–928, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348663.

Koul:1994:WCR

[KO94] Hira L. Koul and Mina Ossiander. Weak convergence of randomly weighted dependent residual empiricals with applications to autoregression. Annals of Statistics, 22(1):540–562, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176325383.

Koch:1996:APM

[Koc96] Inge Koch. On the asymptotic performance of median smoothers in image analysis and nonparametric regression. Annals of Statistics, 24(4):1648–1666, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1032298289.

Kolassa:1996:HOA

[Kol96] John E. Kolassa. Higher-order approximations to conditional distribution functions. Annals of Statistics, 24(1):353–364, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1033066213.

Koltchinskii:1997:MEC

[Kol97] V. I. Koltchinskii. M-estimation, convexity and quantiles. Annals of Statistics, 25(2):435–477, April 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1031833659.

Kolassa:1999:CAG

[Kol99] John E. Kolassa. Convergence and accuracy of Gibbs sampling for conditional distributions in generalized linear models. Annals of Statistics, 27(1):129–142, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1018031104.

Konishi:1991:NTB

[Kon91] Sadanori Konishi. Normalizing transformations and bootstrap confidence intervals. Annals of Statistics, 19(4):2209–2225, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176348393.

Koning:1992:ASI

[Kon92] Alex J. Koning. Approximation of stochastic integrals with applications to goodness-of-fit tests. Annals of Statistics, 20 (1):428–454, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348531.

Koning:1994:ABM

[Kon94] Alex J. Koning. Approximation of the basic martingale. Annals of Statistics, 22(2):565–579, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176325485.

Koo:1993:ORC

[Koo93] Ja-Yong Koo. Optimal rates of convergence for nonparametric statistical inverse problems. Annals of Statistics, 21(2):590–599, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176349138.

Korostelev:1996:MCN

[Kor96] Alexander Korostelev. A minimaxity criterion in nonparametric regression based on large-deviations probabilities. Annals of Statistics, 24(3):1075–1083, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1032526957.

Koster:1996:MPN

[Kos96] J. T. A. Koster. Markov properties of nonrecursive causal models. Annals of Statistics, 24(5):2148–2177, October 1996. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1069362315.

Koul:1996:ASE

[Kou96] Hira L. Koul. Asymptotics of some estimators and sequential residual empiricals in nonlinear time series. Annals of Statistics, 24(1):380–404, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1033066216.

Koyak:1990:CAT

[Koy90] Robert Koyak. Consistency for ACE-type methods. Annals of Statistics, 18(2):742–757, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347623.

Kim:1990:CRA

[KP90a] Jeankyung Kim and David Pollard. Cube root asymptotics. Annals of Statistics, 18(1):191–219, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347498.

Koul:1990:WAE

[KP90b] Hira L. Koul and Georg Ch. Pflug. Weakly adaptive estimators in explosive autoregression. Annals of Statistics, 18(2):939–960, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347634.

Kerkyacharian:1996:ENF

[KP96] Gérard Kerkyacharian and Dominique Picard. Estimating nonquadratic functionals of a density using Haar wavelets. Annals of Statistics, 24(2):485–507, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1032894450.

Konev:1997:GEM

[KP97] V. Konev and S. Pergamenshchikov. On guaranteed estimation of the mean of an autoregressive process. Annals of Statistics, 25(5):2127–2163, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1069362391.

Krieger:1999:DCR

[KPY99] Abba M. Krieger, Moshe Pollak, and Benjamin Yakir. Detecting a change in regression: first-order optimality. Annals of Statistics, 27(6):1896–1913, December 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1017939243.

Kreiss:1990:TLH

[Kre90] Jens-Peter Kreiss. Testing linear hypotheses in autoregressions. Annals of Statistics, 18(3):1470–1482, September 1990. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347762.

Koul:1993:EPA

[KS93] Hira L. Koul and A. K. Md. E. Saleh. R-estimation of the parameters of autoregressive [Ar(p)] models. Annals of Statistics, 21(1):534–551, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349041.

Kass:1994:SDM

[KS94] Robert E. Kass and Elizabeth H. Slate. Some diagnostics of maximum likelihood and posterior nonnormality. Annals of Statistics, 22(2):668–695, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325490.

Koul:1995:AQR

[KS95] Hira L. Koul and A. K. Md. E. Saleh. Autoregression quantiles and related rank-scores processes. Annals of Statistics, 23 (2):670–689, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324541.

Kokonendji:1996:DSD

[KS96a] Célestin C. Kokonendji and V. Seshadri. On the determinant of the second derivative of a Laplace transform. Annals of Statistics, 24(4):1813–1827, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1032298297.

Koul:1996:AER

[KS96b] Hira L. Koul and Anton Schick. Adaptive estimation in a random coefficient autoregressive model. Annals of Statistics, 24 (3):1025–1052, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032526954.

Koul:1997:AEM

[KS97] Hira L. Koul and Donatas Surgailis. Asymptotic expansion of M-estimators with long-memory errors. Annals of Statistics, 25(2):818–850, April 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1031833675.

Koul:1999:NMC

[KS99a] Hira L. Koul and Winfried Stute. Nonparametric model checks for time series. Annals of Statistics, 27(1):204–236, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1018031108.

Kubokawa:1999:INE

[KS99b] T. Kubokawa and M. S. Srivastava. Improved nonnegative estimation of multivariate components of variance. Annals of Statistics, 27(6):2008–2032, December 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1017939248.

Kubokawa:1999:RIE

[KS99c] T. Kubokawa and M. S. Srivastava. Robust improvement in estimation of a covariance matrix in an elliptically contoured distribution. Annals of Statistics, 27(2):600–609, April 1999. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1018031209.

Korostelev:1995:EEM

[KST95] A. P. Korostelev, L. Simar, and A. B. Tsybakov. Efficient estimation of monotone boundaries. Annals of Statistics, 23 (2):476–489, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324531.

Kent:1991:REM

[KT91] John T. Kent and David E. Tyler. Redescending M-estimates of multivariate location and scatter. Annals of Statistics, 19(4): 2102–2119, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348388.

Kent:1996:CME

[KT96a] John T. Kent and David E. Tyler. Constrained M-estimation for multivariate location and scatter. Annals of Statistics, 24 (3):1346–1370, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032526973.

Kokoszka:1996:PEI

[KT96b] Piotr S. Kokoszka and Murad S. Taqqu. Parameter estimation for infinite variance fractional ARIMA. Annals of Statistics, 24(5):1880–1913, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1069362302.

Kubokawa:1994:UAI

[Kub94] Tatsuya Kubokawa. A unified approach to improving equivariant estimators. Annals of Statistics, 22(1):290–299, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325369.

Kuriki:1993:OST

[Kur93] Satoshi Kuriki. One-sided test for the equality of two covariance matrices. Annals of Statistics, 21(3):1379–1384, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349263.

Kushner:1997:ORM

[Kus97] H. B. Kushner. Optimal repeated measurements designs: the linear optimality equations. Annals of Statistics, 25(6):2328– 2344, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1030741075. See correction [Kus98].

Kushner:1998:COR

[Kus98] H. B. Kushner. Correction: “Optimal repeated measurements designs: the linear optimality equations” [Ann. Statist. 25 (1997), no. 6, 2328–2344; MR 99g:62108]. Annals of Statistics, 26(5):2081, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1024691369. See [Kus97].

Kadane:1996:SCC

[KW96] Joseph B. Kadane and Larry Wasserman. Symmetric, coherent, Choquet capacities. Annals of Statistics, 24(3):1250–1264, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032526967.

Klonecki:1992:AEV

[KZ92] Witold Klonecki and Stefan Zontek. Admissible estimators of variance components obtained via submodels. Annals of Statistics, 20(3):1454–1467, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348778.

Kulkarni:1995:GCR

[KZ95] Sanjeev R. Kulkarni and Ofer Zeitouni. A general classification rule for probability measures. Annals of Statistics, 23(4): 1393–1407, August 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324714.

Lahiri:1992:BEM

[Lah92] Soumendra Nath Lahiri. Bootstrapping M-estimators of a multiple linear regression parameter. Annals of Statistics, 20 (3):1548–1570, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348784.

Lahiri:1995:ABM

[Lah95] S. N. Lahiri. On the asymptotic behaviour of the moving block bootstrap for normalized sums of heavy-tail random variables. Annals of Statistics, 23(4):1331–1349, August 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176324711.

Lahiri:1999:TCB

[Lah99] S. N. Lahiri. Theoretical comparisons of block bootstrap methods. Annals of Statistics, 27(1):386–404, February 1999. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1018031117.

Lai:1994:APN

[Lai94] Tze Leung Lai. Asymptotic properties of nonlinear least squares estimates in stochastic regression models. Annals of Statistics, 22(4):1917–1930, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325764.

Lalley:1999:BNC

[Lal99] Steven P. Lalley. Beneath the noise, chaos. Annals of Statistics, 27(2):461–479, April 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1018031203.

Lang:1996:MLM

[Lan96] Joseph B. Lang. Maximum likelihood methods for a generalized class of log-linear models. Annals of Statistics, 24(2):726–752, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1032894462.

Laredo:1990:SCA

[Lar90] Catherine F. Laredo. A sufficient condition for asymptotic sufficiency of incomplete observations of a diffusion process. Annals of Statistics, 18(3):1158–1171, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347744.

Laurent:1996:EEI

[Lau96] Béatrice Laurent. Efficient estimation of integral functionals of a density. Annals of Statistics, 24(2):659–681, April 1996. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1032894458.

Lavine:1992:SAP

[Lav92] Michael Lavine. Some aspects of Pólya tree distributions for statistical modelling. Annals of Statistics, 20(3):1222–1235, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1176348767.

Lavine:1994:MAP

[Lav94] Michael Lavine. More aspects of Pólya tree distributions for statistical modelling. Annals of Statistics, 22(3):1161–1176, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1176325623.

Liu:1993:NIU

[LB93] Richard C. Liu and Lawrence D. Brown. Nonexistence of informative unbiased estimators in singular problems. Annals of Statistics, 21(1):1–13, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349012.

Liu:1995:UMP

[LB95] Huimei Liu and Roger L. Berger. Uniformly more powerful, one-sided tests for hypotheses about linear inequalities. Annals of Statistics, 23(1):55–72, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324455.

Lin:1991:TFB

[LD91] Meily Lin and A. M. Dean. Trend-free block designs for varietal and factorial experiments. Annals of Statistics, 19(3):1582– 1596, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348263.

Li:1993:GPF

[LD93] Gang Li and Hani Doss. Generalized Pearson–Fisher chi-square goodness-of-fit tests, with applications to models with life history data. Annals of Statistics, 21(2):772–797, June 1993. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349151.

Li:1995:ANR

[LD95] Gang Li and Hani Doss. An approach to nonparametric regression for life history data using local linear fitting. Annals of Statistics, 23(3):787–823, June 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324623.

Lee:1996:SEA

[Lee96] Sangyeol Lee. Sequential estimation for the autocorrelations of linear processes. Annals of Statistics, 24(5):2233–2249, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1069362319.

Lele:1993:ARL

[Lel93] Chitra Lele. Admissibility results in loss estimation. Annals of Statistics, 21(1):378–390, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349031.

Leroux:1992:CEM

[Ler92] Brian G. Leroux. Consistent estimation of a mixing distribution. Annals of Statistics, 20(3):1350–1360, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348772.

Leonard:1992:BIC

[LH92] Tom Leonard and John S. J. Hsu. Bayesian inference for a covariance matrix. Annals of Statistics, 20(4):1669–1696, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348885.

Liang:1999:ESP

[LHC99] Hua Liang, Wolfgang Härdle, and Raymond J. Carroll. Estimation in a semiparametric partially linear errors-in-variables model. Annals of Statistics, 27(5):1519–1535, October 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1017939140.

Li:1996:NLR

[LHMY96] Gang Li, Myles Hollander, Ian W. McKeague, and Jie Yang. Nonparametric likelihood ratio confidence bands for quantile functions from incomplete survival data. Annals of Statistics, 24(2):628–640, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032894455.

Li:1995:BDL

[Li95] Ta-Hsin Li. Blind deconvolution of linear systems with multilevel nonstationary inputs. Annals of Statistics, 23(2):690–704, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176324542.

Li:1996:MAC

[Li96] Bing Li. A minimax approach to consistency and efficiency for estimating equations. Annals of Statistics, 24(3):1283–1297, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1032526969.

Li:1997:NCH

[Li97] Ker-Chau Li. Nonlinear confounding in high-dimensional regression. Annals of Statistics, 25(2):577–612, April 1997. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1031833665.

Lindsay:1994:EVR

[Lin94] Bruce G. Lindsay. Efficiency versus robustness: The case for minimum Hellinger distance and related methods. Annals of Statistics, 22(2):1081–1114, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176325512.

Ling:1998:WCS

[Lin98] Shiqing Ling. Weak convergence of the sequential empirical processes of residuals in nonstationary autoregressive models. Annals of Statistics, 26(2):741–754, April 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1028144857.

Liu:1990:NDD

[Liu90] Regina Y. Liu. On a notion of data depth based on random simplices. Annals of Statistics, 18(1):405–414, March 1990. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347507.

Liu:1996:NHB

[Liu96] Jun S. Liu. Nonparametric hierarchical Bayes via sequential imputations. Annals of Statistics, 24(3):911–930, June 1996. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1032526949.

Liu:1997:IFS

[Liu97] Wei Liu. Improving the fully sequential sampling scheme of Anscombe–Chow–Robbins. Annals of Statistics, 25(5):2164– 2171, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362392.

Le:1993:RSE

[LK93] Huiling Le and David G. Kendall. The Riemannian structure of Euclidean shape spaces: A novel environment for statistics. Annals of Statistics, 21(3):1225–1271, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349259.

Lin:1999:GCF

[LK99] Chin-Yu Lin and Michael R. Kosorok. A general class of function-indexed nonparametric tests for survival analysis. Annals of Statistics, 27(5):1722–1744, October 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1017939149.

Lazzeroni:1997:MCM

[LL97a] Laura C. Lazzeroni and Kenneth Lange. Markov chains for Monte Carlo tests of genetic equilibrium in multidimensional contingency tables. Annals of Statistics, 25(1):138–168, February 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1034276624.

Lindsay:1997:SOO

[LL97b] Bruce G. Lindsay and Bing Li. On second-order optimality of the observed Fisher information. Annals of Statistics, 25(5): 2172–2199, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362393.

Ling:1998:LDM

[LL98] Shiqing Ling and W. K. Li. Limiting distributions of maximum likelihood estimators for unstable autoregressive movingaverage time series with general autoregressive heteroscedastic errors. Annals of Statistics, 26(1):84–125, February 1998. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1030563979.

Letac:1990:NRE

[LM90] Gerard Letac and Marianne Mora. Natural real exponential families with cubic variance functions. Annals of Statistics, 18(1):1–37, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347491.

Li:1994:PFC

[LM94] Bing Li and Peter McCullagh. Potential functions and conservative estimating functions. Annals of Statistics, 22(1):340–356, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176325372.

Letac:1998:QIR

[LM98] Gérard Letac and Hélène Massam. Quadratic and inverse regressions for Wishart distributions. Annals of Statistics, 26 (2):573–595, April 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1028144849.

Lepski:1997:OSA

[LMS97] O. V. Lepski, E. Mammen, and V. G. Spokoiny. Optimal spatial adaptation to inhomogeneous smoothness: an approach based on kernel estimates with variable bandwidth selectors. Annals of Statistics, 25(3):929–947, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1069362731.

Lugosi:1996:CDD

[LN96] Gábor Lugosi and Andrew Nobel. Consistency of data-driven histogram methods for density estimation and classification. Annals of Statistics, 24(2):687–706, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1032894460.

Lugosi:1999:AMS

[LN99] Gábor Lugosi and Andrew B. Nobel. Adaptive model selection using empirical complexities. Annals of Statistics, 27(6): 1830–1864, December 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1017939241.

Lo:1992:SPG

[Lo92] Shaw-Hwa Lo. From the species problem to a general coverage problem via a new interpretation. Annals of Statistics, 20(2): 1094–1109, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348672.

Lo:1993:BBC

[Lo93a] Albert Y. Lo. A Bayesian bootstrap for censored data. Annals of Statistics, 21(1):100–123, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349017.

Lo:1993:BMW

[Lo93b] Albert Y. Lo. A Bayesian method for weighted sampling. Annals of Statistics, 21(4):2138–2148, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349414.

Loader:1992:LLM

[Loa92] Clive R. Loader. A log-linear model for a Poisson process change point. Annals of Statistics, 20(3):1391–1411, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348774.

Loader:1996:CPE

[Loa96a] Clive R. Loader. Change point estimation using nonparametric regression. Annals of Statistics, 24(4):1667–1678, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032298290.

Loader:1996:LLD

[Loa96b] Clive R. Loader. Local likelihood density estimation. Annals of Statistics, 24(4):1602–1618, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1032298287.

Loader:1999:BSC

[Loa99] Clive R. Loader. Bandwidth selection: classical or plug-in? Annals of Statistics, 27(2):415–438, April 1999. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1018031201.

Lockhart:1991:OTI

[Loc91] R. A. Lockhart. Overweight tails are inefficient. Annals of Statistics, 19(4):2254–2258, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348397.

Lohr:1990:AME

[Loh90] Sharon L. Lohr. Accurate multivariate estimation using triple sampling. Annals of Statistics, 18(4):1615–1633, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347869.

Loh:1991:ECMa

[Loh91a] Wei-Liem Loh. Estimating covariance matrices. Annals of Statistics, 19(1):283–296, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347982.

Loh:1991:ECMb

[Loh91b] Wei-Liem Loh. Estimating the common mean of two multivariate normal distributions. Annals of Statistics, 19(1):297–313, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347983.

Loh:1996:CCL

[Loh96a] Wei-Liem Loh. A combinatorial central limit theorem for randomized orthogonal array sampling designs. Annals of Statistics, 24(3):1209–1224, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1032526964.

Loh:1996:LHS

[Loh96b] Wei-Liem Loh. On Latin hypercube sampling. Annals of Statistics, 24(5):2058–2080, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1069362310.

Lopuhaa:1992:HEE

[Lop92] Hendrik P. Lopuhaa. Highly efficient estimators of multivariate location with high breakdown point. Annals of Statistics, 20 (1):398–413, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348529.

Lopuhaa:1999:ARE

[Lop99] Hendrik P. Lopuhaä. Asymptotics of reweighted estimators of multivariate location and scatter. Annals of Statistics, 27(5): 1638–1665, October 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1017939145.

Low:1992:NEA

[Low92a] Mark G. Low. Non-existence of an adaptive estimator for the value of an unknown probability density. Annals of Statistics, 20(1):598–602, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348544.

Low:1992:RWN

[Low92b] Mark G. Low. Renormalization and white noise approximation for nonparametric functional estimation problems. Annals of Statistics, 20(1):545–554, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348538.

Low:1993:RUL

[Low93] Mark G. Low. Renormalizing upper and lower bounds for integrated risk in the white noise model. Annals of Statistics, 21 (2):577–589, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349137.

Low:1995:BVT

[Low95] Mark G. Low. Bias-variance tradeoffs in functional estimation problems. Annals of Statistics, 23(3):824–835, June 1995. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176324624.

Low:1997:NCI

[Low97] Mark G. Low. On nonparametric confidence intervals. Annals of Statistics, 25(6):2547–2554, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1030741084.

Liu:1999:MAD

[LPS99] Regina Y. Liu, Jesse M. Parelius, and Kesar Singh. Multivariate analysis by data depth: descriptive statistics, graphics and inference, (with discussion and a rejoinder by Liu and singh). Annals of Statistics, 27(3):783–858, June 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1018031260.

Lopuhaa:1991:BPA

[LR91] Hendrik P. Lopuhaa and Peter J. Rousseeuw. Breakdown points of affine equivariant estimators of multivariate location and covariance matrices. Annals of Statistics, 19(1):229–248, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347978.

Lehmann:1992:IDO

[LR92] E. L. Lehmann and J. Rojo. Invariant directional orderings. Annals of Statistics, 20(4):2100–2110, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348905.

Lu:1993:RD

[LR93] I-Li Lu and Donald Richards. Random discriminants. Annals of Statistics, 21(4):1982–2000, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349406.

Lindsay:1997:MBO

[LR97] Bruce Lindsay and Kathryn Roeder. Moment-based oscillation properties of mixture models. Annals of Statistics, 25(1): 378–386, February 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1034276634.

Liu:1992:ERR

[LS92a] Regina Y. Liu and Kesar Singh. Efficiency and robustness in resampling. Annals of Statistics, 20(1):370–384, March 1992. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348527.

Liu:1992:ODD

[LS92b] Regina Y. Liu and Kesar Singh. Ordering directional data: Concepts of data depth on circles and spheres. Annals of Statistics, 20(3):1468–1484, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348779.

Lepski:1997:OPA

[LS97] O. V. Lepski and V. G. Spokoiny. Optimal pointwise adaptive methods in nonparametric estimation. Annals of Statistics, 25(6):2512–2546, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1030741083.

Lu:1990:DAP

[Lu90] Kun-Liang Lu. Discussion: An ancillarity paradox which appears in multiple linear regression. Annals of Statistics, 18 (2):525–528, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347611.

Luo:1998:BSS

[Luo98] Zhen Luo. Backfitting in smoothing spline ANOVA. Annals of Statistics, 26(5):1733–1759, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691355.

Leslie:1993:CED

[LvE93] Julian Leslie and Constance van Eeden. On a characterization of the exponential distribution based on a type 2 right censored sample. Annals of Statistics, 21(3):1640–1647, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349276.

Lee:1999:REP

[LW99] Sangyeol Lee and Ching-Zong Wei. On residual empirical processes of stochastic regression models with applications to time series. Annals of Statistics, 27(1):237–261, February 1999. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1018031109.

Lai:1991:EDF

[LY91a] Tze Leung Lai and Zhiliang Ying. Estimating a distribution function with truncated and censored data. Annals of Statistics, 19(1):417–442, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347991.

Lai:1991:LST

[LY91b] Tze Leung Lai and Zhiliang Ying. Large sample theory of a modified Buckley–James estimator for regression analysis with censored data. Annals of Statistics, 19(3):1370–1402, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348253.

Lai:1991:RRM

[LY91c] Tze Leung Lai and Zhiliang Ying. Rank regression methods for left-truncated and right-censored data. Annals of Statistics, 19(2):531–556, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348110.

Lai:1994:MIP

[LY94] Tze Leung Lai and Zhiliang Ying. A missing information principle and M-estimators in regression analysis with censored and truncated data. Annals of Statistics, 22(3):1222–1255, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325627.

Lee:1995:AIB

[LY95a] Stephen M. S. Lee and G. Alastair Young. Asymptotic iterated bootstrap confidence intervals. Annals of Statistics, 23(4): 1301–1330, August 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324710.

Lin:1995:SAG

[LY95b] D. Y. Lin and Zhiliang Ying. Semiparametric analysis of general additive-multiplicative hazard models for counting processes. Annals of Statistics, 23(5):1712–1734, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176324320.

Li:1998:LRI

[LZ98a] Gang Li and Cun-Hui Zhang. Linear regression with interval censored data. Annals of Statistics, 26(4):1306–1327, August 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1024691244.

Li:1998:BPL

[LZ98b] Guoying Li and Jian Zhang. Breakdown properties of location M-estimators. Annals of Statistics, 26(3):1170–1189, June 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1024691093.

Machler:1995:VSP

[Mac95] Martin B. Machler. Variational solution of penalized likelihood problems and smooth curve estimation. Annals of Statistics, 23(5):1496–1517, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324309.

Maguluri:1993:SEA

[Mag93] Gangaji Maguluri. Semiparametric estimation of association in a bivariate survival function. Annals of Statistics, 21(3):1648– 1662, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349277.

Majumdar:1992:ODC

[Maj92] Dibyen Majumdar. Optimal designs for comparing test treatments with a control utilizing prior information. Annals of Statistics, 20(1):216–237, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348519.

Majumdar:1996:AOT

[Maj96a] Dibyen Majumdar. On admissibility and optimality of treatment-control designs. Annals of Statistics, 24(5):2097– 2107, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362312.

Majumdar:1996:AOA

[Maj96b] Suman Majumdar. Asymptotically optimal and admissible decision rules in compound compact Gaussian shift experiments. Annals of Statistics, 24(1):196–211, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1033066206.

Mammen:1991:ESM

[Mam91a] Enno Mammen. Estimating a smooth monotone regression function. Annals of Statistics, 19(2):724–740, June 1991. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348117.

Mammen:1991:NRU

[Mam91b] Enno Mammen. Nonparametric regression under qualitative smoothness assumptions. Annals of Statistics, 19(2):741–759, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176348118.

Mammen:1993:BWB

[Mam93] Enno Mammen. Bootstrap and wild bootstrap for high dimensional linear models. Annals of Statistics, 21(1):255–285, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349025.

Mammen:1996:EPR

[Mam96] Enno Mammen. Empirical process of residuals for highdimensional linear models. Annals of Statistics, 24(1):307– 335, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1033066211.

Martinsek:1983:SOA

[Mar83] Adam T. Martinsek. Second order approximation to the risk of a sequential procedure. Annals of Statistics, 11(3):827–836, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176346249. See corrections [Mar86, Mar91b].

Martinsek:1986:CSO

[Mar86] Adam T. Martinsek. Correction: “Second Order Approximation to the Risk of a Sequential Procedure”. Annals of Statistics, 14(1):359, March 1986. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349864. See [Mar83].

Marden:1991:SSV

[Mar91a] John I. Marden. Sensitive and sturdy p-values. Annals of Statistics, 19(2):918–934, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348128.

Martinsek:1991:CSO

[Mar91b] Adam T. Martinsek. Correction: “Second Order Approximation to the Risk of a Sequential Procedure”. Annals of Statistics, 19(4):2283, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348404. See [Mar83].

Martinsek:1992:USR

[Mar92] Adam T. Martinsek. Using stopping rules to bound the mean integrated squared error in density estimation. Annals of Statistics, 20(2):797–806, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348657.

Mashayekhi:1993:ECD

[Mas93] Mostafa Mashayekhi. On equivariance and the compound decision problem. Annals of Statistics, 21(2):736–745, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349147.

Massam:1994:EDT

[Mas94] Helene Massam. An exact decomposition theorem and a unified view of some related distributions for a class of exponential transformation models on symmetric cones. Annals of Statistics, 22(1):369–394, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176325374.

Mattner:1993:SIB

[Mat93] L. Mattner. Some incomplete but boundedly complete location families. Annals of Statistics, 21(4):2158–2162, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349416.

Mattner:1996:COS

[Mat96] L. Mattner. Complete order statistics in parametric models. Annals of Statistics, 24(3):1265–1282, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1032526968.

Meng:1991:LSP

[MBL91] Xiao-Li Meng, Yiannis Bassiakos, and Shaw-Hwa Lo. Largesample properties for a general estimator of the treatment effect in the two-sample problem with right censoring. Annals of Statistics, 19(4):1786–1812, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348371.

Mili:1996:RES

[MC96] Lamine Mili and Clint W. Coakley. Robust estimation in structured linear regression. Annals of Statistics, 24(6):2593– 2607, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032181171.

McCullagh:1996:MTC

[McC96] Peter McCullagh. Möbius transformation and Cauchy parameter estimation. Annals of Statistics, 24(2):787–808, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032894465.

Meeden:1992:ALI

[Mee92] Glen Meeden. The admissibility of the linear interpolation estimator of the population total. Annals of Statistics, 20(1): 510–522, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348535.

Meng:1994:RCE

[Men94a] Xiao-Li Meng. On the rate of convergence of the ECM algorithm. Annals of Statistics, 22(1):326–339, March 1994. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176325371.

Meng:1994:PPV

[Men94b] Xiao-Li Meng. Posterior predictive p-values. Annals of Statistics, 22(3):1142–1160, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325622.

Messer:1991:CSE

[Mes91] K. Messer. A comparison of a spline estimate to its equivalent kernel estimate. Annals of Statistics, 19(2):817–829, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348122.

Messer:1993:NCK

[MG93] Karen Messer and Larry Goldstein. A new class of kernels for nonparametric curve estimation. Annals of Statistics, 21 (1):179–195, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349021.

Mikosch:1995:PEA

[MGKA95] Thomas Mikosch, Tamar Gadrich, Claudia Kluppelberg, and Robert J. Adler. Parameter estimation for ARMA models with infinite variance innovations. Annals of Statistics, 23(1): 305–326, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324469.

Muller-Gronbach:1998:SAP

[MGR98] Thomas Müller-Gronbach and Klaus Ritter. Spatial adaption for predicting random functions. Annals of Statistics, 26(6): 2264–2288, December 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1024691470.

Milbrodt:1992:TSM

[Mil92] Hartmut Milbrodt. Testing stationarity in the mean of autoregressive processes with a nonparametric regression trend. Annals of Statistics, 20(3):1426–1440, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348776.

Minnotte:1997:NTE

[Min97] Michael C. Minnotte. Nonparametric testing of the existence of modes. Annals of Statistics, 25(4):1646–1660, August 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1031594735.

Mathew:1994:ECR

[MK94] Thomas Mathew and Subramanyam Kasala. An exact confidence region in multivariate calibration. Annals of Statistics, 22(1):94–105, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325359.

Mammen:1999:EAP

[MLN99] E. Mammen, O. Linton, and J. Nielsen. The existence and asymptotic properties of a backfitting projection algorithm under weak conditions. Annals of Statistics, 27(5):1443–1490, October 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1017939138.

Marzec:1997:GMR

[MM97] Leszek Marzec and Paweł Marzec. Generalized martingaleresidual processes for goodness-of-fit inference in Cox’s type regression models. Annals of Statistics, 25(2):683–714, April 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1031833669.

Mizera:1999:BPV

[MM99] Ivan Mizera and Christine H. Müller. Breakdown points and variation exponents of robust M-estimators in linear models. Annals of Statistics, 27(4):1164–1177, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1017938920.

Mason:1992:RSA

[MN92] David M. Mason and Michael A. Newton. A rank statistics approach to the consistency of a general bootstrap. Annals of Statistics, 20(3):1611–1624, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348787.

Massam:1998:ETL

[MN98] Hélène Massam and Erhard Neher. Estimation and testing for lattice conditional independence models on Euclidean Jordan algebras. Annals of Statistics, 26(3):1051–1082, June 1998. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1024691088.

McKeague:1995:OTI

[MNS95] Ian W. McKeague, A. M. Nikabadze, and Yanqing Sun. An omnibus test for independence of a survival time from a covariate. Annals of Statistics, 23(2):450–475, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176324530.

Mathew:1998:TRM

[MNS98a] Thomas Mathew, Kenneth Nordström, and Manoj Kumar Sharma. Tolerance regions and multiple-use confidence regions in multivariate calibration. Annals of Statistics, 26(5): 1989–2013, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1024691366.

Miller:1998:MCN

[MNS98b] Forrest R. Miller, James W. Neill, and Brian W. Sherfey. Maximin clusters for near-replicate regression lack of fit tests. Annals of Statistics, 26(4):1411–1433, August 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1024691249.

Morris:1990:DAP

[Mor90] C. N. Morris. Discussion: An ancillarity paradox which appears in multiple linear regression. Annals of Statistics, 18 (2):528–530, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347612.

Mottonen:1997:EMS

[MOT97] Jyrki Möttönen, Hannu Oja, and Juha Tienari. On the efficiency of multivariate spatial sign and rank tests. Annals of Statistics, 25(2):542–552, April 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1031833663.

Marden:1990:ISP

[MP90] John I. Marden and Michael D. Perlman. On the inadmissibility of step-down procedures for the Hotelling T^{2} problem. Annals of Statistics, 18(1):172–190, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347497.

Muller:1992:WCA

[MP92] Hans-Georg Muller and Kathryn Prewitt. Weak convergence and adaptive peak estimation for spectral densities. Annals of Statistics, 20(3):1329–1349, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348771.

Maa:1996:RMT

[MPB96] Jen-Fue Maa, Dennis K. Pearl, and Robert Bartoszyński. Reducing multidimensional two-sample data to one-dimensional interpoint comparisons. Annals of Statistics, 24(3):1069–1074, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1032526956.

Mallat:1998:ACE

[MPZ98] Stéphane Mallat, George Papanicolaou, and Zhifeng Zhang. Adaptive covariance estimation of locally stationary processes. Annals of Statistics, 26(1):1–47, February 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1030563977.

Mykland:1996:ACS

[MR96] Per A. Mykland and Jian-Jian Ren. Algorithms for computing self-consistent and maximum likelihood estimators with doubly censored data. Annals of Statistics, 24(4):1740–1764, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032298293.

Menendez:1992:DLR

[MRS92] J. A. Menendez, C. Rueda, and B. Salvador. Dominance of likelihood ratio tests under cone constraints. Annals of Statistics, 20(4):2087–2099, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348904.

Mason:1983:MKS

[MS83] David M. Mason and John H. Schuenemeyer. A modified Kolmogorov–Smirnov test sensitive to tail alternatives. Annals of Statistics, 11(3):933–946, September 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176346259. See

correction [MS92].

Menendez:1991:ALR

[MS91] J. A. Menendez and B. Salvador. Anomalies of the likelihood ratio test for testing restricted hypotheses. Annals of Statistics, 19(2):889–898, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348126. See correction [MS95].

Mason:1992:CMK

[MS92] David M. Mason and John H. Schuenemeyer. Correction: “A Modified Kolmogorov–Smirnov Test Sensitive to Tail Alternatives”. Annals of Statistics, 20(1):620–621, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/

1176348549. See [MS83].

Messig:1993:MSC

[MS93] Michael A. Messig and William E. Strawderman. Minimal sufficiency and completeness for dichotomous quantal response models. Annals of Statistics, 21(4):2149–2157, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349415.

Menendez:1995:CAL

[MS95] J. A. Menéndez and B. Salvador. Correction: “Anomalies of the likelihood ratio test for testing restricted hypotheses”. Annals of Statistics, 23(6):2292, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1034713658. See [MS91].

Moulines:1999:BLP

[MS99a] Eric Moulines and Philippe Soulier. Broadband log-periodogram regression of time series with long-range dependence. Annals of Statistics, 27(4):1415–1439, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1017938932.

Muller:1999:DVS

[MS99b] Hans-Georg Müller and Ulrich Stadtmüller. Discontinuous versus smooth regression. Annals of Statistics, 27(1):299–337, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1018031113.

Mauldin:1992:PTR

[MSW92] R. Daniel Mauldin, William D. Sudderth, and S. C. Williams. Pólya trees and random distributions. Annals of Statistics, 20(3):1203–1221, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348766.

Mitchell:1994:ABC

[MSY94] Toby Mitchell, Jerome Sacks, and Donald Ylvisaker. Asymptotic Bayes criteria for nonparametric response surface design. Annals of Statistics, 22(2):634–651, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325488.

Mammen:1995:AMR

[MT95a] E. Mammen and A. B. Tsybakov. Asymptotical minimax recovery of sets with smooth boundaries. Annals of Statistics, 23 (2):502–524, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324533.

McCabe:1995:TTS

[MT95b] B. P. M. McCabe and A. R. Tremayne. Testing a time series for difference stationarity. Annals of Statistics, 23(3):1015–1028, June 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176324634.

Mengersen:1996:RCH

[MT96] K. L. Mengersen and R. L. Tweedie. Rates of convergence of the Hastings and Metropolis algorithms. Annals of Statistics, 24(1): 101–121, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1033066201.

Mammen:1999:SDA

[MT99] Enno Mammen and Alexandre B. Tsybakov. Smooth discrimination analysis. Annals of Statistics, 27(6):1808–1829, December 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1017939240.

McKeague:1990:INC

[MU90] Ian W. McKeague and Klaus J. Utikal. Inference for a nonlinear counting process regression model. Annals of Statistics, 18(3):1172–1187, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347745.

Morgan:1991:TDD

[MU91] John P. Morgan and Nizam Uddin. Two-dimensional design for correlated errors. Annals of Statistics, 19(4):2160–2182, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348391.

Morgan:1996:OBM

[MU96] J. P. Morgan and Nizam Uddin. Optimal blocked main effects plans with nested rows and columns and related designs. Annals of Statistics, 24(3):1185–1208, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1032526963.

Mukerjee:1992:ECQ

[Muk92] Hari Mukerjee. Estimating conditional quantiles at the root of a regression function. Annals of Statistics, 20(4):2168–2176, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176348911.

Mukerjee:1993:IMC

[Muk93a] Hari Mukerjee. An improved monotone conditional quantile estimator. Annals of Statistics, 21(2):924–942, June 1993. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349158.

Mukerjee:1993:NNR

[Muk93b] Hari Mukerjee. Nearest neighbor regression with heavy-tailed errors. Annals of Statistics, 21(2):681–693, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349144.

Mukerjee:1999:OOA

[Muk99] Rahul Mukerjee. On the optimality of orthogonal array plus one run plans. Annals of Statistics, 27(1):82–93, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1018031102.

Muller:1992:CPN

[Mul92] Hans-Georg Muller. Change-points in nonparametric regression analysis. Annals of Statistics, 20(2):737–761, June 1992. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348654.

Muller:1998:ORT

[Mül98] Christine Müller. Optimum robust testing in linear models. Annals of Statistics, 26(3):1126–1146, June 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1024691091.

Murphy:1994:CPH

[Mur94] S. A. Murphy. Consistency in a proportional hazards model incorporating a random effect. Annals of Statistics, 22(2): 712–731, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325492.

Murphy:1995:ATF

[Mur95] S. A. Murphy. Asymptotic theory for the frailty model. Annals of Statistics, 23(1):182–198, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324462.

Mammen:1997:LAR

[MvdG97a] Enno Mammen and Sara van de Geer. Locally adaptive regression splines. Annals of Statistics, 25(1):387–413, February 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1034276635.

Mammen:1997:PQL

[MvdG97b] Enno Mammen and Sara van de Geer. Penalized quasilikelihood estimation in partial linear models. Annals of Statistics, 25(3):1014–1035, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1069362736.

Murphy:1997:SLR

[MvdV97] S. A. Murphy and A. W. van der Vaart. Semiparametric likelihood ratio inference. Annals of Statistics, 25(4):1471–1509, August 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1031594729.

Marron:1992:EMI

[MW92] J. S. Marron and M. P. Wand. Exact mean integrated squared error. Annals of Statistics, 20(2):712–736, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348653.

Mukerjee:1995:ESN

[MW95] Rahul Mukerjee and C. F. Jeff Wu. On the existence of saturated and nearly saturated asymmetrical orthogonal arrays. Annals of Statistics, 23(6):2102–2115, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1034713649.

Mizera:1998:NSC

[MW98] Ivan Mizera and Jon A. Wellner. Necessary and sufficient conditions for weak consistency of the median of independent but not identically distributed random variables. Annals of Statistics, 26(2):672–691, April 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1028144854.

Mukerjee:1999:BRF

[MW99a] Rahul Mukerjee and C. F. J. Wu. Blocking in regular fractional factorials: a projective geometric approach. Annals of Statistics, 27(4):1256–1271, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1017938925.

Muliere:1999:CNR

[MW99b] Pietro Muliere and Stephen Walker. A characterization of a neutral to the right prior via an extension of Johnson’s sufficientness postulate. Annals of Statistics, 27(2):589–599, April 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1018031208.

Maronna:1993:BRE

[MY93] Ricardo A. Maronna and Victor J. Yohai. Bias-robust estimates of regression based on projections. Annals of Statistics, 21(2):965–990, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349160.

Morvai:1996:NIE

[MYG96] Gusztáv Morvai, Sidney Yakowitz, and László Györfi. Nonparametric inference for ergodic, stationary time series. Annals of Statistics, 24(1):370–379, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1033066215.

Mykland:1992:AEB

[Myk92] Per Aslak Mykland. Asymptotic expansions and bootstrapping distributions for dependent variables: A martingale approach. Annals of Statistics, 20(2):623–654, June 1992. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348649.

Mykland:1994:BTI

[Myk94] Per Aslak Mykland. Bartlett type identities for martingales. Annals of Statistics, 22(1):21–38, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325355.

Mykland:1995:DL

[Myk95a] Per Aslak Mykland. Dual likelihood. Annals of Statistics, 23 (2):396–421, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324527.

Mykland:1995:MES

[Myk95b] Per Aslak Mykland. Martingale expansions and second order inference. Annals of Statistics, 23(3):707–731, June 1995. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176324617.

Mykland:1999:BIL

[Myk99] Per Aslak Mykland. Bartlett identities and large deviations in likelihood theory. Annals of Statistics, 27(3):1105–1117, June 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1018031270.

Martin:1993:BRE

[MZ93a] R. D. Martin and Ruben H. Zamar. Bias robust estimation of scale. Annals of Statistics, 21(2):991–1017, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349161.

Martin:1993:ECB

[MZ93b] R. Douglas Martin and Ruben H. Zamar. Efficiency-constrained bias-robust estimation of location. Annals of Statistics, 21(1): 338–354, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349029.

McKeague:1994:INT

[MZ94] Ian W. McKeague and Mei-Jie Zhang. Identification of nonlinear time series from first order cumulative characteristics. Annals of Statistics, 22(1):495–514, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325381.

Muller:1995:SVF

[MZ95] Hans-Georg Muller and Peng-Liang Zhao. On a semiparametric variance function model and a test for heteroscedasticity. Annals of Statistics, 23(3):946–967, June 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324630.

Mathew:1996:CCR

[MZ96] Thomas Mathew and Wenxing Zha. Conservative confidence regions in multivariate calibration. Annals of Statistics, 24 (2):707–725, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032894461.

Naiman:1990:VTN

[Nai90] Daniel Q. Naiman. Volumes of tubular neighborhoods of spherical polyhedra and statistical inference. Annals of Statistics, 18 (2):685–716, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347621.

Nakayama:1997:MCP

[Nak97] Marvin K. Nakayama. Multiple-comparison procedures for steady-state simulations. Annals of Statistics, 25(6):2433– 2450, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1030741080.

Nau:1992:IPF

[Nau92] Robert F. Nau. Indeterminate probabilities on finite sets. Annals of Statistics, 20(4):1737–1767, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348888.

Nagaraja:1994:DMC

[ND94] H. N. Nagaraja and H. A. David. Distribution of the maximum of concomitants of selected order statistics. Annals of Statistics, 22(1):478–494, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325380.

Neuhaus:1993:CRT

[Neu93] Georg Neuhaus. Conditional rank tests for the two-sample problem under random censorship. Annals of Statistics, 21(4): 1760–1779, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349396.

Neuhauser:1994:PAS

[Neu94] Claudia Neuhauser. A Poisson approximation for sequence comparisons with insertions and deletions. Annals of Statistics, 22(3):1603–1629, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325645.

Neumann:1995:ABC

[Neu95] Michael H. Neumann. Automatic bandwidth choice and confidence intervals in nonparametric regression. Annals of Statistics, 23(6):1937–1959, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1034713641.

Neumann:1998:SAD

[Neu98] Michael H. Neumann. Strong approximation of density estimators from weakly dependent observations by density estimators from independent observations. Annals of Statistics, 26(5): 2014–2048, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1024691367.

Nagaraj:1991:EPL

[NF91] Neerchal K. Nagaraj and Wayne A. Fuller. Estimation of the parameters of linear time series models subject to nonlinear restrictions. Annals of Statistics, 19(3):1143–1154, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348242.

Niemiro:1992:AED

[Nie92] Wojciech Niemiro. Asymptotics for M-estimators defined by convex minimization. Annals of Statistics, 20(3):1514–1533, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1176348782.

Nishiyama:1999:MIC

[Nis99] Yoichi Nishiyama. A maximal inequality for continuous martingales and M-estimation in a Gaussian white noise model. Annals of Statistics, 27(2):675–696, April 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1018031212.

Niu:1998:EIS

[Niu98] Xu-Feng Niu. Estimating integrals of stochastic processes using space-time data. Annals of Statistics, 26(6):2246–2263, December 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1024691469.

Nielsen:1995:KEN

[NL95] Jens P. Nielsen and Oliver B. Linton. Kernel estimation in a nonparametric marker dependent hazard model. Annals of Statistics, 23(5):1735–1748, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324321.

Nielsen:1998:SSM

[NLB98] Jens P. Nielsen, Oliver Linton, and Peter J. Bickel. On a semiparametric survival model with flexible covariate effect. Annals of Statistics, 26(1):215–241, February 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1030563983.

Naik-Nimbalkar:1994:VBB

[NNR94] U. V. Naik-Nimbalkar and M. B. Rajarshi. Validity of blockwise bootstrap for empirical processes with stationary observations. Annals of Statistics, 22(2):980–994, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325507.

Nogales:1996:SRS

[NO96] A. G. Nogales and J. A. Oyola. Some remarks on sufficiency, invariance and conditional independence. Annals of Statistics, 24(2):906–909, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032894473.

Nobel:1996:HRE

[Nob96] Andrew Nobel. Histogram regression estimation using datadependent partitions. Annals of Statistics, 24(3):1084–1105, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1032526958.

Nobel:1999:LCR

[Nob99] Andrew B. Nobel. Limits to classification and regression estimation from ergodic processes. Annals of Statistics, 27(1): 262–273, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1018031110.

Nabeya:1994:APA

[NT94] Seiji Nabeya and Katsuto Tanaka. Acknowledgment of priority: Asymptotic theory of a test for the constancy of regression coefficients against the random walk alternative. Annals of Statistics, 22(1):563, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325384.

Nussbaum:1996:AED

[Nus96] Michael Nussbaum. Asymptotic equivalence of density estimation and Gaussian white noise. Annals of Statistics, 24(6): 2399–2430, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032181160.

Neumann:1997:WTA

[NvS97] Michael H. Neumann and Rainer von Sachs. Wavelet thresholding in anisotropic function classes and application to adaptive estimation of evolutionary spectra. Annals of Statistics, 25 (1):38–76, February 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1034276621.

Naiman:1992:IEB

[NW92] Daniel Q. Naiman and Henry P. Wynn. Inclusion-exclusion-Bonferroni identities and inequalities for discrete tube-like problems via Euler characteristics. Annals of Statistics, 20 (1):43–76, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348512.

Naiman:1997:ATI

[NW97] Daniel Q. Naiman and Henry P. Wynn. Abstract tubes, improved inclusion-exclusion identities and inequalities and importance sampling. Annals of Statistics, 25(5):1954–1983, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1069362380.

Nicoleris:1997:RCE

[NY97] Theodoros Nicoleris and Yannis G. Yatracos. Rates of convergence of estimates, Kolmogorov’s entropy and the dimensionality reduction principle in regression. Annals of Statistics, 25(6): 2493–2511, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1030741082.

Nychka:1990:APV

[Nyc90] Douglas Nychka. The average posterior variance of a smoothing spline and a consistent estimate of the average squared error. Annals of Statistics, 18(1):415–428, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347508.

Nychka:1995:SLS

[Nyc95] Douglas Nychka. Splines as local smoothers. Annals of Statistics, 23(4):1175–1197, August 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324704.

Okamoto:1991:ATS

[OAT91] Ichi Okamoto, Shun-Ichi Amari, and Kei Takeuchi. Asymptotic theory of sequential estimation: Differential geometrical approach. Annals of Statistics, 19(2):961–981, June 1991. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348131.

Oller:1995:IAS

[OC95] J. M. Oller and J. M. Corcuera. Intrinsic analysis of statistical estimation. Annals of Statistics, 23(5):1562–1581, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176324312.

Oehlert:1992:RBS

[Oeh92] Gary W. Oehlert. Relaxed boundary smoothing splines. Annals of Statistics, 20(1):146–160, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348516.

Opsomer:1997:FBA

[OR97] Jean D. Opsomer and David Ruppert. Fitting a bivariate additive model by local polynomial regression. Annals of Statistics, 25(1):186–211, February 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1034276626.

OSullivan:1991:DMA

[O’S91] Finbarr O’Sullivan. Discussion: Multivariate adaptive regression splines. Annals of Statistics, 19(1):99–102, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347968.

OSullivan:1993:NEC

[O’S93] Finbarr O’Sullivan. Nonparametric estimation in the Cox model. Annals of Statistics, 21(1):124–145, March 1993. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176349018.

OSullivan:1995:SLS

[O’S95] Finbarr O’Sullivan. A study of least squares and maximum likelihood for image reconstruction in positron emission tomography. Annals of Statistics, 23(4):1267–1300, August 1995. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176324709.

Owen:1990:ELR

[Owe90] Art Owen. Empirical likelihood ratio confidence regions. Annals of Statistics, 18(1):90–120, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347494.

Owen:1991:DMA

[Owe91a] Art Owen. Discussion: Multivariate adaptive regression splines. Annals of Statistics, 19(1):102–112, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347969.

Owen:1991:ELL

[Owe91b] Art Owen. Empirical likelihood for linear models. Annals of Statistics, 19(4):1725–1747, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348368.

Owen:1994:LSR

[Owe94] Art Owen. Lattice sampling revisited: Monte Carlo variance of means over randomized orthogonal arrays. Annals of Statistics, 22(2):930–945, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325504.

Owen:1997:SNV

[Owe97] Art B. Owen. Scrambled net variance for integrals of smooth functions. Annals of Statistics, 25(4):1541–1562, August 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1031594731.

Parner:1998:ATC

[Par98] Erik Parner. Asymptotic theory for the correlated gammafrailty model. Annals of Statistics, 26(1):183–214, February 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1030563982.

Patil:1993:LSC

[Pat93] P. N. Patil. On the least squares cross-validation bandwidth in hazard rate estimation. Annals of Statistics, 21(4):1792– 1810, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349398.

Pena:1998:SGF

[Peñ98] Edsel A. Peña. Smooth goodness-of-fit tests for composite hypothesis in hazard based models. Annals of Statistics, 26(5):

1935–1971, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1024691364.

Peres:1992:INP

[Per92] Yuval Peres. Iterating von Neumann’s procedure for extracting random bits. Annals of Statistics, 20(1):590–597, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348543;   http://www.jstor.org/stable/2242181.

Pfanzagl:1990:LDP

[Pfa90] J. Pfanzagl. Large deviation probabilities for certain nonparametric maximum likelihood estimators. Annals of Statistics, 18(4):1868–1877, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347884.

Pfanzagl:1993:IVR

[Pfa93] J. Pfanzagl. Incidental versus random nuisance parameters. Annals of Statistics, 21(4):1663–1691, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349392.

Praestgaard:1996:ATN

[PH96] Jens Thomas Præstgaard and Jian Huang. Asymptotic theory for nonparametric estimation of survival curves under order restrictions. Annals of Statistics, 24(4):1679–1716, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032298291.

Phelan:1990:BEM

[Phe90] Michael J. Phelan. Bayes estimation from a Markov renewal process. Annals of Statistics, 18(2):603–616, June 1990. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347618.

Phelan:1996:ALE

[Phe96] Michael J. Phelan. Asymptotic likelihood estimation from birth and death on a flow. Annals of Statistics, 24(3):1161–1184, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032526962.

Pinelis:1994:EPP

[Pin94] Iosif Pinelis. Extremal probabilistic problems and Hotelling’s T^{2} test under a symmetry condition. Annals of Statistics, 22 (1):357–368, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325373.

Pitts:1994:NES

[Pit94] Susan M. Pitts. Nonparametric estimation of the stationary waiting time distribution function for the GI/G/1 queue. Annals of Statistics, 22(3):1428–1446, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325635.

Portnoy:1989:AEL

[PK89] Stephen Portnoy and Roger Koenker. Adaptive L-estimation for linear models. Annals of Statistics, 17(1):362–381, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/

1176347022. See correction [PK90].

Portnoy:1990:CAE

[PK90] Stephen Portnoy and Roger Koenker. Correction: “Adaptive L-Estimation for Linear Models”. Annals of Statistics, 18(2):986, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347638. See [PK89].

Prentice:1995:SCP

[PM95] Michael J. Prentice and Kanti V. Mardia. Shape changes in the plane for landmark data. Annals of Statistics, 23(6):1960– 1974, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1034713642.

Polonik:1995:MMC

[Pol95] Wolfgang Polonik. Measuring mass concentrations and estimating density contour clusters — an excess mass approach. Annals of Statistics, 23(3):855–881, June 1995. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324626.

Polonik:1998:SCF

[Pol98] Wolfgang Polonik. The silhouette, concentration functions and ML-density estimation under order restrictions. Annals of Statistics, 26(5):1857–1877, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691360.

Polonik:1999:CGF

[Pol99] Wolfgang Polonik. Concentration and goodness-of-fit in higher dimensions: (asymptotically) distribution-free methods. Annals of Statistics, 27(4):1210–1229, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1017938922.

Poor:1998:QDE

[Poo98] H. Vincent Poor. Quickest detection with exponential penalty for delay. Annals of Statistics, 26(6):2179–2205, December 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1024691466.

Portnoy:1985:ABE

[Por85] Stephen Portnoy. Asymptotic behavior of M estimators of p regression parameters when p^{2}/n is large; II. normal approximation. Annals of Statistics, 13(4):1403–1417, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/

1176349744. See correction [Por91].

Portnoy:1991:CAB

[Por91] Stephen Portnoy. Correction: “Asymptotic Behavior of M Estimators of p Regression Parameters when p^{2}/n is Large: II. Normal Approximation”. Annals of Statistics, 19(4):2282, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176348403. See [Por85].

Portnoy:1997:LAQ

[Por97] Stephen Portnoy. Local asymptotics for quantile smoothing splines. Annals of Statistics, 25(1):414–434, February 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1034276636.

Poskitt:1992:IEC

[Pos92] D. S. Poskitt. Identification of echelon canonical forms for vector linear processes using least squares. Annals of Statistics, 20 (1):195–215, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348518.

Politis:1992:GRS

[PR92] Dimitris N. Politis and Joseph P. Romano. A general resampling scheme for triangular arrays of α-mixing random variables with application to the problem of spectral density estimation. Annals of Statistics, 20(4):1985–2007, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348899.

Politis:1994:LSC

[PR94] Dimitris N. Politis and Joseph P. Romano. Large sample confidence regions based on subsamples under minimal assumptions. Annals of Statistics, 22(4):2031–2050, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325770.

Proschan:1991:NBH

[Pro91] Michael Proschan. A note on Blackwell and Hodges (1957) and Diaconis and Graham (1981). Annals of Statistics, 19(2): 1106–1108, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348144.

Pruitt:1991:NMA

[Pru91] Ronald C. Pruitt. On negative mass assigned by the bivariate Kaplan–Meier estimator. Annals of Statistics, 19(1):443–453, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347992.

Pollak:1991:SDC

[PS91] Moshe Pollak and D. Siegmund. Sequential detection of a change in a normal mean when the initial value is unknown.

Annals of Statistics, 19(1):394–416, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347990.

Phillips:1992:ALP

[PS92] Peter C. B. Phillips and Victor Solo. Asymptotics for linear processes. Annals of Statistics, 20(2):971–1001, June 1992. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348666.

Pukelsheim:1993:ODP

[PS93] Friedrich Pukelsheim and William J. Studden. E-optimal designs for polynomial regression. Annals of Statistics, 21(1): 402–415, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349033.

Pistone:1995:IDG

[PS95] Giovanni Pistone and Carlo Sempi. An infinite-dimensional geometric structure on the space of all the probability measures equivalent to a given one. Annals of Statistics, 23(5): 1543–1561, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324311.

Pukelsheim:1991:OWE

[PT91] Friedrich Pukelsheim and Ben Torsney. Optimal weights for experimental designs on linearly independent support points. Annals of Statistics, 19(3):1614–1625, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348265.

Pensky:1999:AWE

[PV99] Marianna Pensky and Brani Vidakovic. Adaptive wavelet estimator for nonparametric density deconvolution. Annals of Statistics, 27(6):2033–2053, December 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1017939249.

Putter:1996:RCS

[PvZ96] Hein Putter and Willem R. van Zwet. Resampling: consistency of substitution estimators. Annals of Statistics, 24(6): 2297–2318, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032181156.

Putter:1998:EEE

[PvZ98] Hein Putter and Willem R. van Zwet. Empirical Edgeworth expansions for symmetric statistics. Annals of Statistics, 26(4): 1540–1569, August 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1024691253.

Qin:1993:ELB

[Qin93] Jing Qin. Empirical likelihood in biased sample problems. Annals of Statistics, 21(3):1182–1196, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349257.

Qin:1999:ELR

[Qin99] Jing Qin. Empirical likelihood ratio based confidence intervals for mixture proportions. Annals of Statistics, 27(4):1368–1384, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1017938930.

Qiu:1998:DRS

[Qiu98] Peihua Qiu. Discontinuous regression surfaces fitting. Annals of Statistics, 26(6):2218–2245, December 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691468.

Qin:1994:ELG

[QL94] Jin Qin and Jerry Lawless. Empirical likelihood and general estimating equations. Annals of Statistics, 22(1):300–325, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325370.

Raimondo:1998:MES

[Rai98] Marc Raimondo. Minimax estimation of sharp change points. Annals of Statistics, 26(4):1379–1397, August 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1024691247.

Ritov:1990:AIB

[RB90] Y. Ritov and P. J. Bickel. Achieving information bounds in non and semiparametric models. Annals of Statistics, 18 (2):925–938, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347633.

Ruppert:1994:BRK

[RC94] David Ruppert and Daren B. H. Cline. Bias reduction in kernel density estimation by smoothed empirical transformations. Annals of Statistics, 22(1):185–210, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325365.

Resnick:1997:HTM

[Res97] Sidney I. Resnick. Heavy tail modeling and teletraffic data: special invited paper. Annals of Statistics, 25(5):1805–1869, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1069362376.

Ritov:1991:ORR

[RG91] Ya’acov Ritov and Zvi Gilula. The order-restricted RC model for ordered contingency tables: Estimation and testing for fit. Annals of Statistics, 19(4):2090–2101, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348387.

Ren:1997:RME

[RG97] Jian-Jian Ren and Minggao Gu. Regression M-estimators with doubly censored data. Annals of Statistics, 25(6):2638– 2664, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1030741089.

Robinson:1997:TSR

[RH97] P. M. Robinson and F. J. Hidalgo. Time series regression with long-range dependence. Annals of Statistics, 25(1):77– 104, February 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1034276622.

Richards:1995:EAS

[Ric95] Donald St. P. Richards. Exact asymptotics for some probability distributions on compact manifolds. Annals of Statistics, 23(5): 1582–1586, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324313.

Riedel:1997:PCF

[Rie97] Kurt S. Riedel. Piecewise convex function estimation: pilot estimators. Annals of Statistics, 25(6):2592–2606, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1030741086.

Ritov:1990:AEE

[Rit90a] Y. Ritov. Asymptotic efficient estimation of the change point with unknown distributions. Annals of Statistics, 18(4):1829– 1839, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347881.

Ritov:1990:DTO

[Rit90b] Y. Ritov. Decision theoretic optimality of the Cusum procedure. Annals of Statistics, 18(3):1464–1469, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347761.

Ritov:1990:ELR

[Rit90c] Y. Ritov. Estimation in a linear regression model with censored data. Annals of Statistics, 18(1):303–328, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347502.

Ritter:1996:AOR

[Rit96] Klaus Ritter. Asymptotic optimality of regular sequence designs. Annals of Statistics, 24(5):2081–2096, October 1996. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1069362311.

Robert:1994:DMC

[Rob94a] Christian P. Robert. Discussion: Markov chains for exploring posterior distributions. Annals of Statistics, 22(4):1742– 1747, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325753.

Robinson:1994:SAL

[Rob94b] P. M. Robinson. Semiparametric analysis of long-memory time series. Annals of Statistics, 22(1):515–539, March 1994. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176325382.

Robinson:1995:GSE

[Rob95a] P. M. Robinson. Gaussian semiparametric estimation of long range dependence. Annals of Statistics, 23(5):1630–1661, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176324317.

Robinson:1995:LPR

[Rob95b] P. M. Robinson. Log-periodogram regression of time series with long range dependence. Annals of Statistics, 23(3):1048–1072, June 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176324636.

Robinson:1997:LSI

[Rob97] P. M. Robinson. Large-sample inference for nonparametric regression with dependent errors. Annals of Statistics, 25(5): 2054–2083, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362387.

Rocke:1996:RPE

[Roc96] David M. Rocke. Robustness properties of S-estimators of multivariate location and shape in high dimension. Annals of Statistics, 24(3):1327–1345, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1032526972.

Roeder:1992:SEN

[Roe92] Kathryn Roeder. Semiparametric estimation of normal mixture densities. Annals of Statistics, 20(2):929–943, June 1992. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348664.

Rojo:1992:PTO

[Roj92] Javier Rojo. A pure-tail ordering based on the ratio of the quantile functions. Annals of Statistics, 20(1):570–579, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348541.

Rosenbaum:1991:SPS

[Ros91] Paul R. Rosenbaum. Some poset statistics. Annals of Statistics, 19(2):1091–1097, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348141.

Rosenberger:1993:AIR

[Ros93] William F. Rosenberger. Asymptotic inference with responseadaptive treatment allocation designs. Annals of Statistics, 21(4):2098–2107, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349412.

Rosenthal:1995:RCG

[Ros95] Jeffrey S. Rosenthal. Rates of convergence for Gibbs sampling for variance component models. Annals of Statistics, 23 (3):740–761, June 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324619.

Rubin:1995:ECA

[RS95] Herman Rubin and Kai-Sheng Song. Exact computation of the asymptotic efficiency of maximum likelihood estimators of a discontinuous signal in a Gaussian white noise. Annals of Statistics, 23(3):732–739, June 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176324618.

Robinson:1998:BPS

[RS98] J. Robinson and Ib M. Skovgaard. Bounds for probabilities of small relative errors for empirical saddlepoint and bootstrap tail approximations. Annals of Statistics, 26(6):2369–2394, December 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1024691475.

Riccomagno:1997:LBD

[RSW97] Eva Riccomagno, Rainer Schwabe, and Henry P. Wynn. Lattice-based D-optimum design for Fourier regression. Annals of Statistics, 25(6):2313–2327, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1030741074.

Roussas:1992:ANR

[RT92a] George G. Roussas and Lanh T. Tran. Asymptotic normality of the recursive kernel regression estimate under dependence conditions. Annals of Statistics, 20(1):98–120, March 1992. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348514.

Rubin:1992:AIM

[RT92b] Donald B. Rubin and Neal Thomas. Affinely invariant matching methods with ellipsoidal distributions. Annals of Statistics, 20 (2):1079–1093, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348671.

Routledge:1997:RBT

[RT97] Rick Routledge and Min Tsao. On the relationship between two asymptotic expansions for the distribution of sample mean and its applications. Annals of Statistics, 25(5):2200–2209, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1069362394.

Rublik:1996:LDT

[Rub96] František Rublík. A large deviation theorem for the q-sample likelihood ratio statistic. Annals of Statistics, 24(5):2280–2287, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1069362322.

Rukhin:1990:KCM

[Ruk90] Andrew L. Rukhin. Kolmogorov’s contributions to mathematical statistics. Annals of Statistics, 18(3):1011–1016, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347737.

Rukhin:1994:RTM

[Ruk94] Andrew L. Rukhin. Recursive testing of multiple hypotheses: Consistency and efficiency of the Bayes rule. Annals of Statistics, 22(2):616–633, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325487.

Ruschendorf:1995:CIP

[Rus95] Ludger Ruschendorf. Convergence of the iterative proportional fitting procedure. Annals of Statistics, 23(4):1160–1174, August 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176324703.

Ruppert:1994:MLW

[RW94] D. Ruppert and M. P. Wand. Multivariate locally weighted least squares regression. Annals of Statistics, 22(3):1346– 1370, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325632.

Ryden:1994:CAN

[Ryd94] Tobias Ryden. Consistent and asymptotically normal parameter estimates for hidden Markov models. Annals of Statistics, 22(4):1884–1895, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325762.

Sato:1996:IIB

[SA96] Michikazu Sato and Masafumi Akahira. An information inequality for the Bayes risk. Annals of Statistics, 24(5):2288– 2295, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362323.

Salinelli:1998:NPC

[Sal98] Ernesto Salinelli. Nonlinear principal components I. Absolutely continuous random variables with positive bounded densities. Annals of Statistics, 26(2):596–616, April 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1028144850.

Samaranayake:1992:SWR

[Sam92] K. Samaranayake. Stay-with-a-winner rule for dependent Bernoulli bandits. Annals of Statistics, 20(4):2111–2123, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176348906.

Sarkar:1991:OAB

[Sar91] Jyotirmoy Sarkar. One-armed bandit problems with covariates. Annals of Statistics, 19(4):1978–2002, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348382.

Sarkar:1998:SPI

[Sar98] Sanat K. Sarkar. Some probability inequalities for ordered MTP_{2} random variables: a proof of the Simes conjecture. Annals of Statistics, 26(2):494–504, April 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1028144846.

Sasieni:1993:SNE

[Sas93] Peter Sasieni. Some new estimators for Cox regression. Annals of Statistics, 21(4):1721–1759, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349395.

Sriram:1991:SEB

[SBH91] T. N. Sriram, I. V. Basawa, and R. M. Huggins. Sequential estimation for branching processes with immigration. Annals of Statistics, 19(4):2232–2243, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348395.

Stefanski:1991:DBS

[SC91] Leonard A. Stefanski and Raymond J. Carroll. Deconvolutionbased score tests in measurement error models. Annals of Statistics, 19(1):249–259, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347979.

Schumaker:1991:DMA

[Sch91] Larry L. Schumaker. Discussion: Multivariate adaptive regression splines. Annals of Statistics, 19(1):112–113, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347970.

Schick:1993:EER

[Sch93] Anton Schick. On efficient estimation in regression models. Annals of Statistics, 21(3):1486–1521, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349269. See

correction [Sch95].

Schick:1995:CAE

[Sch95] Anton Schick. Correction and addendum: “On Efficient Estimation in Regression Models”. Annals of Statistics, 23(5): 1862–1863, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324327. See [Sch93].

Suen:1997:SID

[SCW97] Chung-Yi Suen, Hegang Chen, and C. F. J. Wu. Some identities on q^{n−m} designs with application to minimum aberration designs. Annals of Statistics, 25(3):1176–1188, June 1997. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1069362743.

Sinha:1999:EPR

[SdH99] Ashoke Kumar Sinha and Laurens de Haan. Estimating the probability of a rare event. Annals of Statistics, 27(2):732–759, April 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1018031214.

Sen:1996:RRS

[Sen96] Pranab Kumar Sen. Regression rank scores estimation in ANOCOVA. Annals of Statistics, 24(4):1586–1601, August 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032298286.

Seoh:1990:BET

[Seo90] Munsup Seoh. Berry–Esseen-type bounds for signed linear rank statistics with a broad range of scores. Annals of Statistics, 18(3):1483–1490, September 1990. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347763.

Sadrolhefazi:1994:FDD

[SF94] Amir Sadrolhefazi and Terrence L. Fine. Finite-dimensional distributions and tail behavior in stationary interval-valued probability models. Annals of Statistics, 22(4):1840–1870, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325760.

Shao:1993:DSF

[Sha93] Jun Shao. Differentiability of statistical functionals and consistency of the jackknife. Annals of Statistics, 21(1):61–75, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349015.

Shao:1994:SCS

[Sha94] Jun Shao. L-statistics in complex survey problems. Annals of Statistics, 22(2):946–967, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325505.

Sherman:1994:MID

[She94] Robert P. Sherman. Maximal inequalities for degenerate Uprocesses with applications to optimization estimators. Annals of Statistics, 22(1):439–459, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176325377.

Shen:1995:ORI

[She95] Larry Z. Shen. On optimal B-robust influence functions in semiparametric models. Annals of Statistics, 23(3):968–989, June 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176324631.

Shen:1997:MSP

[She97] Xiaotong Shen. On methods of sieves and penalization. Annals of Statistics, 25(6):2555–2591, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1030741085.

Shi:1991:SAR

[Shi91] Xiquan Shi. Some asymptotic results for jackknifing the sample quantile. Annals of Statistics, 19(1):496–503, March 1991. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347996.

Stone:1997:PST

[SHKT97] Charles J. Stone, Mark H. Hansen, Charles Kooperberg, and Young K. Truong. Polynomial splines and their tensor products in extended linear modeling: 1994 Wald memorial lecture. Annals of Statistics, 25(4):1371–1470, August 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1031594728.

Shorrock:1990:ICI

[Sho90] Glenn Shorrock. Improved confidence intervals for a normal variance. Annals of Statistics, 18(2):972–980, June 1990. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347636.

Shorack:1996:LRS

[Sho96] Galen R. Shorack. Linear rank statistics, finite sampling, permutation tests and Winsorizing. Annals of Statistics, 24(3): 1371–1385, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032526974.

Siegmund:1993:SCT

[Sie93] D. Siegmund. A sequential clinical trial for comparing three treatments. Annals of Statistics, 21(1):464–483, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349037.

Silverman:1996:SFP

[Sil96] Bernard W. Silverman. Smoothed functional principal components analysis by choice of norm. Annals of Statistics, 24 (1):1–24, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1033066196.

Singh:1998:BTB

[Sin98] Kesar Singh. Breakdown theory for bootstrap quantiles. Annals of Statistics, 26(5):1719–1732, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1024691354.

Skovgaard:1990:DMC

[Sko90] Ib M. Skovgaard. On the density of minimum contrast estimators. Annals of Statistics, 18(2):779–789, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347625.

Sun:1994:SCB

[SL94] Jiayang Sun and Clive R. Loader. Simultaneous confidence bands for linear regression and smoothing. Annals of Statistics, 22(3):1328–1345, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325631.

Simpson:1990:NTD

[SM90] Douglas G. Simpson and Barry H. Margolin. Nonparametric testing for dose-response curves subject to downturns: Asymptotic power considerations. Annals of Statistics, 18(1):373–390, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347505.

Schick:1994:RRE

[SM94] Irvin C. Schick and Sanjoy K. Mitter. Robust recursive estimation in the presence of heavy-tailed observation noise. Annals of Statistics, 22(2):1045–1080, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176325511.

Small:1997:MMA

[Sma97] Christopher G. Small. Multidimensional medians arising from geodesics on graphs. Annals of Statistics, 25(2):478–494, April 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1031833660.

Spokoiny:1996:AHT

[Spo96] V. G. Spokoiny. Adaptive hypothesis testing using wavelets. Annals of Statistics, 24(6):2477–2498, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1032181163.

Spokoiny:1998:EFD

[Spo98] V. G. Spokoiny. Estimation of a function with discontinuities via local polynomial fit with an adaptive window choice. Annals of Statistics, 26(4):1356–1378, August 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691246.

Smith:1996:BMS

[SQ96] Jim Q. Smith and Catriona M. Queen. Bayesian models for sparse probability tables. Annals of Statistics, 24(5):2178– 2198, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362316.

Sriram:1992:ISP

[Sri92] T. N. Sriram. An improved sequential procedure for estimating the regression parameter in regression models with symmetric errors. Annals of Statistics, 20(3):1441–1453, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348777.

Sriram:1994:IBC

[Sri94] T. N. Sriram. Invalidity of bootstrap for critical branching processes with immigration. Annals of Statistics, 22(2):1013–1023, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176325509.

Shao:1994:IJS

[SS94] Peter Yi-Shi Shao and William E. Strawderman. Improving on the James–Stein positive-part estimator. Annals of Statistics, 22(3):1517–1538, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325640.

Schmid:1997:SPE

[SS97] Wolfgang Schmid and Alexander Schöne. Some properties of the EWMA control chart in the presence of autocorrelation. Annals of Statistics, 25(3):1277–1283, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1069362748.

Seidenfeld:1995:RPO

[SSK95] Teddy Seidenfeld, Mark J. Schervish, and Joseph B. Kadane. A representation of partially ordered preferences. Annals of Statistics, 23(6):2168–2217, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1034713653.

Shaked:1992:CED

[ST92] Moshe Shaked and Y. L. Tong. Comparison of experiments via dependence of normal variables with a common marginal distribution. Annals of Statistics, 20(1):614–618, March 1992. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176348547.

Strawderman:1996:APF

[ST96] Robert L. Strawderman and Anastasios A. Tsiatis. On the asymptotic properties of a flexible hazard estimator. Annals of Statistics, 24(1):41–63, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1033066198.

Stafford:1996:RAP

[Sta96] James E. Stafford. A robust adjustment of the profile likelihood. Annals of Statistics, 24(1):336–352, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1033066212.

Stein:1990:DAP

[Ste90a] Charles Stein. Discussion: An ancillarity paradox which appears in multiple linear regression. Annals of Statistics, 18 (2):530–533, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347613.

Stein:1990:UAO

[Ste90b] Michael Stein. Uniform asymptotic optimality of linear predictions of a random field using an incorrect second-order structure. Annals of Statistics, 18(2):850–872, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347629.

Stein:1990:BEL

[Ste90c] Michael L. Stein. Bounds on the efficiency of linear predictions using an incorrect covariance function. Annals of Statistics, 18(3):1116–1138, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347742.

Stein:1990:CGC

[Ste90d] Michael L. Stein. A comparison of generalized cross validation and modified maximum likelihood for estimating the parameters of a stochastic process. Annals of Statistics, 18(3):1139– 1157, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347743.

Stein:1993:SSE

[Ste93] Michael L. Stein. Spline smoothing with an estimated order parameter. Annals of Statistics, 21(3):1522–1544, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349270.

Stepniak:1994:NCG

[Ste94] C. Stepniak. A note on comparison of genetic experiments. Annals of Statistics, 22(3):1630–1632, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325646.

Stefanov:1995:ELR

[Ste95a] Valeri T. Stefanov. Explicit limit results for minimal sufficient statistics and maximum likelihood estimators in some Markov processes: Exponential families approach. Annals of Statistics, 23(4):1073–1101, August 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324699.

Stein:1995:LLS

[Ste95b] Michael L. Stein. Locally lattice sampling designs for isotropic random fields. Annals of Statistics, 23(6):1991–2012, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/

1034713644. See correction [Ste99].

Stein:1995:PIR

[Ste95c] Michael L. Stein. Predicting integrals of random fields using observations on a lattice. Annals of Statistics, 23(6):1975– 1990, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1034713643.

Steif:1997:CEJ

[Ste97] Jeffrey E. Steif. Consistent estimation of joint distributions for sufficiently mixing random fields. Annals of Statistics, 25(1): 293–304, February 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1034276630.

Stein:1999:CLL

[Ste99] Michael L. Stein. Correction: “Locally lattice sampling designs for isotropic random fields”. Annals of Statistics, 27(4):1440, August 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1017938933. See [Ste95b].

Stone:1990:LSI

[Sto90] Charles J. Stone. Large-sample inference for log-spline models. Annals of Statistics, 18(2):717–741, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347622.

Stone:1991:ADF

[Sto91a] Charles J. Stone. Asymptotics for doubly flexible logspline response models. Annals of Statistics, 19(4):1832–1854, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348373.

Stone:1991:DMA

[Sto91b] Charles J. Stone. Discussion: Multivariate adaptive regression splines. Annals of Statistics, 19(1):113–115, March 1991. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176347971.

Stone:1994:RUP

[Sto94a] Charles J. Stone. Rejoinder: The use of polynomial splines and their tensor products in multivariate function estimation. Annals of Statistics, 22(1):179–184, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325364.

Stone:1994:UPS

[Sto94b] Charles J. Stone. The use of polynomial splines and their tensor products in multivariate function estimation. Annals of Statistics, 22(1):118–171, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176325361.

Streitberg:1990:LIR

[Str90] Bernd Streitberg. Lancaster interactions revisited. Annals of Statistics, 18(4):1878–1885, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347885.

Strasser:1996:AEE

[Str96] Helmut Strasser. Asymptotic efficiency of estimates for models with incidental nuisance parameters. Annals of Statistics, 24 (2):879–901, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032894471.

Streitberg:1999:EIH

[Str99] Bernd Streitberg. Exploring interactions in high-dimensional tables: a bootstrap alternative to log-linear models. Annals of Statistics, 27(1):405–413, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1018031118.

Shirakura:1996:SPN

[STS96] Teruhiro Shirakura, Tadashi Takahashi, and Jagdish N. Srivastava. Searching probabilities for nonzero effects in search designs for the noisy case. Annals of Statistics, 24(6):2560– 2568, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032181169.

Stute:1993:ASR

[Stu93] Winfried Stute. Almost sure representations of the productlimit estimator for truncated data. Annals of Statistics, 21 (1):146–156, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349019.

Stute:1994:UCC

[Stu94] Winfried Stute. Universally consistent conditional U-statistics. Annals of Statistics, 22(1):460–473, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325378.

Stute:1995:CLT

[Stu95] Winfried Stute. The central limit theorem under random censorship. Annals of Statistics, 23(2):422–439, April 1995. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176324528.

Stute:1996:JEV

[Stu96] Winfried Stute. The jackknife estimate of variance of a Kaplan– Meier integral. Annals of Statistics, 24(6):2679–2704, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032181175.

Stute:1997:NMC

[Stu97] Winfried Stute. Nonparametric model checks for regression. Annals of Statistics, 25(2):613–641, April 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1031833666.

Stute:1998:MCR

[STZ98] Winfried Stute, Silke Thies, and Li-Xing Zhu. Model checks for regression: an innovation process approach. Annals of Statistics, 26(5):1916–1934, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691363.

Sun:1994:IEP

[Sun94] Dongchu Sun. Integrable expansions for posterior distributions for a two-parameter exponential family. Annals of Statistics, 22(4):1808–1830, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325758.

Sun:1996:SET

[Sun96] Li Sun. Shrinkage estimation in the two-way multivariate normal model. Annals of Statistics, 24(2):825–840, April 1996. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1032894468.

Siegmund:1995:UGL

[SV95] D. Siegmund and E. S. Venkatraman. Using the generalized likelihood ratio statistic for sequential detection of a changepoint. Annals of Statistics, 23(1):255–271, February 1995. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176324466.

Snapp:1998:AEN

[SV98] Robert R. Snapp and Santosh S. Venkatesh. Asymptotic expansions of the k nearest neighbor risk. Annals of Statistics, 26(3):850–878, June 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1024691080.

Severini:1992:PLC

[SW92a] Thomas A. Severini and Wing Hung Wong. Profile likelihood and conditionally parametric models. Annals of Statistics, 20(4):1768–1802, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348889.

Shao:1992:APB

[SW92b] Jun Shao and C. F. J. Wu. Asymptotic properties of the balanced repeated replication method for sample quantiles. Annals of Statistics, 20(3):1571–1593, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348785.

Seidenfeld:1993:DSP

[SW93a] Teddy Seidenfeld and Larry Wasserman. Dilation for sets of probabilities. Annals of Statistics, 21(3):1139–1154, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349254.

Srivastava:1993:CEC

[SW93b] M. S. Srivastava and Yanhong Wu. Comparison of EWMA, CUSUM and Shiryayev–Roberts procedures for detecting a shift in the mean. Annals of Statistics, 21(2):645–670, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349142.

Stute:1993:SLU

[SW93c] W. Stute and J.-L. Wang. The strong law under random censorship. Annals of Statistics, 21(3):1591–1607, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349273.

Shen:1994:CRS

[SW94a] Xiaotong Shen and Wing Hung Wong. Convergence rate of sieve estimates. Annals of Statistics, 22(2):580–615, June 1994. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176325486.

Srivastava:1994:DSP

[SW94b] M. S. Srivastava and Yanhong Wu. Dynamic sampling plan in Shiryayev–Roberts procedure for detecting a change in the drift of Brownian motion. Annals of Statistics, 22(2):805–823, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325496.

Siegmund:1995:TSU

[SW95] David O. Siegmund and Keith J. Worsley. Testing for a signal with unknown location and scale in a stationary Gaussian random field. Annals of Statistics, 23(2):608–639, April 1995. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176324539.

Sun:1999:TUV

[SW99] Jiayang Sun and Michael Woodroofe. Testing uniformity versus a monotone density. Annals of Statistics, 27(1):338–360, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1018031114.

Swanepoel:1999:LBM

[Swa99] Jan W. H. Swanepoel. The limiting behavior of a modified maximal symmetric 2s-spacing with applications. Annals of Statistics, 27(1):24–35, February 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1018031099.

Sweeting:1992:AAC

[Swe92] Trevor J. Sweeting. Asymptotic ancillarity and conditional inference for stochastic processes. Annals of Statistics, 20(1): 580–589, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348542.

Shen:1998:LAR

[SWZ98] X. Shen, D. A. Wolfe, and S. Zhou. Local asymptotics for regression splines and confidence regions. Annals of Statistics, 26(5):1760–1782, October 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1024691356.

Simpson:1998:FSO

[SY98] Douglas G. Simpson and Victor J. Yohai. Functional stability of one-step GM-estimators in approximately linear regression. Annals of Statistics, 26(3):1147–1169, June 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1024691092.

Simons:1999:AWN

[SY99] Gordon Simons and Yi-Ching Yao. Asymptotics when the number of parameters tends to infinity in the Bradley–Terry model for paired comparisons. Annals of Statistics, 27(3):1041–1060, June 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1018031267.

Schumacher:1993:UPI

[SZ93a] Peter Schumacher and James V. Zidek. Using prior information in designing intervention detection experiments. Annals of Statistics, 21(1):447–463, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349036.

Siegmund:1993:ENL

[SZ93b] David Siegmund and Heping Zhang. The expected number of local maxima of a random field and the volume of tubes. Annals of Statistics, 21(4):1948–1966, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349404.

Scheike:1998:CRF

[SZ98] Thomas H. Scheike and Mei-Jie Zhang. Cumulative regression function tests for regression models for longitudinal data. Annals of Statistics, 26(4):1328–1355, August 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1024691245.

Szkutnik:1992:SCH

[Szk92] Zbigniew Szkutnik. Special capacities, the Hunt–Stein theorem and transformation groups. Annals of Statistics, 20(2): 1120–1128, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348674.

Thomas-Agnan:1991:SFS

[TA91] Christine Thomas-Agnan. Spline functions and stochastic filtering. Annals of Statistics, 19(3):1512–1527, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348259.

Takada:1992:SPA

[Tak92] Yoshikazu Takada. A sequential procedure with asymptotically negative regret for estimating a normal mean. Annals of Statistics, 20(1):562–569, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348540.

Takagi:1994:LAM

[Tak94] Yoshiji Takagi. Local asymptotic minimax risk bounds for asymmetric loss functions. Annals of Statistics, 22(1):39–48, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176325356.

Tang:1993:MRD

[Tan93] Dei-In Tang. Minimax regression designs under uniform departure models. Annals of Statistics, 21(1):434–446, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349035.

Tseng:1997:GEC

[TB97] Yu-Ling Tseng and Lawrence D. Brown. Good exact confidence sets for a multivariate normal mean. Annals of Statistics, 25(5): 2228–2258, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362396.

Tsai:1985:LSS

[TC85] Wei-Yann Tsai and John Crowley. A large sample study of generalized maximum likelihood estimators from incomplete data via self-consistency. Annals of Statistics, 13(4):1317–1334, December 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176349740. See correction [TC90].

Tsai:1990:CLS

[TC90] Wei-Yann Tsai and John Crowley. Correction: “A Large Sample Study of Generalized Maximum Likelihood Estimators from Incomplete Data via Self-Consistency”. Annals of Statistics, 18(1):470, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347514. See [TC85].

Thompson:1993:GPP

[Tho93] G. L. Thompson. Generalized permutation polytopes and exploratory graphical methods for ranked data. Annals of Statistics, 21(3):1401–1430, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176349265.

Thompson:1996:LLF

[Tho96] E. A. Thompson. Likelihood and linkage: from Fisher to the future. Annals of Statistics, 24(2):449–465, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1032894448.

Tierney:1994:MCE

[Tie94a] Luke Tierney. Markov chains for exploring posterior distributions. Annals of Statistics, 22(4):1701–1728, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325750;   http://www.jstor.org/stable/2242477.

Tierney:1994:RMC

[Tie94b] Luke Tierney. Rejoinder: Markov chains for exploring posterior distributions. Annals of Statistics, 22(4):1758–1762, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325755.

Tjur:1991:BDE

[Tju91] Tue Tjur. Block designs and electrical networks. Annals of Statistics, 19(2):1010–1027, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348134.

Takemura:1997:WDS

[TK97] Akimichi Takemura and Satoshi Kuriki. Weights of χ̄^{2} distribution for smooth or piecewise smooth cone alternatives. Annals of Statistics, 25(6):2368–2387, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1030741077.

Tarpey:1995:PPS

[TLF95] Thaddeus Tarpey, Luning Li, and Bernard D. Flury. Principal points and self-consistent points of elliptical distributions. Annals of Statistics, 23(1):103–112, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176324457.

Tang:1993:ESC

[TM93] S. M. Tang and I. B. MacNeill. The effect of serial correlation on tests for parameter change at unknown time. Annals of Statistics, 21(1):552–575, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176349042.

Tran:1993:NFE

[Tra93] Lanh Tat Tran. Nonparametric function estimation for time series by local average estimators. Annals of Statistics, 21(2): 1040–1057, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349163.

Tran:1996:FDR

[TRYV96] Lanh Tran, George Roussas, Sidney Yakowitz, and B. Truong Van. Fixed-design regression for linear time series. Annals of Statistics, 24(3):975–991, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1032526952.

Terrell:1992:VKD

[TS92a] George R. Terrell and David W. Scott. Variable kernel density estimation. Annals of Statistics, 20(3):1236–1265, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348768.

Truong:1992:NFE

[TS92b] Young K. Truong and Charles J. Stone. Nonparametric function estimation involving time series. Annals of Statistics, 20 (1):77–97, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348513.

Tsiatis:1990:ERP

[Tsi90] Anastasios A. Tsiatis. Estimating regression parameters using linear rank tests for censored data. Annals of Statistics, 18 (1):354–372, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347504.

Tsybakov:1997:NED

[Tsy97] A. B. Tsybakov. On nonparametric estimation of density level sets. Annals of Statistics, 25(3):948–969, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1069362732.

Tsybakov:1998:PSN

[Tsy98] A. B. Tsybakov. Pointwise and sup-norm sharp adaptive estimation of functions on the Sobolev classes. Annals of Statistics, 26(6):2420–2469, December 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1024691478.

Tsay:1990:APM

[TT90] Ruey S. Tsay and George C. Tiao. Asymptotic properties of multivariate nonstationary processes with applications to autoregressions. Annals of Statistics, 18(1):220–250, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347499.

Tang:1996:CMA

[TW96] Boxin Tang and C. F. J. Wu. Characterization of minimum aberration 2^{n−k} designs in terms of their complementary designs. Annals of Statistics, 24(6):2549–2559, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032181168.

Tyler:1994:FSB

[Tyl94] David E. Tyler. Finite sample breakdown points of projection based multivariate location and scatter statistics. Annals of Statistics, 22(2):1024–1044, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176325510.

Uhlig:1994:SWS

[Uhl94] Harald Uhlig. On singular Wishart and singular multivariate beta distributions. Annals of Statistics, 22(1):395–405, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325375.

Uddin:1997:UOD

[UM97] Nizam Uddin and John P. Morgan. Universally optimal designs with blocksize p × 2 and correlated observations. Annals of Statistics, 25(3):1189–1207, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1069362744.

VanKeilegom:1999:TTI

[VA99] Ingrid Van Keilegom and Michael G. Akritas. Transfer of tail information in censored regression models. Annals of Statistics, 27(5):1745–1784, October 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1017939150.

VanDerVaart:1991:DF

[Van91] Aad Van Der Vaart. On differentiable functionals. Annals of Statistics, 19(1):178–204, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347976.

VanderVaart:1996:EML

[Van96] Aad Van der Vaart. Efficient maximum likelihood estimation in semiparametric mixture models. Annals of Statistics, 24 (2):862–878, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032894470.

vandeGeer:1990:ERF

[vdG90] Sara van de Geer. Estimating a regression function. Annals of Statistics, 18(2):907–924, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347632.

vandeGeer:1993:HCC

[vdG93] Sara van de Geer. Hellinger-consistency of certain nonparametric maximum likelihood estimators. Annals of Statistics, 21(1):14–44, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349013.

vandeGeer:1995:EIM

[vdG95] Sara van de Geer. Exponential inequalities for martingales, with application to maximum likelihood estimation for counting processes. Annals of Statistics, 23(5):1779–1801, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176324323.

vanderLaan:1996:EEB

[vdL96] Mark J. van der Laan. Efficient estimation in the bivariate censoring model and repairing NPMLE. Annals of Statistics, 24(2):596–627, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032894454.

vanderLaan:1998:EER

[vdLM98] Mark J. van der Laan and Ian W. McKeague. Efficient estimation from right-censored data when failure indicators are missing at random. Annals of Statistics, 26(1):164–182, February 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1030563981.

vanderVaart:1994:MLE

[vdV94] Aad van der Vaart. Maximum likelihood estimation with partially censored data. Annals of Statistics, 22(4):1896–1916, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176325763.

vanderVaart:1996:MLE

[vdV96] Aad van der Vaart. Maximum likelihood estimation under a spatial sampling scheme. Annals of Statistics, 24(5):2049– 2057, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362309.

vanEs:1992:ALS

[vE92] Bert van Es. Asymptotics for least squares cross-validation bandwidths in nonsmooth cases. Annals of Statistics, 20 (3):1647–1657, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348790.

Vellaisamy:1992:IRS

[Vel92] P. Vellaisamy. Inadmissibility results for the selected scale parameters. Annals of Statistics, 20(4):2183–2191, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348913.

Valz:1995:CGF

[VMT95] Paul D. Valz, A. Ian McLeod, and Mary E. Thompson. Cumulant generating function and tail probability approximations for Kendall’s score with tied rankings. Annals of Statistics, 23(1): 144–160, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324460.

Vos:1991:GAD

[Vos91] Paul W. Vos. A geometric approach to detecting influential cases. Annals of Statistics, 19(3):1570–1581, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348262.

Verdinelli:1998:BGF

[VW98] Isabella Verdinelli and Larry Wasserman. Bayesian goodnessof-fit testing using infinite-dimensional exponential families. Annals of Statistics, 26(4):1215–1241, August 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1024691240.

vanZuijlen:1991:NTB

[vZ91] Martien C. A. van Zuijlen. Note on the tail behavior of general weighted empirical processes. Annals of Statistics, 19(2): 1102–1105, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348143.

Vardi:1992:LSS

[VZ92] Y. Vardi and Cun-Hui Zhang. Large sample study of empirical distributions in a random-multiplicative censoring model. Annals of Statistics, 20(2):1022–1039, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348668.

Vu:1997:GLR

[VZ97] H. T. V. Vu and S. Zhou. Generalization of likelihood ratio tests under nonstandard conditions. Annals of Statistics, 25 (2):897–916, April 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1031833677.

Walther:1997:GS

[Wal97a] Guenther Walther. Granulometric smoothing. Annals of Statistics, 25(6):2273–2299, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1030741072.

Walther:1997:MCS

[Wal97b] Guenther Walther. Monte Carlo sampling in dual space for approximating the empirical halfspace distance. Annals of Statistics, 25(5):1926–1953, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1069362379.

Wang:1996:ALS

[Wan96a] Jinde Wang. Asymptotics of least-squares estimators for constrained nonlinear regression. Annals of Statistics, 24(3):1316– 1326, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/
euclid.aos/1032526971.

Wang:1996:FEW

[Wan96b] Yazhen Wang. Function estimation via wavelet shrinkage for long-memory data. Annals of Statistics, 24(2):466–484, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032894449.

Wang:1999:BPE

[Wan99] Hsiuying Wang. Brown’s paradox in the estimated confidence approach. Annals of Statistics, 27(2):610–626, April 1999. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1018031210.

Wasserman:1990:PEB

[Was90] Larry Alan Wasserman. Prior envelopes based on belief functions. Annals of Statistics, 18(1):454–464, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176347511.

Wasserman:1992:IPD

[Was92] Larry Wasserman. Invariance properties of density ratio priors. Annals of Statistics, 20(4):2177–2182, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348912.

Wu:1993:KTE

[WC93] J. S. Wu and C. K. Chu. Kernel-type estimators of jump points and values of a regression function. Annals of Statistics, 21(3):1545–1566, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176349271.

Wefelmeyer:1996:QLM

[Wef96] Wolfgang Wefelmeyer. Quasi-likelihood models and optimal inference. Annals of Statistics, 24(1):405–422, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1033066217.

Wei:1992:PLS

[Wei92] C. Z. Wei. On predictive least squares principles. Annals of Statistics, 20(1):1–42, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348511.

Wei:1995:AEE

[Wei95] Xiaoying Wei. Asymptotically efficient estimation of the index of regular variation. Annals of Statistics, 23(6):2036–2058, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1034713646.

Welsh:1989:PE

[Wel89] A. H. Welsh. On M-processes and M-estimation. Annals of Statistics, 17(1):337–361, March 1989. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347021. See correction

[Wel90].

Welsh:1990:CPE

[Wel90] A. H. Welsh. Correction: “On M-Processes and M-Estimation”. Annals of Statistics, 18(3):1500, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/

1176347766. See [Wel89].

Wang:1996:CES

[WG96] Zhiming Wang and Joseph C. Gardiner. A class of estimators of the survival function from interval-censored data. Annals of Statistics, 24(2):647–658, April 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1032894457.

Wang:1997:ACD

[WG97] Kongming Wang and Theo Gasser. Alignment of curves by dynamic time warping. Annals of Statistics, 25(3):1251–1276, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1069362747.

Woodroofe:1990:SAE

[WH90] Michael Woodroofe and Janis Hardwick. Sequential allocation for an estimation problem with ethical costs. Annals of Statistics, 18(3):1358–1377, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347754.

Walter:1991:BEB

[WH91] G. G. Walter and G. G. Hamedani. Bayes empirical Bayes estimation for natural exponential families with quadratic variance functions. Annals of Statistics, 19(3):1191–1224, September 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348245.

Wijsman:1985:PAS

[Wij85] Robert A. Wijsman. Proper action in steps, with application to density ratios of maximal invariants. Annals of Statistics, 13 (1):395–402, March 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346600. See correction [Wij93].

Wijsman:1993:CPA

[Wij93] Robert A. Wijsman. Correction: “Proper Action in Steps, with Application to Density Ratios of Maximal Invariants”. Annals of Statistics, 21(4):2168–2169, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349418. See

[Wij85].

Wasserman:1990:BTC

[WK90] Larry A. Wasserman and Joseph B. Kadane. Bayes’ Theorem for Choquet capacities. Annals of Statistics, 18(3):1328– 1339, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176347752.

Wasserman:1992:SUP

[WK92] Larry Wasserman and Joseph B. Kadane. Symmetric upper probabilities. Annals of Statistics, 20(4):1720–1736, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348887.

Walker:1997:BSP

[WM97] Stephen Walker and Pietro Muliere. Beta-Stacy processes and a generalization of the Pólya-urn scheme. Annals of Statistics, 25(4):1762–1780, August 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1031594741.

Wang:1998:AOO

[WMC98] Jane-Ling Wang, Hans-Georg Müller, and William B. Capra. Analysis of oldest-old mortality: lifetables revisited. Annals of Statistics, 26(1):126–163, February 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1030563980.

Wood:1999:BMG

[Woo99] G. R. Wood. Binomial mixtures: geometric estimation of the mixing distribution. Annals of Statistics, 27(5):1706–1721, October 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1017939148.

Worsley:1995:ENP

[Wor95] K. J. Worsley. Estimating the number of peaks in a random field using the Hadwiger characteristic of excursion sets, with applications to medical images. Annals of Statistics, 23(2):640–669, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176324540.

Wong:1991:MLE

[WS91] Wing Hung Wong and Thomas A. Severini. On maximum likelihood estimation in infinite dimensional parameter spaces. Annals of Statistics, 19(2):603–632, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176348113.

Wong:1995:PIL

[WS95] Wing Hung Wong and Xiaotong Shen. Probability inequalities for likelihood ratios and convergence rates of sieve MLES. Annals of Statistics, 23(2):339–362, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176324524.

Wick:1997:EDA

[WS97] David Wick and Steven G. Self. Estimating disease attack rates in heterogeneous interacting populations, with applications to HIV vaccine trials. Annals of Statistics, 25(2):642–661, April 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1031833667.

Wu:1990:APJ

[Wu90] C. F. J. Wu. On the asymptotic properties of the jackknife histogram. Annals of Statistics, 18(3):1438–1452, September 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347759.

Wu:1995:ERP

[Wu95a] Colin O. Wu. Estimating the real parameter in a two-sample proportional odds model. Annals of Statistics, 23(2):376–395, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176324526.

Wu:1995:ARE

[Wu95b] Tiee-Jian Wu. Adaptive root n estimates of integrated squared density derivatives. Annals of Statistics, 23(5):1474–1495, October 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176324308.

Wu:1997:OED

[Wu97] Huaiqing Wu. Optimal exact designs on a circle or a circular arc. Annals of Statistics, 25(5):2027–2043, October 1997. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1069362385.

Wegkamp:1996:CLS

[WvdG96] Marten Wegkamp and Sara van de Geer. Consistency for the least squares estimator in nonparametric regression. Annals of Statistics, 24(6):2513–2523, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1032181165.

Wei:1990:EMB

[WW90] C. Z. Wei and J. Winnicki. Estimation of the means in the branching process with immigration. Annals of Statistics, 18(4):1757–1773, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176347876.

Weisberg:1994:AML

[WW94] S. Weisberg and A. H. Welsh. Adapting for the missing link. Annals of Statistics, 22(4):1674–1700, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325749.

Wahba:1995:SSA

[WWG^{+}95] Grace Wahba, Yuedong Wang, Chong Gu, Ronald Klein, and Barbara Klein. Smoothing spline ANOVA for exponential families, with application to the Wisconsin epidemiological study of diabetic retinopathy: the 1994 Neyman Memorial Lecture. Annals of Statistics, 23(6):1865–1895, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1034713638.

Xiang:1994:BEB

[Xia94] Xiaojing Xiang. On the Berry–Esseen bound for L-statistics in the non-I.D. case with applications to the estimation of location parameters. Annals of Statistics, 22(2):968–979, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176325506.

Xiang:1995:DSQ

[Xia95] Xiaojing Xiang. Deficiency of the sample quantile estimator with respect to kernel quantile estimators for censored data. Annals of Statistics, 23(3):836–854, June 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176324625.

Xu:1995:SCB

[XM95] Yi Xu and Adam T. Martinsek. Sequential confidence bands for densities. Annals of Statistics, 23(6):2218–2240, December 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1034713654.

Xu:1997:SEL

[Xu97] Jian-Lun Xu. Simultaneous estimation of location parameters for sign-invariant distributions. Annals of Statistics, 25(5): 2259–2272, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362397.

Xu:1995:NCE

[XY95] Jian-Lun Xu and Grace L. Yang. A note on a characterization of the exponential distribution based on a type II censored sample. Annals of Statistics, 23(3):769–773, June 1995. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176324621.

Yajima:1991:APL

[Yaj91] Yoshihiro Yajima. Asymptotic properties of the LSE in a regression model with long-memory stationary errors. Annals of Statistics, 19(1):158–177, March 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176347975.

Yakir:1995:NRL

[Yak95] Benjamin Yakir. A note on the run length to false alarm of a change-point detection policy. Annals of Statistics, 23(1): 272–281, February 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324467.

Yakir:1996:DSP

[Yak96a] Benjamin Yakir. Dynamic sampling policy for detecting a change in distribution, with a probability bound on false alarm. Annals of Statistics, 24(5):2199–2214, October 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1069362317.

Yakir:1996:LBA

[Yak96b] Benjamin Yakir. A lower bound on the ARL to detection of a change with a probability constraint on false alarm. Annals of Statistics, 24(1):431–435, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1033066219.

Yakir:1997:NOD

[Yak97] Benjamin Yakir. A note on optimal detection of a change in distribution. Annals of Statistics, 25(5):2117–2126, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1069362390.

Yakir:1998:ARL

[Yak98] Benjamin Yakir. On the average run length to false alarm in surveillance problems which possess an invariance structure. Annals of Statistics, 26(3):1198–1214, June 1998. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1024691095.

Yang:1991:MHD

[Yan91] Song Yang. Minimum Hellinger distance estimation of parameter in the random censorship model. Annals of Statistics, 19 (2):579–602, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348112.

Yang:1992:SIA

[Yan92] Song Yang. Some inequalities about the Kaplan–Meier estimator. Annals of Statistics, 20(1):535–544, March 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348537.

Yang:1997:GPL

[Yan97] Song Yang. A generalization of the product-limit estimator with an application to censored regression. Annals of Statistics, 25 (3):1088–1108, June 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362739.

Yang:1994:ECM

[YB94] Ruoyong Yang and James O. Berger. Estimation of a covariance matrix using the reference prior. Annals of Statistics, 22(3):1195–1211, September 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176325625.

Yang:1999:ITD

[YB99] Yuhong Yang and Andrew Barron. Information-theoretic determination of minimax rates of convergence. Annals of Statistics, 27(5):1564–1599, October 1999. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1017939142.

Ye:1997:NCE

[YD97] Jianming Ye and Naihua Duan. Nonparametric n^{−1/2}- consistent estimation for the general transformation models. Annals of Statistics, 25(6):2682–2717, December 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1030741091.

Yeh:1994:BVS

[Yeh94] Lam Yeh. Bayesian variable sampling plans for the exponential distribution with Type I censoring. Annals of Statistics, 22 (2):696–711, June 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176325491.

Yu:1995:AMU

[YG95] Qiqing Yu and Z. Govindarajulu. Admissibility and minimaxity of the UMVU estimator of P {X < Y }. Annals of Statistics, 23(2):598–607, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176324538.

Ying:1990:NSA

[Yin90] Zhiliang Ying. Nonlinear stochastic approximation procedures for L_{p} loss functions. Annals of Statistics, 18(4):1817–1828, December 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347880.

Ying:1992:MHT

[Yin92] Zhiliang Ying. Minimum Hellinger-type distance estimation for censored data. Annals of Statistics, 20(3):1361–1390, September 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348773.

Ying:1993:LSS

[Yin93a] Zhiliang Ying. A large sample study of rank estimation for censored regression data. Annals of Statistics, 21(1):76–99, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176349016.

Ying:1993:MLE

[Yin93b] Zhiliang Ying. Maximum likelihood estimation of parameters under a spatial sampling scheme. Annals of Statistics, 21(3):1567–1590, September 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176349272.

Yu:1992:MBI

[YP92] Qiqing Yu and Eswar Phadia. Minimaxity of the best invariant estimator of a distribution function under the Kolmogorov– Smirnov loss. Annals of Statistics, 20(4):2192–2195, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348914.

Yu:1997:ICM

[YS97] Bin Yu and T. P. Speed. Information and the clone mapping of chromosomes. Annals of Statistics, 25(1):169–185, February 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1034276625.

Yu:1991:MED

[YsC91] Qiqing Yu and Mo suk Chow. Minimaxity of the empirical distribution function in invariant estimation. Annals of Statistics, 19(2):935–951, June 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348129.

Yu:1993:DEN

[Yu93] Bin Yu. Density estimation in the L^{∞} norm for dependent data with applications to the Gibbs sampler. Annals of Statistics, 21(2):711–735, June 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349146.

Yohai:1993:MBP

[YZ93] Victor J. Yohai and Ruben H. Zamar. A minimax-bias property of the least α-quantile estimates. Annals of Statistics, 21(4): 1824–1842, December 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176349400.

Zaman:1991:ASA

[Zam91] Asad Zaman. Asymptotic suprema of averaged random functions. Annals of Statistics, 19(4):2145–2159, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348390.

Zamar:1992:BRE

[Zam92] Ruben H. Zamar. Bias robust estimation in orthogonal regression. Annals of Statistics, 20(4):1875–1888, December 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176348893.

Zhao:1997:EBC

[ZBCL97] Lincheng Zhao, Zhidong Bai, Chern-Ching Chao, and Wen-Qi Liang. Error bound in a central limit theorem of doubleindexed permutation statistics. Annals of Statistics, 25(5): 2210–2227, October 1997. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1069362395.

Zhu:1996:AKE

[ZF96] Li-Xing Zhu and Kai-Tai Fang. Asymptotics for kernel estimate of sliced inverse regression. Annals of Statistics, 24(3): 1053–1068, June 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1032526955.

Zhang:1990:FME

[Zha90] Cun-Hui Zhang. Fourier methods for estimating mixing densities and distributions. Annals of Statistics, 18(2):806–831, June 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176347627.

Zhang:1991:VSN

[Zha91] Ping Zhang. Variable selection in nonparametric regression with continuous covariates. Annals of Statistics, 19(4):1869– 1882, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176348375.

Zhang:1993:MSM

[Zha93] Ping Zhang. Model selection via multifold cross validation. Annals of Statistics, 21(1):299–313, March 1993. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176349027.

Zhang:1994:CRL

[Zha94] Heping Zhang. Confidence regions in linear functional relationships. Annals of Statistics, 22(1):49–66, March 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325357.

Zhang:1995:EMD

[Zha95] Cun-Hui Zhang. On estimating mixing densities in discrete exponential family models. Annals of Statistics, 23(3):929–945, June 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176324629.

Zhao:1997:MLE

[Zha97] Linda H. Zhao. Minimax linear estimation in a white noise problem. Annals of Statistics, 25(2):745–755, April 1997. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1031833671.

Zhou:1991:SPK

[Zho91] Mai Zhou. Some properties of the Kaplan–Meier estimator for independent nonidentically distributed random variables. Annals of Statistics, 19(4):2266–2274, December 1991. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176348399.

Zhou:1992:ANS

[Zho92] Mai Zhou. Asymptotic normality of the ‘synthetic data’ regression estimator for censored survival data. Annals of Statistics, 20(2):1002–1021, June 1992. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176348667.

Zucker:1990:NSA

[ZK90] David M. Zucker and Alan F. Karr. Nonparametric survival analysis with time-dependent covariate effects: A penalized partial likelihood approach. Annals of Statistics, 18(1):329–353, March 1990. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176347503.

Zhang:1996:LRD

[ZL96] Cun-Hui Zhang and Xin Li. Linear regression with doubly censored data. Annals of Statistics, 24(6):2720–2743, December 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1032181177.

Zhou:1996:DUR

[ZP96] Kenneth Q. Zhou and Stephen L. Portnoy. Direct use of regression quantiles to construct confidence sets in linear models. Annals of Statistics, 24(1):287–306, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1033066210.